### v1

# Estimating treatment effects with competing intercurrent events in randomized controlled trials

Abstract
The analysis of randomized controlled trials is often complicated by intercurrent events—events that occur after treatment initiation and may impact outcome assessment. These events may lead to patients discontinuing their assigned treatment or dropping out of the trial entirely. In an analysis of data from two recent immunology trials, we categorize intercurrent events into two broad types: those unrelated to treatment (e.g., withdrawal from the study due to external factors like pandemics or relocation) and those related to treatment (e.g., adverse events or lack of efficacy). We adopt distinct strategies to handle each type, aiming to target a clinically more relevant estimand. For treatment-related intercurrent events, they often meaningfully describe the patient’s outcome, we employ a composite variable strategy, where we attribute an outcome value that reflects the lack of treatment success. For treatment-unrelated intercurrent events, we adopt a hypothetical strategy that assumes these event times are conditionally independent of the outcome, given treatment and covariates, and envisions a scenario in which the intercurrent events do not occur. We establish the nonparametric identification and semiparametric estimation theory for the causal estimand and introduce doubly robust estimators. We illustrate our methods through the re-analysis of two randomized trials on baricitinib for Systemic Lupus Erythematosus. We classify intercurrent events, apply four estimators, and compare our approach with common ad-hoc methods, highlighting the robustness and practical implications of our framework.
Keywords: Causal inference; Censoring; Estimand; Post-treatment variable; Potential outcome
1 Intercurrent events in randomized controlled trials
Randomized controlled trials (RCTs) are considered the gold standard for evaluating treatment efficacy. However, after treatment initiation, various events—referred to as intercurrent events (ICEs)—can arise, impacting the interpretation or availability of the outcome of interest, and posing significant challenges to the analysis of RCTs. These ICEs can include treatment discontinuation due to adverse events, lack of efficacy, or study withdrawal due to patient relocation. Addressing these ICEs carefully is crucial to ensure the validity and reliability of the causal conclusions drawn from RCTs.
Given the complexities introduced by ICEs, clear guidance on how to handle them is crucial for drawing valid causal conclusions. The International Conference on Harmonization (ICH) has established global standards for drug development to align practices among pharmaceutical companies and regulatory agencies. Recognizing the importance of well-defined estimands in the presence of ICEs, the ICH introduced the E9 (R1) Addendum (ICH E9 , 2019, R1), a key resource that provides updated statistical principles to support clear and consistent approaches to analyzing and interpreting RCT results. The guidance has become essential for companies in drug research and development, and various strategies have been proposed and utilized to address ICEs following the guidance (Qu et al.,, 2021; Han and Zhou,, 2023).
One straightforward approach to address the complications from the ICEs is the intention-to-treat (ITT) analysis. The ITT strategy includes all participants in their originally assigned groups and uses the observed outcome value, regardless of whether they completed the treatment or adhered to the study protocol. This strategy, also known as the treatment policy strategy in ICH E9 , 2019 (R1), is relevant for policy decisions and widely accepted. However, it fails to address ICEs that are terminal events, such as death, since these events directly impact the existence of the outcome variable.
To address the limitation of ITT analyses and approach the issue from a different perspective, the composite variable strategy modifies the causal parameter of interest from the commonly used average causal effect to a comparison between two composite potential outcomes across the treatment arms. The composite potential outcome is defined as the endpoint of interest if no ICE occurs before its measurement, and a pre-specified, clinically meaningful value indicates treatment failure otherwise. The composite variable strategy is widely used in practice across different types of endpoints: binary endpoints (e.g., the non-responder imputation approach), ordinal or continuous endpoints (e.g., Rosenbaum,, 2006), and time-to-event endpoints (e.g., progression-free survival). In addition to the treatment policy strategy and the composite outcome strategy, three additional strategies for addressing ICEs were provided in ICH E9 , 2019 (R1). The hypothetical strategy considers treatment effects under a counterfactual scenario where ICEs do not occur or occur differently. The remaining two strategies–the while-on-treatment strategy and the principal stratification strategy–are less commonly used in practice. ICH E9 , 2019 (R1) recommends using different strategies based on the specific type of ICEs involved. However, very few papers have formally explored how to combine different strategies and rigorously addressed the pitfalls of doing so naively (Lipkovich et al.,, 2020; Qu et al.,, 2021).
1.1 A phase-3 immunology trial
To illustrate our setup and notation, we first introduce a real phase-3 immunology trial (Morand et al.,, 2023; Petri et al.,, 2023), which we will revisit in Section 6. This trial investigates the causal effect of baricitinib versus placebo on Systemic Lupus Erythematosus. Patients in the RCT were assigned either to a baricitinib treatment group or to a placebo control group.
The primary endpoint of interest is an immune response index measured 52 weeks after treatment initiation. Ideally, we would compare this outcome across the two groups at week 52. However, 429 out of 1535 (27.95%) patients experienced ICEs between treatment initiation and the outcome measurement, leading to missingness of the outcome. Figure 1 shows the types and proportions of these ICEs. During the 52 weeks, various ICEs happened. For instance, some patients discontinued the assigned treatment due to adverse events or lack of efficacy, some had protocol deviation and were excluded from the study, some of them withdrew from the study for unknown reasons, and some were lost to follow-up.
Figure 1: Pie chart showing the ICE types and proportions
1.2 Our proposal and contribution
We propose to classify ICEs into two broad types: (1) treatment-unrelated ICEs, such as treatment discontinuation due to relocation or COVID-19 lockdown, which are assumed to be independent of treatment efficacy conditional on the observed covariates; and (2) treatment-related ICEs, such as treatment discontinuation due to adverse events or lack of efficacy, use of rescue medication, and terminating events such as death, which are likely correlated with both treatment and outcomes. Figure 1 illustrates these classifications: yellow solid charts represent treatment-related ICEs, the green solid chart represents a treatment-unrelated ICE, and blue stripped charts represent a mix of both types. For instance, patients withdrew from the study for various reasons–some due to relocation or the pandemic, which can be considered as unrelated to the treatment, while others withdrew due to concerns about potential side effects, which are likely to be treatment-related. When patients did not disclose a detailed reason, we adopt a conservative approach and classify their ICEs as treatment-related.
Both types of ICEs may occur between treatment initiation and outcome measurement. Figure 2 depicts the timeline and three scenarios in our motivating example. Before week 52, patient 1 withdrew from the study due to relocation, which is arguably unrelated to the treatment, patient 2 dropped out of the study due to an adverse event, and patient 3 completed the study through the endpoint at week 52, when the outcome of interest was measured.
Figure 2: Illustration of the motivating immunology trial example
Both types of ICEs lead to missing primary endpoints measured at a landmark time because the outcome cannot be observed or is not even well-defined when an ICE occurs. For the first type of ICE, we apply the hypothetical strategy and assume the ICE event happens at random conditional on the observed covariates. It considers the hypothetical scenario in which the ICE did not occur. For the second category, we utilize the composite variable strategy, incorporating the ICE itself as indicative of treatment failure, for which we provide a formal definition and discussion in Section 3.
A main challenge in combining the two strategies and our key insight is that the two types of ICEs can compete with each other, meaning that patients discontinue due to the first reason that occurs. Use study withdrawal due to relocation and adverse events as examples to represent the two types of ICEs, respectively. For example, suppose a patient discontinues due to relocation like patient 1 in Figure 2. In that case, it remains unknown whether an adverse event would have occurred before the final endpoint measurement in the counterfactual scenario had the withdrawal due to relocation ICE not occurred. It is important to carefully deal with the issue of competing ICEs, which, to the best of our knowledge, has never been pointed out in the literature. We consider this feature and provide the nonparametric identification of our causal parameter of interest under two assumptions: (1) treatment assignment satisfies the ignorability and overlap assumptions, and (2) the treatment-unrelated ICE time is independent of the potential outcomes and the potential values of the treatment-related ICE time, given observed covariates. We present two identification formulas relying on different sets of nuisance models. We then propose two basic estimators based on these two identification formulas and a third estimator that combines them in a similar way to the construction of the augmented inverse probability weighting estimator to achieve double robustness.
The combined estimator can be further improved in terms of both robustness and asymptotic efficiency. To construct a more principled estimator that simultaneously achieves robustness and semiparametric efficiency, we derive the efficient influence function (EIF) thus the semiparametric efficiency bound of our causal parameter. We propose an EIF-based estimator and study its asymptotic properties. We show the EIF estimator is doubly robust, meaning its consistency does not require all nuisance parameters to be consistently estimated—only one of the two sets of nuisance models needs to be correctly specified. Moreover, the EIF estimator achieves the semiparametric efficiency bound under certain conditions on nuisance estimation.
The remaining of the paper proceeds as follows. In Section 2, we provide a motivating example, introduce the basic setup of our research question, describe the composite variable strategy, define our causal parameter of interest, and illustrate the identification challenge. In Section 3, we present the key identification assumptions, provide the nonparametric identification of our causal parameter, and construct three basic estimators. In Section 4, we derive the EIF for our causal parameter, propose a doubly robust and asymptotically efficient estimator based on the EIF, and study its asymptotic properties. In Section 5, we run Monte Carlo simulations to evaluate the finite sample performance of our proposed estimation procedures. In Section 6, we implement the estimators to re-analyze two randomized controlled trials that study the causal effect of baricitinib treatment on systemic lupus erythematosus. In Section 7, we discuss two future research directions.
We use the following notation. Let $\lVert r\rVert_{2}=\{\int r(o)^{2}\textnormal{d}P(o)\}^{1/2}$ denote the $L_{2}(P)$ norm where $P(\cdot)$ denotes the distribution of the observed data $O=o$. For the survival and censoring functions, let $\lVert r\rVert_{2}=\{\iint r(t,o)^{2}\textnormal{d}P(o)\textnormal{d}t\}^{1/2}$ denote the $L_{2}(P)$ norm. We write $b_{n}=O_{P}(a_{n})$ if $b_{n}/a_{n}$ is bounded in probability and $b_{n}=o_{P}(a_{n})$ if $b_{n}/a_{n}$ converges to 00 in probability.
2 Basic setup
2.1 Setup and notation
In this subsection, we introduce our formal setup and mathematical notation. Let $A$ denote the binary treatment indicator, with $A=1$ for patients in the treatment group, and $A=0$ for those in the control group. The primary outcome, denoted as $Y$, is measured at a pre-specified time point $k$. In our motivating example, $Y$ is the immune response index and $k$ is 52 weeks. We adopt the potential outcomes framework (Neyman,, 1923; Rubin,, 1974), where $Y(a)$ denotes the potential outcome under treatment assignment $A=a$ for $a=0,1$, and $Y=Y(A)=AY(1)+(1-A)Y(0)$ is the observed outcome.
As discussed, two types of ICEs may occur between the treatment initiation and the measurement of the outcome: treatment-unrelated ICEs, assumed independent of the potential outcomes conditional on covariates, and treatment-related ICEs, which likely correlate with the endpoints. To formally define the two types of ICEs, let $C$ and $T$ denote time to treatment-unrelated and time-to-treatment-related ICEs, respectively. Both are post-treatment variables, thus we use $C(a)$ and $T(a)$ to denote their potential values under treatment $a$, for $a=0,1$, with observed values $C=C(A)=AC(1)+(1-A)C(0)$ and $T=T(A)=AT(1)+(1-A)T(0)$.
2.2 Causal parameter of interest and challenge in identification
To address treatment-related ICEs, we define a composite potential outcome as $Y^{\textup{cp}}(a)=Y(a)1\{T(a)>k\}$ for $a=0,1$ where the superscript stands for “composite” and define our parameter of interest as the mean contrast, i.e.,
$\tau=E\{Y^{\textup{cp}}(1)-Y^{\textup{cp}}(0)\}=E[Y(1)1\{T(1)>k\}]-E[Y(0)1\{T(0)>k\}].$ (1)
This target parameter is widely used for binary outcomes to capture success or failure, where treatment-related ICEs, such as adverse events, are treated as failure (ICH E9 , 2019, R1). In our motivating example, $Y(a)$ measures whether a patient is an immune responder at week 52 after taking the treatment $A=a$, which is unobserved for patients experiencing a treatment-related ICE such as an adverse event or lack of efficacy. Such patients are defined as non-responders under the composite variable strategy.
The composite variable approach extends beyond binary outcomes to accommodate scenarios where a clinically significant value can denote treatment failure. More specifically, for outcomes such as quality of life, chronic pain, physical functioning, and cognitive performance that are measured on the ordinal and continuous scale, if there is a predefined clinically meaningful value $v$ that can signal treatment failure, then a composite outcome under treatment arm $A=a$ may be constructed as $Y(a)1\{T(a)>k\}+v1\{T(a)\leq k\}$, in which case the causal parameter is defined as in (1) with $Y(a)$ replaced by $Y(a)-v$ for $a=0,1$. Rosenbaum, (2006) discussed similar causal parameters when the ICE is death and the outcome of interest is some measure of life quality after a certain period.
Identifying $\tau$ in the presence of both ICE types is challenging. For clarity, in the remainder of this subsection, we focus on withdrawal due to relocation and adverse events to represent the treatment-unrelated and treatment-related ICEs, respectively. The definition of the causal parameter $\tau$ in (1) essentially treats $Y^{\textup{cp}}(a)$ as the relevant outcome, assigning 0 when adverse events occur, i.e., when $T(a)<k$. If adverse events were the only reason for missingness in the outcome, we could identify $\tau$ using the difference in means of the observed composite outcomes $Y^{\textup{cp}}=Y^{\textup{cp}}(A)=Y1(T>k)$, i.e.,
$\tau$ $=$ $E\{Y1(T>k)\mid A=1\}-E\{Y1(T>k)\mid A=0\}.$
However, this formula is infeasible to learn from observed data because treatment-unrelated ICEs, such as withdrawal due to relocation, could happen before treatment-related ICEs thus censoring $T(a)$. Figure 2 shows how withdrawal due to relocation in patient 1 censored their adverse event time and the outcome, creating missingness in the composite outcome $Y1(T>k)$ for such patients. The key challenge is that the two types of ICEs compete with each other, and patients discontinue due to the first ICE that occurs. Table 1 summarizes observed ICE types and composite outcomes, highlighting that censoring complicates the identification of $\tau$. When there is no ICE, we observe the outcome $Y$, which equals the composite outcome because $T>k$. When we observe an adverse event, the composite outcome is by definition 0. When we observe a withdrawal due to relocation, both the outcome of interest and the time-to-adverse event are missing, and so is the composite outcome. We address the challenge and present identification results in the following section.
Table 1: Summary of the observed ICE types and outcome. $C$ is the time to withdrawal due to relocation (WR), $T$ is the time to adverse event (AE), $Y$ is the outcome of interest, and $k$ is the pre-specified time point when the measurement of $Y$ is taken.
observed ICE type $(T,C,k)$-relationship $Y$ $1(T>k)$ composite outcome $Y^{\textup{cp}}$
no AE/WR $C\wedge T>k$ $Y$ 1 $Y$
AE $C\wedge k>T$ 0 0
WR $T\wedge k>C$
3 Identification and three basic estimators
3.1 Assumptions
We assume the joint distribution of $\{X_{i},A_{i},T_{i}(1),T_{i}(0),C_{i}(1),C_{i}(0),Y_{i}(1),Y_{i}(0)\}$ for patient $i$ is independently and identically distributed from a super population and will drop the subscript $i$ whenever there is no confusion.
We first assume the following ignorability and overlap assumption for the identification of our causal parameter of interest.
Assumption 1 (Treatment assignment)
We assume the following conditional independence and overlap conditions for the treatment assignment:
  • (a)
    $A\perp\!\!\!\perp\{Y(a),T(a),C(a)\}\mid X$ for $a=0,1$.
  • (b)
    For some constant $\eta\in(0,0.5)$, $\eta<e(X)<1-\eta$ with probability 1, where $e(X)=\textup{pr}(A=1\mid X)$ denotes the propensity score.
Assumption 1 is guaranteed by the design of an RCT, allowing our results to apply directly to RCTs like our motivating example. However, our formulation is more general as it can also accommodate observational studies. Under Assumption 1, we have the following equations for $\tau$
$\tau$ $=$ $E[E\{Y1(T>k)\mid A=1,X\}-E\{Y1(T>k)\mid A=0,X\}]$
$=$ $E\left\{\frac{AY1(T>k)}{e(X)}-\frac{(1-A)Y1(T>k)}{1-e(X)}\right\},$
which are the classic outcome regression and the inverse probability weighting identification formulas when we treat the composite potential outcomes $Y^{\textup{cp}}(a)=Y(a)1\{T(a)>k\}$ as the “new” outcome variable of interest with $Y^{\textup{cp}}=Y1(T>k)$. However, as discussed in Section 2.2, they do not serve as identification formulas anymore with the presence of the competing treatment-unrelated ICEs because $Y1(T>k)$ is no longer fully observed if treatment-unrelated ICE happened before both $T$ and $k$. Therefore, we need further identification assumptions on the time of treatment-unrelated ICE. Specifically, we assume it is independent of all other potential values, including the potential outcomes, which is equivalent to the censoring at random assumption in the survival analysis literature when we treat treatment-unrelated ICEs as censoring. In addition, we need the censoring probability to be bounded away from 1. We summarize them in the following assumption.
Assumption 2 (Time of treatment-unrelated ICE)
We assume the following conditional independence and positivity conditions for the time of treatment-unrelated ICE.
  • (a)
    $C(a)\perp\!\!\!\perp\{Y(a),T(a)\}\mid X$ for $a=0,1$.
  • (b)
    For some constant $\eta_{C}>0$, $\textup{pr}\{C(a)>k\mid X\}>\eta_{C}$ with probability 1.
In our motivating example, Assumption 2 states that the time to treatment-unrelated ICEs, such as withdrawal due to relocation, is independent of both the potential outcome of whether a patient will be a responder at week 52 and the time to an adverse event, conditional on observed baseline covariates. In other words, after accounting for factors such as region, baseline corticosteroid use, and the physician global assessment score, the decision to withdraw due to relocation does not provide additional information about the patients’ eventual response or the time to adverse events. Additionally, it assumes that every patient has a nonzero probability of not experiencing a treatment-unrelated ICE.
3.2 Two identification formulas
Using the composite variable strategy for the ICEs that are related to the treatment, recall that we define the parameter of interest as:
$\tau=\mu_{1}-\mu_{0}=E[Y(1)1\{T(1)>k\}]-E[Y(0)1\{T(0)>k\}].$
Under Assumptions 1 and 2, we show the parameter of interest $\tau$ is nonparametrically identified and provide two identification formulas in the following theorem.
Theorem 1 (Nonparametric identification of $\tau$)
Under Assumptions 1 and 2, $\tau$ is nonparametrically identified by two distinct formulas. First,
$\tau$ $=$ $E\left\{\mu_{1}(X)S_{1}(k\mid X)-\mu_{0}(X)S_{0}(k\mid X)\right\},$ (2)
where $\mu_{a}(X)=E(Y\mid T\wedge C>k,X,A=a)$ is the conditional mean of observed outcome among those with no ICE and $A=a$ and $S_{a}(t\mid X)=\textup{pr}(T>t\mid X,A=a)$ is the conditional survival function of treatment-related ICE time among those with $A=a$. Second,
$\tau$ $=$ $E\left[\frac{AY1(T\wedge C>k)}{e(X)G_{1}(k\mid X)}-\frac{(1-A)Y1(T\wedge C>k)}{\left\{1-e(X)\right\}G_{0}(k\mid X)}\right],$ (3)
where $e(X)=\textup{pr}(A=1\mid X)$ is the propensity score and $G_{a}(t\mid X)=\textup{pr}(C>t\mid X,A=a)$ for $a=0,1$ is the survival function of $C$ in the subsample $A=a$.
The first formula (2) expresses $\tau$ using a standard outcome regression approach. It models the conditional mean of the outcome among individuals who have not experienced ICE by $k$, adjusting for the probability of this event occurring. Equation (3) provides an alternative identification using inverse probability weighting. It identifies $\tau$ by averaging observed outcomes only among individuals who have not experienced ICE by time $k$ while correcting for the bias due to differences in the probability of observing a composite outcome. The weights adjust for two key factors: $e(X)$ that accounts for difference in treatment assignment probabilities, and $G_{a}(k\mid X)$ that adjusts for the likelihood of censoring in each treatment group. For instance, units with a lower probability of being observed are up-weighted, while those more likely to remain observed are down-weighted.
Both equations (2) and (3) are valid identification formulas because the nuisance parameters involved are either functions of observed data or identified from observed data. More specifically, nuisance parameters $e(X)$ and $\mu_{a}(X)$ are functions of the observed data. The survival functions of the treatment-related time $S_{a}(t\mid X)$ and the censoring mechanism $G_{a}(t\mid X)$ for $t\leq k$, although not direct functions of observed data, are both identifiable due to the classic result on survival data with censoring at random (Robins and Rotnitzky,, 1992; Robins and Finkelstein,, 2000). For example, $S_{1}(t\mid X)=\textup{pr}(T>t\mid X,A=1)$ for $t\leq k$ can be estimated using observations in the treatment subgroup from treating $T\wedge C\wedge k$ as the observed event time and $1(C\wedge k\geq T)$ as the event status for the treatment-related ICE, and estimating using a Cox proportional hazard model (Cox,, 1972). We can use a similar procedure to estimate $G_{1}(t\mid X)$ by switching the roles of $T$ and $C$, and use similar procedures to estimate $S_{0}(t\mid X)$ and $G_{0}(t\mid X)$ by using observations in control subgroup.
3.3 Two basic estimators
Identification formulas (2) and (3) in Theorem 1 motivate two basic estimators based on estimating the corresponding nuisance functions. Let $\hat{S}_{a}(t\mid X)$ and $\hat{G}_{a}(t\mid X)$ denote the fitted models for $t\leq k$ and $a=0,1$, $\hat{\mu}_{a}(X)$ denote the fitted outcome models for $a=0,1$, and $\hat{e}(X)$ denote the fitted propensity score model. Plugging in the fitted values of the nuisance parameters and taking empirical analogs of equations (2) and (3) motivate the following two estimators, respectively,
$\hat{\tau}^{\textup{out}}$ $=$ $n^{-1}\sum_{i=1}^{n}\hat{\mu}_{1}(X_{i})\hat{S}_{1}(k\mid X_{i})-n^{-1}\sum_{i=1}^{n}\hat{\mu}_{0}(X_{i})\hat{S}_{0}(k\mid X_{i}),$
$\hat{\tau}^{\textup{ipw}}$ $=$ $n^{-1}\sum_{i=1}^{n}\frac{A_{i}Y_{i}1(T_{i}\wedge C_{i}>k)}{\hat{e}(X_{i})\hat{G}_{1}(k\mid X_{i})}-n^{-1}\sum_{i=1}^{n}\frac{(1-A_{i})Y_{i}1(T_{i}\wedge C_{i}>k)}{\left\{1-\hat{e}(X_{i})\right\}\hat{G}_{0}(k\mid X_{i})}.$
The outcome regression estimator $\hat{\tau}^{\textup{out}}$ is consistent if the outcome model and the survival function are correctly specified. The weighting estimator $\hat{\tau}^{\textup{ipw}}$ is consistent if the propensity score model and the censoring mechanism are correctly specified.
3.4 Propensity score-augmented weighting estimator
Combining the two identification formulas (2) and (3), we have the following augmented weighting identification formula,
$\tau$ $=$ $E\left\{\frac{AY1(T\wedge C>k)}{e(X)G_{1}(k\mid X)}-\frac{A-e(X)}{e(X)}\mu_{1}(X)S_{1}(k\mid X)\right\}$ (4)
$-E\left[\frac{(1-A)Y1(T\wedge C>k)}{\left\{1-e(X)\right\}G_{0}(k\mid X)}-\frac{e(X)-A}{1-e(X)}\mu_{0}(X)S_{0}(k\mid X)\right].$
Based on (4), we construct an estimator that augments the weighting estimator $\hat{\tau}^{\textup{ipw}}$ using the estimated outcome models,
$\hat{\tau}^{\textup{aug}}$ $=$ $\hat{\tau}^{\textup{ipw}}-n^{-1}\sum_{i=1}^{n}\left\{\frac{A_{i}-\hat{e}(X_{i})}{\hat{e}(X_{i})}\hat{\mu}_{1}(X_{i})\hat{S}_{1}(k\mid X_{i})+\frac{A_{i}-\hat{e}(X_{i})}{1-\hat{e}(X_{i})}\hat{\mu}_{0}(X_{i})\hat{S}_{0}(k\mid X_{i})\right\}.$
It has a similar mathematical form to the classic augmented inverse probability weighting estimator for the average treatment effect (Bang and Robins,, 2005). We provide the properties of $\hat{\tau}^{\textup{aug}}$ in the following Proposition 1.
Proposition 1 (Double robustness of $\hat{\tau}^{\textup{aug}}$)
Suppose Assumptions 1 and 2 hold, and assume that $G_{a}(k\mid X)$ is correctly specified for $a=0,1$. $\hat{\tau}^{\textup{aug}}$ is a consistent estimator for $\tau$ if either the propensity score model $e(X)$ is correct, or both the outcome regression $\mu_{a}(X)$ and the survival models $S_{a}(k\mid X)$ are correct for $a=0,1$.
Proposition 1 shows that the consistency of $\hat{\tau}^{\textup{aug}}$ depends on the correct model of $G_{a}(k\mid X)$ for $a=0,1$ but remains robust to misspecification of the propensity score model $e(X)$. To build intuition, Equation (4) can be viewed as a projection of the weighting identification formula (3) onto the nuisance tangent space of the propensity score model $A\mid X$, thus we refer to $\hat{\tau}^{\textup{aug}}$ as a propensity score-augmented estimator. However, since it is not further projected onto the tangent space of the $G_{a}(k\mid X)$ model, the estimator does not retain robustness to misspecification of $G_{a}(k\mid X)$. We will further address the issue in Section 4, where we introduce an estimator that is robust to the misspecification of both models.
$\hat{\tau}^{\textup{aug}}$ dominates the weighting estimator $\hat{\tau}^{\textup{ipw}}$ in terms of robustness. The consistency of $\hat{\tau}^{\textup{aug}}$ is contingent on less restrictive requirements on nuisance parameter estimation in the sense that $\hat{\tau}^{\textup{aug}}$ is consistent whenever $\hat{\tau}^{\textup{ipw}}$ is. Consistency of $\hat{\tau}^{\textup{aug}}$ requires the correct specification of the censoring mechanism $G_{a}(k\mid X)$ for $a=0,1$. Given that $G_{a}(k\mid X)$ is correctly modeled, $\hat{\tau}^{\textup{aug}}$ is doubly robust because it is consistent if either the propensity score model is correct, or the outcome regression and the survival models are both correct. There is no clear dominance between $\hat{\tau}^{\textup{aug}}$ and $\hat{\tau}^{\textup{out}}$ as they require consistent estimation of different sets of nuisance parameters. For all three estimators $\hat{\tau}^{\textup{out}}$, $\hat{\tau}^{\textup{ipw}}$, and $\hat{\tau}^{\textup{aug}}$, we can construct variance estimators using nonparametric bootstrap.
4 Another doubly robust and semiparametrically efficient estimator
The propensity score-augmented weighting estimator $\hat{\tau}^{\textup{aug}}$ improves robustness. However, it is not robust to the misspecification of the censoring mechanism, nor does it achieve the semiparametric efficiency bound. In this section, we show that $\hat{\tau}^{\textup{aug}}$ can be further improved. We derive the semiparametric efficient influence function (EIF) for $\tau$ to learn the best asymptotic efficiency a consistent estimator of $\tau$ can achieve and propose an asymptotically efficient and doubly robust estimator based on the EIF.
We first describe the full and observed data structures in our setting and introduce some additional notation. Ideally, we want to observe the full data $(X,Y^{\textup{cp}}(1),Y^{\textup{cp}}(0))$. The missingness of the full data comes from two strings. First, for a given treatment arm $a=0,1$, the treatment-unrelated ICE time $C(a)$ censors the full data because $Y^{\textup{cp}}(a)$ is only observable if $C(a)\geq\{T(a)\wedge k\}$. Within each treatment group, we do not observe the full data due to censoring and only observe $(\Delta(a)=1\{C(a)\geq T(a)\wedge k\},\tilde{T}(a)=C(a)\wedge T(a)\wedge k,\Delta(a)Y^{\textup{cp}}(a))$. Second, the treatment assignment generates another level of missingness, because, for each observation, we never simultaneously observe both composite potential outcomes $\{Y^{\textup{cp}}(1),Y^{\textup{cp}}(0)\}$ even without censoring. Therefore, the observed data is $O=(A,\Delta=\Delta(A),\tilde{T}=\tilde{T}(A),\Delta Y^{\textup{cp}}=\Delta(A)Y^{\textup{cp}}(A))$.
4.1 EIF and estimator
In the following theorem, we provide the EIF for $\tau$.
Theorem 2 (EIF for $\tau$)
Under the nonparametric model, the EIF for $\mu_{1}$ is
$D_{1}(O)$ $=$ $\frac{A}{e(X)}\left\{\frac{Y1(T\wedge C>k)}{G_{1}(k\mid X)}+\mu_{1}(X)S_{1}(k\mid X)\int_{0}^{\tilde{T}}\frac{\textnormal{d}M_{G_{1}}(t)}{S_{1}(t\mid X)G_{1}(t\mid X)}\right\}$ (5)
$-\frac{A-e(X)}{e(X)}\mu_{1}(X)S_{1}(k\mid X)-\mu_{1},$
the EIF for $\mu_{0}$ is
$D_{0}(O)$ $=$ $\frac{1-A}{1-e(X)}\left\{\frac{Y1(T\wedge C>k)}{G_{0}(k\mid X)}+\mu_{0}(X)S_{0}(k\mid X)\int_{0}^{\tilde{T}}\frac{\textnormal{d}M_{G_{0}}(t)}{S_{0}(t\mid X)G_{0}(t\mid X)}\right\}$ (6)
$-\frac{e(X)-A}{1-e(X)}\mu_{0}(X)S_{0}(k\mid X)-\mu_{0},$
and thus, the EIF for $\tau$ is $D_{\tau}(O)=D_{1}(O)-D_{0}(O)$, where $\textnormal{d}M_{G_{a}}(t)=1(C\in\textnormal{d}t,\Delta=0)-1(\tilde{T}\geq t)\textnormal{d}\Lambda_{a}(t\mid X)$ with $\Lambda_{a}(t\mid X)$ denoting the conditional cumulative hazard function for the censoring $C$ in the treatment group $A=a$ for $a=0,1$.
The $M_{G_{a}}(t)$ in the EIF is the martingale constructed from the censoring counting process. Intuitively, $1(C\in\textnormal{d}t,\Delta=0)$ is the actual observed increment in the censoring counting process at time $t$, which records whether a censoring event has occurred, while $1(\tilde{T}\geq t)\textnormal{d}\Lambda_{a}(t\mid X)$ represents the expected increment in the counting process, given the history up to time $t$. The martingale $M_{G_{a}}(t)$ is essentially capturing the difference between the actual observed events and their expected occurrences. The EIF implies another identification formula for $\tau$ by the property that $E\{D_{\tau}(O)\}=0$. Rearranging terms, we have
$\tau$ $=$ $E\left\{\frac{AY1(T\wedge C>k)}{e(X)G_{1}(k\mid X)}-\frac{A-e(X)}{e(X)}\mu_{1}(X)S_{1}(k\mid X)\right\}$ (7)
$-E\left[\frac{(1-A)Y1(T\wedge C>k)}{\left\{1-e(X)\right\}G_{0}(k\mid X)}-\frac{e(X)-A}{1-e(X)}\mu_{0}(X)S_{0}(k\mid X)\right]$
$+E\left[\frac{A}{e(X)}\mu_{1}(X)S_{1}(k\mid X)\int_{0}^{\tilde{T}}\frac{\textnormal{d}M_{G_{1}}(t)}{S_{1}(t\mid X)G_{1}(t\mid X)}\right]$
$-E\left[\frac{1-A}{1-e(X)}\mu_{0}(X)S_{0}(k\mid X)\int_{0}^{\tilde{T}}\frac{\textnormal{d}M_{G_{0}}(t)}{S_{0}(t\mid X)G_{0}(t\mid X)}\right],$
where the first two lines are the same as in the propensity score-augmented inverse probability weighting identification formula (4). Intuitively, we further augment (4) by the term (7) to achieve robustness to misspecification of the censoring model. The EIFs in Theorem 2 motivate the following estimators for $\tau$,
$\hat{\tau}^{\textup{eif}}$ $=$ $\hat{\tau}^{\textup{aug}}+n^{-1}\sum_{i=1}^{n}\frac{A_{i}}{\hat{e}(X_{i})}\hat{\mu}_{1}(X_{i})\hat{S}_{1}(k\mid X_{i})\int_{0}^{\tilde{T}_{i}}\frac{\textnormal{d}M_{\hat{G}_{1}}(t)}{\hat{S}_{1}(t\mid X_{i})\hat{G}_{1}(t\mid X_{i})}$
$-n^{-1}\sum_{i=1}^{n}\frac{1-A_{i}}{1-\hat{e}(X_{i})}\hat{\mu}_{0}(X_{i})\hat{S}_{0}(k\mid X_{i})\int_{0}^{\tilde{T}_{i}}\frac{\textnormal{d}M_{\hat{G}_{0}}(t)}{\hat{S}_{0}(t\mid X_{i})\hat{G}_{0}(t\mid X_{i})}.$
More explicitly, taking the first modification term in $\hat{\tau}^{\textup{eif}}$ as an example, because the observed time points are discrete, the estimator of the integration part for observation $i$ is
$\int_{0}^{\tilde{T}_{i}}\frac{\textnormal{d}M_{\hat{G}_{1}}(t)}{\hat{S}_{1}(t\mid X_{i})\hat{G}_{1}(t\mid X_{i})}$ $=$ $\sum_{t\leq\tilde{T}_{i}}\frac{1(\Delta_{i}=0,C_{i}=t)-\hat{\lambda}_{1}(t\mid X_{i})}{\hat{S}_{1}(t\mid X_{i})\hat{G}_{1}(t\mid X_{i})}$ (8)
$=$ $-\sum_{t\leq\tilde{T}_{i}}\frac{\hat{\lambda}_{1}(t\mid X_{i})}{\hat{S}_{1}(t\mid X_{i})\hat{G}_{1}(t\mid X_{i})}+\frac{1(\Delta_{i}=0)}{\hat{S}_{1}(\tilde{T}_{i}\mid X_{i})\hat{G}_{1}(\tilde{T}_{i}\mid X_{i})},$
where $\hat{\lambda}_{1}(t\mid X_{i})$ denotes the estimator of the conditional hazard function of censoring $\lambda_{1}(t\mid X)$. The first term in (8) is a summation of the ratio $-\hat{\lambda}_{1}(t\mid X_{i})/\{\hat{S}_{1}(t\mid X_{i})\hat{G}_{1}(t\mid X_{i})\}$, where the summation is over all observed event time points before $\tilde{T}_{i}$. The second term is 0 for observations that are not right-censored by the treatment-unrelated ICE, and is $1/\{\hat{S}_{1}(\tilde{T}_{i}\mid X_{i})\hat{G}_{1}(\tilde{T}_{i}\mid X_{i})\}$ for observations with a treatment-unrelated ICE happened at time $\tilde{T}_{i}$.
To construct the estimator $\hat{\tau}^{\textup{eif}}$, we need to estimate the following nuisance parameters: the propensity score model $e(X)$, the outcome model $\mu_{a}(X)$ for $a=0,1$, the survival function $S_{a}(t\mid X)$, and the censoring mechanism $G_{a}(t\mid X)$ for $t\leq k$ and $a=0,1$. Nonetheless, the consistency of $\hat{\tau}^{\textup{eif}}$ does not require the correct specification of all four nuisance parameters, as shown in Theorem 3.
4.2 Asymptotic properties of the EIF-based estimator
We next discuss the asymptotic properties of $\hat{\tau}^{\textup{eif}}$. We first introduce additional notation. For $t\leq k$ and $a=0,1$, let $e^{*}$, $G_{a}^{*}$, $\mu_{a}^{*}$, and $S_{a}^{*}$ denote the probability limit of the estimated propensity score, estimated censoring mechanism, estimated outcome model, and estimated survival function of the AE time, respectively, i.e., $\lVert\hat{e}-e^{*}\rVert=o_{P}(1)$, $\lVert\hat{G}_{a}-G_{a}^{*}\rVert=o_{P}(1)$, $\lVert\hat{\mu}_{a}-\mu_{a}^{*}\rVert=o_{P}(1)$, and $\lVert\hat{S}_{a}-S_{a}^{*}\rVert=o_{P}(1)$. If the propensity score model is consistently estimated, we have $e^{*}=e$, and analogous results for the other three nuisance models.
In the following theorem, we provide the double robustness of the EIF estimator $\hat{\tau}^{\textup{eif}}$.
Theorem 3 (Double robustness of $\hat{\tau}^{\textup{eif}}$)
Under Assumptions 1 and 2, $\hat{\tau}^{\textup{eif}}$ is doubly robust in the sense that it is consistent for $\tau$ if either $\{e^{*}(X)=e(X),G_{a}^{*}(t\mid X)=G_{a}(t\mid X)\}$ or $\{\mu_{a}^{*}(X)=\mu_{a}(X),S_{a}^{*}(t\mid X)=S_{a}(t\mid X)\}$ for $t\leq k$ and $a=0,1$.
Theorem 3 shows that $\hat{\tau}^{\textup{eif}}$ is consistent if at least one of the following two sets of nuisance parameters are consistently estimated: (1) the outcome model $\mu_{a}(X)$ and the survival function $S_{a}(t\mid X)$; (2) the propensity score model $e(X)$ and the censoring mechanism $G_{a}(t\mid X)$, for $t\leq k$ and $a=0,1$. When our data is from an RCT, i.e., the propensity score $e(X)$ is known, and thus $\hat{e}(X)$ is always correctly specified, the EIF estimator $\hat{\tau}^{\textup{eif}}$ is a consistent estimator for $\tau$ if either the censoring mechanism $G_{a}(t\mid X)$ is consistently estimated, or both the outcome model $\mu_{a}(X)$ and the survival function $S_{a}(t\mid X)$ are consistently estimated, for $t\leq k$ and $a=0,1$.
To conduct statistical inference, we provide the asymptotic distribution of $\hat{\tau}^{\textup{eif}}$. We first introduce three technical conditions and then state the asymptotic result in a theorem.
Assumption 3 (Consistency of the nuisance parameters)
Assume that either $\{e^{*}(X)=e(X),\allowbreak G_{a}^{*}(t\mid X)=G_{a}(t\mid X)\}$ or $\{\mu_{a}^{*}(X)=\mu_{a}(X),S_{a}^{*}(t\mid X)=S_{a}(t\mid X)\}$ for $t\leq k$ and $a=0,1$ is satisfied.
Assumption 4 (Donsker Condition)
The class of functions $\{(e,G_{a},\mu_{a},S_{a}):||e-e^{*}||<\delta,||G_{a}-G_{a}^{*}||<\delta,||\mu_{a}-\mu_{a}^{*}||<\delta,||S_{a}-S_{a}^{*}||<\delta\}$ is Donsker for some $\delta>0$.
Assumption 5 (Convergence rates of nuisance parameters)
The convergence rate of the nuisance parameters estimation satisfy $\{||\hat{e}-e^{*}||+||\hat{G}_{a}-G_{a}^{*}||\}\{||\hat{\mu}_{a}-\mu_{a}^{*}||+||\hat{S}_{a}-S_{a}^{*}||\}=o_{P}(n^{-1/2})$ for $a=0,1$.
Assumption 3 requires either the propensity score and censoring models or the survival and outcome models to be consistently estimated. It guarantees the consistency of $\hat{\tau}^{\textup{eif}}$. Assumption 4 imposes restrictions on the nuisance model complexity and is a standard regularity condition (Van der Vaart,, 2000). Recent literature documents that we can use cross-fitting in the estimation of nuisance parameters to relax the Donsker condition (van der Laan et al.,, 2011; Chernozhukov et al.,, 2018; Díaz,, 2019). Assumption 5 imposes additional restrictions on the rate of convergence of the nuisance parameters, in addition to all of them being consistently estimated.
Theorem 4 (Asymptotic distribution)
Under Assumptions 1– 5, the EIF-based estimator satisfies
$n^{1/2}(\hat{\tau}^{\textup{eif}}-\tau)\ =\ n^{-1/2}\sum_{i=1}^{n}D_{\tau}(O_{i})+o_{P}(1).$
$\hat{\tau}^{\textup{eif}}$ is a consistent and asymptotically Gaussian estimator of $\tau$ with asymptotic variance equal to $E\{D_{\tau}^{2}(O)\}$, thus achieves the semiparametric efficiency bound.
Therefore, we can construct the variance estimator based on the semiparametric efficiency bound by taking the empirical analog of the plug-in estimation $n^{-1}\sum_{i=1}^{n}\hat{D}_{\tau}^{2}(O_{i})$. We can also use a nonparametric bootstrap to estimate variance and conduct statistical inference if the nuisance estimators satisfy certain smoothness conditions.
Compare $\hat{\tau}^{\textup{eif}}$ with the propensity score-augmented weighting estimator $\hat{\tau}^{\textup{aug}}$. On the one hand, $\hat{\tau}^{\textup{eif}}$ is more robust in the sense that it allows for misspecification of the censoring model and is asymptotically more efficient. Continuing the intuition discussed in Section 3.4, $\hat{\tau}^{\textup{aug}}$ only orthogonalizes the part corresponding to $A\mid X$ but not $C\mid X,A$, making it robust only to misspecification of the $A\mid X$ model, not the $C\mid X,A$ model. In contrast, $\hat{\tau}^{\textup{eif}}$ accounts for both possible sources of misspecification, leading to greater robustness and efficiency. On the other hand, $\hat{\tau}^{\textup{aug}}$ is more straightforward to implement in practice, as it only requires the estimation of $S_{a}(k\mid X)$ and $G_{a}(k\mid X)$ at time $k$ for $a=0,1$, while $\hat{\tau}^{\textup{eif}}$ requires consistent estimation of the entire survival curves $S_{a}(t\mid X)$ and $G_{a}(t\mid X)$ for $t\leq k$ and $a=0,1$. We consider the choice between $\hat{\tau}^{\textup{eif}}$ and $\hat{\tau}^{\textup{aug}}$ as a trade-off between implementation simplicity and robustness and efficiency of the estimators.
5 Simulation
In this section, we conduct Monte Carlo simulations to study the finite sample performance of our proposed estimators. We first generate the covariates $X=(X_{1},X_{2},X_{3})^{\tiny\textsc{t}}\in\mathbb{R}^{3}$ from three independent standard Gaussian distributions and denote $\tilde{X}_{j}=\{(X_{j}+2)^{2}-1\}/\sqrt{12}$ for $j=1,2,3$. Next, we generate the treatment assignment following $A\mid X\sim\textup{Bernoulli}(\{e(X)\})$ with $e(X)$ being the propensity score model, and generate the potential outcomes following $Y(a)\mid X\sim\mathcal{N}(\mu_{a}(X),\sigma_{a}^{2})$ with $\mu_{a}(X)$ being the outcome model for $a=0,1$. We then generate the potential values of the treatment-related ICE survival time $T(a)$ from the distribution with a survival function $S_{a}(t\mid X)=\exp[-0.002t^{1.2}\exp\{\gamma_{a}(X)\}]$ and the potential values of the treatment-unrelated ICE survival time $C(a)$ from the distributions with a survival function $G_{a}(t\mid X)=\exp[-\varrho_{a}(t)\exp\{\delta_{a}(X)\}]$, for $a=0,1$. For each of the nuisance parameters, we consider two scenarios when it will be correctly modeled and misspecified, therefore, we generate data following two different choices of each. We summarize the detailed data-generating choices in Table 2.
Table 2: Model choices for the nuisance parameters
Correctly specified Misspecified
$\textup{logit}\{e(X)\}$ $(X_{1}+X_{2}+X_{3})/5$ $1(X_{1}\geq 0)\{\exp(\tilde{X}_{2})-X_{2}(1+\tilde{X}_{3})\}-\exp(\tilde{X}_{2})$
$\mu_{1}(X)$ $2(X_{1}+X_{2}+X_{3})$ $1(X_{1}\geq 0)\{X_{2}+\exp(X_{2})\tilde{X}_{3}-\tilde{X}_{2}\}+\tilde{X}_{2}$
$\mu_{0}(X)$ $X_{1}+X_{2}+X_{3}$ $-\tilde{X}_{1}-1(X_{1}>0.5)\tilde{X}_{2}+1(X_{1}<-0.5)X_{2}^{2}\log(|X_{3}|+1)$
$\sigma_{a}$ $0.1(a+1)$ $1$
$\gamma_{1}(X)$ $0.1(X_{1}+2X_{2}-2X_{3})$ $0.1(X_{1}^{2}X_{2}-X_{2}-1)+1(X_{3}\neq 0)X_{2}\log(10X_{3}^{2})$
$\gamma_{0}(X)$ $0.1(X_{1}-2X_{2}+2X_{3})$ $0.01(-\tilde{X}_{1}+\tilde{X}_{2}+\tilde{X}_{3})$
$\delta_{1}(X)$ 00 $0.1(X_{1}^{2}X_{2}-X_{2}-1)+1(X_{3}\neq 0)X_{2}\log(10X_{3}^{2})$
$\delta_{0}(X)$ 00 00
$\varrho_{a}(t)$ $-0.01t^{1.2}$ $-0.01at^{1.2}+0.6*0.01^{1/1.2}(1-a)t$
We consider five data-generating regimes: all four models are correctly specified, misspecified $e$ and correct $(G,\mu,S)$, misspecified $(e,G)$ and correct $(\mu,S)$, misspecified $(\mu,S)$ and correct $(e,G)$, and all four models are misspecified. Let $Y=AY(1)+(1-A)Y(0)$, $T=AT(1)+(1-A)T(0)$, and $C=AC(1)+(1-A)C(0)$. Generate the observed event time as $T\wedge C\wedge k$, the observed event type indicator, and the observed outcome $Y$ if $T\wedge C>k$.
We compare the finite sample performance of the outcome estimator $\hat{\tau}^{\textup{out}}$, the weighting estimator $\hat{\tau}^{\textup{ipw}}$, the augmented weighting estimator $\hat{\tau}^{\textup{aug}}$, and the EIF estimator $\hat{\tau}^{\textup{eif}}$ based on $1000$ Monte Carlo samples with a sample size of $n=1000$. When fitting the nuisance models, we use logistic regression of $A$ on $X$ to estimate the propensity score model, the linear regression of $Y$ on $X$ on the subsample $A=a,T\wedge C>k$ to estimate the outcome model, and the Cox proportional hazard regression of $T$ and $C$ on $X$ on the subsample $A=a$ to estimate the treatment-related ICE survival model and the censoring model, respectively, for $a=0,1$. We use $X$ in all regression models, therefore, when the true data-generating processes involve non-linear functions of $X$, the fitted nuisance models suffer from misspecification. We evaluate the performance of the four estimators with reported bias, standard deviation, and coverage probability using a nonparametric bootstrap with $200$ bootstrap samples in Table 3.
Table 3: Finite sample performance of the four estimators. For each estimator, we report the finite-sample bias, standard deviation (SD), and the coverage rate (CR) of a 95% confidence interval, which is constructed using a nonparametric bootstrap with 200 bootstrap iterations. Each row corresponds to a different data-generating regime.
$\hat{\tau}^{\textup{out}}$ $\hat{\tau}^{\textup{ipw}}$ $\hat{\tau}^{\textup{aug}}$ $\hat{\tau}^{\textup{eif}}$
Bias SD CR Bias SD CR Bias SD CR Bias SD CR
all_correct 0.003 0.100 0.989 0.005 0.154 0.977 0.004 0.152 0.980 0.004 0.110 0.982
$e$_wrong $-0.002$ 0.134 0.978 $-0.338$ 0.527 0.916 $-0.003$ 0.479 0.961 0.000 0.234 0.979
$e\_G$_wrong $-0.002$ 0.130 0.985 $-0.194$ 1.591 0.831 0.141 1.559 0.972 $-0.013$ 0.354 0.969
$\mu\_S$_wrong $-0.085$ 0.170 0.936 $-0.009$ 0.186 0.974 $-0.009$ 0.186 0.974 $-0.012$ 0.176 0.972
all_wrong 0.230 0.233 0.837 0.660 5.799 0.898 0.850 5.782 0.967 0.573 3.022 0.919
Consistent with our theoretical results, all estimators have a finite sample bias close to zero when all models are correctly specified. When $e$ is misspecified and $(G,\mu,S)$ are correct, the weighting estimator $\hat{\tau}^{\textup{ipw}}$ is inconsistent while the other three estimators have small finite sample bias. Under the regime when $e$ and $G$ are misspecified, both $\hat{\tau}^{\textup{ipw}}$ and $\hat{\tau}^{\textup{aug}}$ perform poorly as they have larger finite sample biases, $\hat{\tau}^{\textup{out}}$ and $\hat{\tau}^{\textup{eif}}$ show small finite sample biases as expected. When $(\mu,S)$ are misspecified, the outcome estimator $\hat{\tau}^{\textup{out}}$ has a relatively large bias, while all other three estimators have near-zero biases. All estimators have non-negligible finite sample bias when all nuisance models are misspecified. $\hat{\tau}^{\textup{eif}}$ is most robust to model misspecifications across all different regimes.
$\hat{\tau}^{\textup{eif}}$ shows a relatively smaller standard deviation compared to $\hat{\tau}^{\textup{ipw}}$ and $\hat{\tau}^{\textup{aug}}$ in the regimes where all estimators are consistent. The outcome estimator $\hat{\tau}^{\textup{out}}$ always has the smallest standard deviation when $(\mu,S)$ is correctly specified. Both $\hat{\tau}^{\textup{ipw}}$ and $\hat{\tau}^{\textup{aug}}$ have larger standard deviations across all regimes in which they are consistent. Furthermore, confidence intervals constructed using the nonparametric bootstrap yield valid coverage rates for the corresponding consistent estimators in their respective regimes.
Next, we evaluate two ad-hoc methods commonly used in clinical trials with ICEs, demonstrating through simulations that they are biased in estimating the clinically relevant causal parameter. The first method is the non-responder imputation method that assigns an outcome as 00 whenever an ICE occurs. It is equivalent to using the composite outcome strategy that treats all ICEs as treatment-related, leading to a composite outcome equal to 00. The second method naively applies a hypothetical strategy to all ICEs, assuming they are independent of potential outcomes conditional on observed covariates. We generate data under the “$\mu\_S$_wrong regime, ensuring correct specification of both the propensity score $e(X)$ and the treatment-unrelated ICE survival model $G_{a}(t\mid X)$ for $a=0,1$. Based on $1000$ Monte Carlo simulations with a sample size of $n=1000$, the inverse probability-weighting estimator using the non-responder imputation method exhibits a finite sample bias of $-1.195$ with a standard deviation of $0.077$, and that using the hypothetical strategy for all ICEs has a bias of $1.188$ with a standard deviation of $0.388$. Both ad-hoc methods yield inconsistent estimators.
6 Application
In this section, we re-analyze data from two immunology trials to study the effect of baricitinib on Systemic Lupus Erythematosus (Morand et al.,, 2023; Petri et al.,, 2023). Both trials are double-blinded, randomized, placebo-controlled phase-3 trials. There are three treatment arms: 2mg baritinib, 4mg baritinib, and placebo. The primary endpoint of interest in both studies is the Systemic Lupus Erythematosus Responder Index 4 (SRI4) measured at week 52. It is a binary composite responder index based on both the level of improvement in disease activity and an indicator of no worsening of the overall condition or the development of substantial disease activity in new organ systems. The baseline covariates used in model-fitting include the region each patient is from, the baseline corticosteroid usage, and the baseline value of the Physician Global Assessment score measured at week 0.
There are significant numbers of observations with ICEs happening in both trials (218/760 and 211/775). We classify the observed intercurrent events into two types: those likely to be correlated with the potential outcomes such as adverse events (82.6% and 84.4%), and those conditionally independent of the potential outcomes given covariates and the treatment arm such as withdrawal due to relocation (17.4% and 15.6%). We compare both 2mg and 4mg treatment arms with the placebo group separately. We implemented all four estimators $\hat{\tau}^{\textup{out}}$, $\hat{\tau}^{\textup{ipw}}$, $\hat{\tau}^{\textup{aug}}$, and $\hat{\tau}^{\textup{eif}}$. We used logistic regression for estimating the propensity score and outcome models and used Cox proportional hazards regression to estimate the survival function and censoring model. We used the nonparametric bootstrap with 500 iterations to compute the variance estimator and conduct inference. We summarize the point estimators, their estimated standard errors, and the $p$-values in Table 4.
Table 4: Treatment effect of Baricitinib on the primary endpoint SRI4. The first four columns report the results of Trial 1 and the last four columns report those of Trial 2. The first three rows report the point estimators, standard errors, and the $p$-value of the composite outcome average treatment effect corresponding to the 2mg treatment arm, and the last three rows report those of the 4mg treatment arm. For each trial and each treatment arm, we report the results based on all four estimators.
Trial 1 (Petri et al.,, 2023) Trial 2 (Morand et al.,, 2023)
$\hat{\tau}^{\textup{out}}$ $\hat{\tau}^{\textup{ipw}}$ $\hat{\tau}^{\textup{aug}}$ $\hat{\tau}^{\textup{eif}}$ $\hat{\tau}^{\textup{out}}$ $\hat{\tau}^{\textup{ipw}}$ $\hat{\tau}^{\textup{aug}}$ $\hat{\tau}^{\textup{eif}}$
2mg point 0.030 0.029 0.026 0.026 0.019 0.019 0.022 0.022
se 0.042 0.043 0.042 0.043 0.042 0.043 0.043 0.042
$p$-value 0.479 0.504 0.534 0.540 0.643 0.662 0.602 0.606
4mg point 0.113 0.120 0.115 0.113 $-0.002$ $-0.002$ $-0.002$ $-0.002$
se 0.046 0.046 0.046 0.046 0.042 0.042 0.042 0.042
$p$-value 0.013 0.008 0.012 0.013 0.961 0.962 0.966 0.969
In both trials and both treatment arms, the point estimators are close to each other in most of the cases, and the statistical inference based on different estimators is remarkably robust across methods. Our results suggest a significantly positive effect of 4mg baricitinib treatment on SRI4 in trial 1, although that in trial 2 is insignificant with a slightly negative point estimator. Neither trial observes a significant effect of 2mg baricitinib treatment on SRI4.
Next, we use the two commonly used ad-hoc methods discussed at the end of Section 5 and compare them with our proposed methods. Table 5 reports the estimated treatment effect of both treatment arms across the two trials using the inverse probability-weighting estimator. The non-responder imputation method yields smaller estimates than our proposed methods most of the time, although the difference is modest due to the relatively low proportion of treatment-unrelated ICEs in both trials. Nonetheless, our proposed method targets a clinically more meaningful causal parameter. The second ad-hoc method naively applies the hypothetical strategy to all ICEs, assuming they are independent of potential outcomes conditional on the observed covariates. Such independence assumption is untestable and likely to be violated in clinical trials. For instance, patients discontinued treatment due to adverse events or lack of efficacy. As shown in Table 5, this method produces estimated treatment effects substantially different from the obtained results using our proposed methods, leading to misleading clinical conclusions. It aligns with the large sample bias it exhibits in the simulation study.
Table 5: Treatment effect of Baricitinib on the primary endpoint SRI4 using non-responder imputation (NRI) and hypothetical strategy (HS). The first two columns report the results of Trial 1 and the last two columns report those of Trial 2. The first three rows report the point estimators, standard errors, and the $p$-value of the estimated treatment effect corresponding to the 2mg treatment arm, and the last three rows report those of the 4mg treatment arm. For each trial and each treatment arm, we report the results based on the inverse probability-weighting estimator.
Trial 1 Trial 2
NRI HS NRI HS
2mg point 0.039 0.011 0.010 0.022
se 0.043 0.049 0.042 0.050
$p$-value 0.362 0.815 0.817 0.654
4mg point 0.093 0.073 $-0.008$ 0.044
se 0.044 0.050 0.041 0.051
$p$-value 0.036 0.147 0.847 0.382
7 Discussion
In this study, we address the challenges posed by ICEs in RCTs by proposing methods to handle competing ICEs. We classify ICEs into treatment-related and treatment-unrelated events and apply different strategies for the estimation of a clinically meaningful causal effect. For treatment-related ICEs, we use the composite outcome approach, assigning values that capture treatment failure. For treatment-unrelated ICEs, we adopt a hypothetical strategy that assumes their occurrence is conditionally independent of the potential outcome given covariates. Our key insight is that patients discontinue treatment following the first ICE that occurs, which in turn censors the potential occurrence of ICEs of another type. In this paper, we rigorously address the issue of competing ICEs by employing a formal causal inference framework. We provide identification formulas for the causal parameter of interest, derive its semiparametric efficient influence function, and propose multiple estimators. While our proposed framework provides a rigorous and flexible approach to handling competing ICEs, several challenges and extensions remain open for future research.
7.1 Data collection and ICE classification
Our proposed methods have broad applicability across various therapeutic areas, including immunology, oncology, and cardiology, where treatment discontinuation and other ICEs frequently occur. However, the effectiveness of these approaches relies on the accurate classification of ICEs, which should be performed in collaboration with clinicians to ensure clinical relevance in trial analyses. In our application studies, the two types of ICEs were classified based on a manual review of the detailed comments collected at the clinic. A modernized case report form for a more accurate collection of reasons for treatment discontinuation is required for each application of the proposed method in clinical trials. Fortunately, a cross-industry PHUSE working group is tackling this problem, and a recommended new case report form will be available soon (PHUSE Working Group,, 2024).
7.2 Random $K$
In practice, the measurement of the endpoint of interest $Y$ does not necessarily happen at a fixed point $k$. For instance, the time $K$ that a patient visits the clinic and takes the measurement can be treated as a random variable independent of treatment $A$, potential outcomes $Y(a)$, and both types of ICEs $T(a)$ and $C(a)$ for $a=0,1$. For a random $K$, the analogous formulas to both the previous identification formulas (2) and (3) hold, with a replacement of $k$ to $K$, i.e.,
$\tau$ $=$ $E\left\{\mu_{1}(X)S_{1}(K\mid X)-\mu_{0}(X)S_{0}(K\mid X)\right\}$ (9)
$=$ $E\left[\frac{AY1(T\wedge C>K)}{e(X)G_{1}(K\mid X)}-\frac{(1-A)Y1(T\wedge C>K)}{\left\{1-e(X)\right\}G_{0}(K\mid X)}\right].$ (10)
If $K$ is a pre-treatment covariate that we observe, the previous analyses carry over. If $K$ is also treated as a post-treatment variable, we can construct a weighting estimator following (10). However, (9) ceases to serve as an identification formula unless $S_{a}(K\mid X)$ is identified for $a=0,1$. Considering the presence of three competing time-to-event random variables, it becomes necessary to employ competing risks models. We defer this analysis to future work.
7.3 Comparison between composite variable and principal stratification
We propose to use a combination of the composite variable and hypothetical strategies to deal with ICEs in this project. Following ICH E9 , 2019 (R1), we can also use the combination of principal stratification and hypothetical strategies to deal with the treatment-related and treatment-unrelated ICEs, respectively. Using the principal stratification strategy will target different causal parameters and rely on additional identification assumptions such as monotonicity and principal ignorability of the two potential values of the AE time (Ding and Lu,, 2017; Jiang and Ding,, 2021; Jiang et al.,, 2022). We leave this for future work.
Supplementary Material
We provide proof of all theorems and propositions in the supplementary material.
Appendix A Proofs
A.1 Proof of Theorem 1
First, equation (2) holds because
LHS $=$ $E[E\{Y(1)1(T(1)>k)\mid X\}]$
$=$ $E[E\{Y1(T>k)\mid X,A=1\}]$
$=$ $E\{E(Y\mid T>k,X,A=1)\textup{pr}(T>k\mid X,A=1)\}$
$=$ $E\{E(Y\mid T>k,C>k,X,A=1)\textup{pr}(T>k\mid X,A=1\}=\textup{RHS},$
where the first equality is by the law of iterated expectations, the second equality is by the randomization assumption 1, and the fourth equality is by the censoring at random assumption 2.
Next, equation (3) holds because
RHS $=$ $E\left[E\left\{\frac{AY1(T\wedge C>k)}{e(X)\textup{pr}(C>k\mid X,A=1)}\mid X\right\}\right]$
$=$ $E\left(\frac{1}{e(X)\textup{pr}(C>k\mid X,A=1)}E\left[AY(1)1\{T(1)\wedge C(1)>k\}\mid X\right]\right)$
$=$ $E\left(\frac{1}{e(X)\textup{pr}(C>k\mid X,A=1)}E(A\mid X)E\left[Y(1)1\{T(1)>k\}1\{C(1)>k\}\mid X\right]\right)$
$=$ $E\left(\frac{1}{\textup{pr}(C>k\mid X,A=1)}E[1\{C(1)>k\}\mid X]E\left[Y(1)1\{T(1)>k\}\mid X\right]\right)$
$=$ $E[E\{Y(1)1(T(1)>k)\mid X\}]=\textup{LHS},$
where the first equality is by the law of iterated expectations, the third equality is by the randomization assumption 1, the fourth equality is by the censoring at random assumption 2, and the last equality is again by the law of iterated expectations. $\square$
A.2 Proof of Proposition 1
We prove that when the censoring model is correctly specified, $\hat{\tau}^{\textup{aug}}$ is doubly robust in the sense that it is consistent to $\tau$ if either $e^{*}=e$ or $(\mu_{1}^{*}=\mu_{1},S_{1}^{*}=S_{1})$. Want to show
$E\left\{\frac{AY1(T\wedge C>k)}{e^{*}(X)G_{1}(k\mid X)}-\frac{A-e^{*}(X)}{e^{*}(X)}\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)-\mu_{1}\right\}$ $=$ 00 0
if either $e^{*}=e$ or $(\mu_{1}^{*}=\mu_{1},S_{1}^{*}=S_{1})$ holds.
$E\left\{\frac{AY1(T\wedge C>k)}{e^{*}(X)G_{1}(k\mid X)}-\frac{A-e^{*}(X)}{e^{*}(X)}\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)\right\}$
$=$ $E\left[\frac{E\{AY1(T\wedge C>k)\mid X\}}{e^{*}(X)G_{1}(k\mid X)}-\frac{e(X)-e^{*}(X)}{e^{*}(X)}\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)\right]$
$=$ $E\left\{\frac{e(X)}{e^{*}(X)}\mu_{1}(X)S_{1}(k\mid X)-\frac{e(X)-e^{*}(X)}{e^{*}(X)}\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)\right\},$
where the first equality is by the law of integrated expectations and the second equality follows from similar derivations as in the proof of Theorem 1. The last line is equal to $\mu_{1}$ if either $e^{*}=e$ or $(\mu_{1}^{*}=\mu_{1},S_{1}^{*}=S_{1})$. $\square$
A.3 Proof of Theorem 2
We follow the semiparametric theory in Bickel et al., (1993) to derive the EIF for $\mu_{1}$. As discussed in Section 4, we have two levels of coarsening of the full data $(X,Y^{\textup{cp}}(1),Y^{\textup{cp}}(0))$, one is because the treatment-unrelated ICE time $C(a)$ is censoring $T(a)\wedge k$, and the other is due to the treatment assignment. Following the steps in Hubbard et al., (2000), we first consider the case when every observation is assigned to the treatment group, i.e., when there is no missingness generated by the treatment assignment, and then project the derived EIF onto the nuisance tangent space of the propensity score to get the final form of EIF for $\mu_{1}$.
With full data $(X,Y^{\textup{cp}}(1),Y^{\textup{cp}}(0))$, the estimating equation for $\mu_{1}$ is $D_{\textup{full}}=Y^{\textup{cp}}(1)-\mu_{1}$. Since $Y^{\textup{cp}}(1)$ is only observable if $C(1)>T(1)\wedge k$, the full data is not available even if all corresponding potential outcomes under treatment $A=1$ are observed, and the observed data is
$O(1)$ $=$ $(X,\Delta(1),\tilde{T}(1),\Delta(1)Y^{\textup{cp}}(1))\sim P_{1},$
where $\Delta(1)=1(C(1)>T(1)\wedge k)$ is the missing indicator with $\Delta(1)=1$ if $Y^{\textup{cp}}(1)$ is observed and $\Delta(1)=0$ otherwise, and $\tilde{T}(1)=T(1)\wedge C(1)\wedge k$ is the observed event time with $\tilde{T}(1)=T(1)\wedge k$ if $\Delta(1)=1$, $Y^{\textup{cp}}(1)$ is observed, and $\tilde{T}(1)=C(1)$ if $\Delta(1)=0$, $Y^{\textup{cp}}(1)$ is missing. An identification formula for $\mu_{1}$ is
$\mu_{1}=E\left\{\frac{\Delta(1)Y^{\textup{cp}}(1)}{G_{1}(\tilde{T}(1)\mid X)}\right\},$
where $G_{1}(t\mid X)=\textup{pr}(C(1)>t\mid X)$ is the probability of not censoring up to time $t$ conditioning on the covariates. We can write the IPCW estimating equation as
$D_{\textup{IPCW}}=\frac{\Delta(1)\{Y^{\textup{cp}}(1)-\mu_{1}\}}{G_{1}(\tilde{T}(1)\mid X)},$
and have $E(D_{\textup{IPCW}})=0$ by the law of iterated expectations and the censoring at random assumption.
Next, we follow the steps in van der Laan and Rubin, (2007) to derive the EIF for $\mu_{1}$ when the observed data is $O(1)=(X,\Delta(1),\tilde{T}(1),\Delta(1)Y^{\textup{cp}}(1))$. For ease of notation, we omit the subscripts, superscripts, or numbers in parenthesis $(1)$ that indicate the potential outcomes under the treatment $A=1$ in the following derivation, so the dependence on the treatment arm is implicit. For ease of notation, denote $T^{1}=T\wedge k$.
Let $T(P)$ denote the tangent space which is the whole Hilbert space $L_{0}^{2}(P)$ since our model is nonparametric. The tangent space can be written as a direct sum of three components $T=T_{X}\oplus T_{\textup{F}}\oplus T_{\textup{CAR}}$ with
$T_{X}$ $=$ $\{h(X)\in L_{0}^{2}(P):E\{h(X)\}=0\},$
$T_{\textup{F}}$ $=$ $\{h(O)\in L_{0}^{2}(P):E\{h(O)\mid C,X\}=0\},$
$T_{\textup{CAR}}$ $=$ $\{h(O)\in L_{0}^{2}(P):E\{h(O)\mid Y^{\textup{cp}},T^{1},X\}=0\},$
where $T_{X}$, $T_{\textup{F}}$, and $T_{\textup{CAR}}$ are orthogonal to each other due to the censoring at random assumption and the factorization of the observed data. These tangent spaces are generated from scores of submodels that perturb the marginal distribution of $X$, the conditional distribution of $Y^{1}\mid X$, and the conditional censoring probability $C\mid X$, respectively. Due to the orthogonality, we have the decomposition of any $h(O)\in L_{0}^{2}(P)$ as
$h(O)$ $=$ $\Pi\{h\mid T(P)\}$
$=$ $\Pi\{h\mid T_{X}(P)\}+\Pi\{h\mid T_{\textup{F}}(P)\}+\Pi\{h\mid T_{\textup{CAR}}(P)\}.$
Let $D(O)$ denote the efficient influence function for $\mu_{1}$. By the results in Chapter 1.4 of van der Laan and Robins, (2003), we have: (1) $D(O)$ should be orthogonal to the nuisance tangent space $T_{\textup{CAR}}$, thus $\Pi\{D(O)\mid T_{\textup{CAR}}(P)\}=0$; (2) $D(O)$ can be rewritten as
$D=D_{\textup{IPCW}}-\Pi\{D_{\textup{IPCW}}\mid T_{\textup{CAR}}(P)\},$
thus
$\Pi\{D\mid T_{\textup{F}}(P)\}$ $=$ $\Pi\{D_{\textup{IPCW}}\mid T_{\textup{F}}(P)\}-\Pi\{\Pi\{D_{\textup{IPCW}}\mid T_{\textup{CAR}}(P)\}\mid T_{\textup{F}}(P)\}$
$=$ $\Pi\{D_{\textup{IPCW}}\mid T_{\textup{F}}(P)\}$
where the last equality is by the fact that $T_{\textup{F}}$ and $T_{\textup{CAR}}$ are orthogonal to each other, and similarly, $\Pi\{D\mid T_{X}(P)\}=\Pi\{D_{\textup{IPCW}}\mid T_{X}(P)\}$.
These projections provide us with two different ways to compute the EIF $D$: (1) directly compute the projections $\Pi\{D_{\textup{IPCW}}\mid T_{X}(P)\}$ and $\Pi\{D_{\textup{IPCW}}\mid T_{\textup{F}}(P)\}$ and sum them up; (2) compute the projection $\Pi\{D_{\textup{IPCW}}\mid T_{\textup{CAR}}(P)\}$ and subtract it from $D_{\textup{IPCW}}$. In the classic survival outcome problem, these two methods are symmetric since $T$ and $C$ are censoring each other thus the projections on $T_{\textup{F}}(P)$ and $T_{\textup{CAR}}(P)$ are very similar. However, due to the complication generated by $Y$ in our setting, the second approach is easier, since the tangent space $T_{\textup{F}}(P)$ is hard to compute.
We next compute $\Pi\{D_{\textup{IPCW}}\mid T_{\textup{CAR}}(P)\}$. By Theorem 1.1 in van der Laan and Robins, (2003),
$T_{\textup{CAR}}(P)=\overline{\left\{\int H(t,\mathcal{F}(t))\textnormal{d}M_{G}(t)\textup{ for all functions }H(t,\mathcal{F}(t))\right\}\cap L_{0}^{2}(P)},$
and the projection of a function $h(O)$ onto $T_{\textup{CAR}}(P)$ is
$\Pi\{h(O)\mid T_{\textup{CAR}}(P)\}=\int_{0}^{\tilde{T}}\left\{E(h(O)\mid dA(t)=1,\mathcal{F}(t))-E(h(O)\mid dA(t)=0,\mathcal{F}(t))\right\}\textnormal{d}M_{G}(t),$
where $A(t)=1(C\leq t)$ is the indicator of whether censoring happens up until time $t$ (define $C=\infty$ if $C>T^{1}$ so that $C$ is always observed), $\mathcal{F}(t)=(\bar{A}(t-),X)$ is the history observed up to time $t$, and $\textnormal{d}M_{G}(t)=1(C\in dt,\Delta=0)-1(\tilde{T}\geq t)d\Lambda(t\mid X)$ is the Doob-Meyer martingale of the counting process of censoring $C$. $E(D_{\textup{IPCW}}\mid dA(t)=1,\mathcal{F}(t))=0$ by the definition of $D_{\textup{IPCW}}$, thus we only need to compute $E(D_{\textup{IPCW}}\mid dA(t)=0,\mathcal{F}(t))$, which plus $\mu_{1}$ is equal to
$E\left\{\frac{\Delta Y^{\textup{cp}}}{G(\tilde{T}\mid X)}\mid dA(t)=0,\mathcal{F}(t)\right\}$
$=$ $E\left\{\frac{\Delta Y^{\textup{cp}}}{G(\tilde{T}\mid X)}\mid T^{1}\geq t,C\geq t,X\right\}$
$=$ $E\left\{\frac{\Delta Y^{\textup{cp}}}{G(\tilde{T}\mid X)}\mid T^{1}\geq t,X\right\}\big{/}G(t\mid X)$
$=$ $E\left\{\frac{Y^{\textup{cp}}}{G(T\wedge k\mid X)}\textup{pr}(C>T\wedge k\mid T^{1}\geq t,T,Y^{1},X)\mid T^{1}\geq t,X\right\}\big{/}G(t\mid X)$
$=$ $E\left\{\frac{Y^{\textup{cp}}}{G(T\wedge k\mid X)}\textup{pr}(C>T\wedge k\mid X)\mid T^{1}\geq t,X\right\}\big{/}G(t\mid X)$
$=$ $E\left\{Y^{\textup{cp}}\mid T^{1}\geq t,X\right\}\big{/}G(t\mid X).$
By the fact that the integration is over $t:t<\tilde{T}$, we have $t<T\wedge k\wedge C$ thus $t>k$. The conditional expectation
$E\left\{Y^{\textup{cp}}\mid T^{1}\geq t,X\right\}$ $=$ $E\left\{Y^{\textup{cp}}\mid T^{1}\geq t,T^{1}=T,X\right\}\textup{pr}(T^{1}=T\mid T^{1}\geq t,X)$
$+E\left\{Y^{\textup{cp}}\mid T^{1}\geq t,T^{1}=k,X\right\}\textup{pr}(T^{1}=k\mid T^{1}\geq t,X)$
$=$ $E(Y\mid T>k\geq t,X)\textup{pr}(T>k\mid T^{1}\geq t,X)$
$=$ $E(Y\mid T>k,X)\textup{pr}(T>k\mid T\wedge k\geq t,X)$
$=$ $E(Y\mid T>k,X)\textup{pr}(T>k\mid T\geq t,X)$
$=$ $\frac{E(Y1(T>k)\mid X)}{\textup{pr}(T>k\mid X)}\frac{\textup{pr}(T>k\mid X)}{\textup{pr}(T\geq t\mid X)}$
$=$ $\frac{E(Y1(T>k)\mid X)}{\textup{pr}(T\geq t\mid X)}.$
Therefore, we have
$\Pi\{D_{\textup{IPCW}}\mid T_{\textup{CAR}}(P)\}+\mu_{1}$ $=$ $-\int_{0}^{\tilde{T}}\frac{E(Y1(T>k)\mid X)}{\textup{pr}(T\geq t\mid X)}\frac{\textnormal{d}M_{G}(t)}{G(t\mid X)}$
$=$ $-E(Y1(T>k)\mid X)\int_{0}^{\tilde{T}}\frac{\textnormal{d}M_{G}(t)}{S(t\mid X)G(t\mid X)},$
and thus the EIF assuming $O(1)$ is the observed data is
$D$ $=$ $\frac{\Delta Y^{\textup{cp}}}{G(\tilde{T}\mid X)}+E(Y^{\textup{cp}}\mid X)\int_{0}^{\tilde{T}}\frac{\textnormal{d}M_{G}(t)}{S(t\mid X)G(t\mid X)}-\mu_{1}$ (S1)
Next, we follow steps in Section 3 of Hubbard et al., (2000) and compute the EIF when the real observed data is $O=(X,A,\Delta,\tilde{T},\Delta Y^{\textup{cp}})$. We need to construct a weighting estimating equation and then subtract its projection onto the nuisance tangent space of the propensity score to get the final form of the EIF. To be clear on the distinction between potential outcomes and the observed values, we add back the dependence on the treatment assignment in (S1) and write it as
$D(1)$ $=$ $\frac{\Delta(1)Y^{\textup{cp}}(1)}{G_{1}(\tilde{T}(1)\mid X)}+E(Y^{\textup{cp}}(1)\mid X)\int_{0}^{\tilde{T}(1)}\frac{\textnormal{d}M_{G_{1}}(t)}{S_{1}(t\mid X)G_{1}(t\mid X)}-\mu_{1}.$
A valid weighting estimating equation is
$D_{\textup{IPW}}$ $=$ $\frac{A}{e(X)}\left\{\frac{\Delta Y^{\textup{cp}}}{G_{1}(\tilde{T}\mid X)}+E(Y^{\textup{cp}}\mid X,A=1)\int_{0}^{\tilde{T}}\frac{\textnormal{d}M_{G_{1}}(t)}{S_{1}(t\mid X)G_{1}(t\mid X)}\right\}-\mu_{1}.$
Further, project this onto the nuisance tangent space of the propensity score, the projection is
$\Pi\{D_{\textup{IPW}}\mid T_{\textup{pscore}}\}$ $=$ $E\left\{D_{\textup{IPW}}\mid A,X\right\}-E\left\{D_{\textup{IPW}}\mid X\right\}$
$=$ $\frac{A-e(X)}{e(X)}E\left\{\frac{\Delta Y^{\textup{cp}}}{G_{1}(\tilde{T}\mid X)}\mid X,A=1\right\}$
$=$ $\frac{A-e(X)}{e(X)}E(Y^{\textup{cp}}\mid X,A=1),$
where the second equality is by the fact that $E\{\textnormal{d}M_{G_{1}}(t)\mid X,A=1\}=E\{\textnormal{d}M_{G_{1}}(t)\mid X\}=0$, and the last equality is by the censoring at random assumption. Therefore, the EIF
$D_{1}$ $=$ $D_{\textup{IPW}}-\Pi\{D_{\textup{IPW}}\mid T_{\textup{pscore}}\}$
$=$ $\frac{A}{e(X)}\left\{\frac{\Delta Y^{\textup{cp}}}{G_{1}(\tilde{T}\mid X)}+E(Y1(T>k)\mid X,A=1)\int_{0}^{\tilde{T}}\frac{\textnormal{d}M_{G_{1}}(t)}{S_{1}(t\mid X)G_{1}(t\mid X)}\right\}$
$-\frac{A-e(X)}{e(X)}E(Y1(T>k)\mid X,A=1)-\mu_{1}$
has the given form in Theorem 2. $\square$
A.4 Proof of Theorem 3
We prove the result by showing
$\mu_{1}=E\left[\frac{A}{e^{*}(X)}\left\{\frac{Y1(T\wedge C>k)}{G_{1}^{*}(k\mid X)}+\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)\int_{0}^{\tilde{T}}\frac{\textnormal{d}M_{G_{1}^{*}}(t)}{S_{1}^{*}(t\mid X)G_{1}^{*}(t\mid X)}\right\}-\frac{A-e^{*}(X)}{e^{*}(X)}\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)\right]$ (S2)
if either $\{e^{*}(X)=e(X),G_{1}^{*}(t\mid X)=G_{1}(t\mid X)\}$ or $\{\mu_{1}^{*}(X)=\mu_{1}(X),S_{1}^{*}(t\mid X)=S_{1}(t\mid X)\}$ for $t\leq k$.
Denote
$\mathcal{T}_{1}$ $=$ $\frac{A}{e^{*}(X)}\frac{Y1(T\wedge C>k)}{G_{1}^{*}(k\mid X)}-\frac{A-e^{*}(X)}{e^{*}(X)}\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)-\mu_{1}(X)S_{1}(k\mid X),$
$\mathcal{T}_{2}$ $=$ $\frac{A}{e^{*}(X)}\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)\int_{0}^{\tilde{T}}\frac{\textnormal{d}M_{G_{1}^{*}}(t)}{S_{1}^{*}(t\mid X)G_{1}^{*}(t\mid X)}.$
By the identification formula (2), to prove (S2), it suffices to show $E(\mathcal{T}_{1}+\mathcal{T}_{2})=0$ if either $\{e^{*}(X)=e(X),G_{1}^{*}(t\mid X)=G_{1}(t\mid X)\}$ or $\{\mu_{1}^{*}(X)=\mu_{1}(X),S_{1}^{*}(t\mid X)=S_{1}(t\mid X)\}$ for $t\leq k$.
First, we have
$E\{\textnormal{d}M_{G_{1}^{*}}(t)\mid X\}$ $=$ $E\{1(C\in\textnormal{d}t,\Delta=0)\mid X\}-E\{1(\tilde{T}\geq t)\textnormal{d}\Lambda_{1}^{*}(t\mid X)\mid X\}$
$=$ $S_{1}(t\mid X)G_{1}(t\mid X)\textnormal{d}\Lambda_{1}(t\mid X)-S_{1}(t\mid X)G_{1}(t\mid X)\textnormal{d}\Lambda_{1}^{*}(t\mid X)$
$=$ $S_{1}(t\mid X)G_{1}(t\mid X)\{\textnormal{d}\Lambda_{1}(t\mid X)-\textnormal{d}\Lambda_{1}^{*}(t\mid X)\},$
therefore,
$E(\mathcal{T}_{2}\mid X)$ $=$ $\frac{e(X)}{e^{*}(X)}\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)E\left\{\int_{0}^{k}\frac{\textnormal{d}M_{G_{1}^{*}}(t)}{S_{1}^{*}(t\mid X)G_{1}^{*}(t\mid X)}\mid X\right\}$
$=$ $\frac{e(X)}{e^{*}(X)}\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)\int_{0}^{k}\frac{S_{1}(t\mid X)G_{1}(t\mid X)}{S_{1}^{*}(t\mid X)G_{1}^{*}(t\mid X)}\{\textnormal{d}\Lambda_{1}(t\mid X)-\textnormal{d}\Lambda_{1}^{*}(t\mid X)\},$
where the first equality follows from Assumption 1. Next, we have
$E(\mathcal{T}_{1}\mid X)$ $=$ $\frac{e(X)G_{1}(k\mid X)}{e^{*}(X)G_{1}^{*}(k\mid X)}\mu_{1}(X)S_{1}(k\mid X)-\frac{e(X)-e^{*}(X)}{e^{*}(X)}\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)-\mu_{1}(X)S_{1}(k\mid X)$
$=$ $\frac{e(X)}{e^{*}(X)}\left\{\frac{G_{1}(k\mid X)}{G_{1}^{*}(k\mid X)}-1\right\}\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)$
+$+\left\{\frac{e(X)G_{1}(k\mid X)}{e^{*}(X)G_{1}^{*}(k\mid X)}-1\right\}\left\{\mu_{1}(X)S_{1}(k\mid X)-\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)\right\}$
$=$ $-\frac{e(X)}{e^{*}(X)}\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)\int_{0}^{k}\frac{G_{1}(t\mid X)}{G_{1}^{*}(t\mid X)}\{\textnormal{d}\Lambda_{1}(t\mid X)-\textnormal{d}\Lambda_{1}^{*}(t\mid X)\}$
+$+\left\{\frac{e(X)G_{1}(k\mid X)}{e^{*}(X)G_{1}^{*}(k\mid X)}-1\right\}\left\{\mu_{1}(X)S_{1}(k\mid X)-\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)\right\},$
where the last equality follows from the Duhamel equation (Gill and Johansen,, 1990; Westling et al.,, 2024). Combing $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$, we have
$E(\mathcal{T}_{1}+\mathcal{T}_{2})$ $=$ $E\{E(\mathcal{T}_{1}+\mathcal{T}_{2}\mid X)\}$
$=$ $\frac{e(X)}{e^{*}(X)}\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)\int_{0}^{k}\left\{\frac{S_{1}(t\mid X)}{S_{1}^{*}(t\mid X)}-1\right\}\frac{G_{1}(t\mid X)}{G_{1}^{*}(t\mid X)}\{\textnormal{d}\Lambda_{1}(t\mid X)-\textnormal{d}\Lambda_{1}^{*}(t\mid X)\}$
+$+\left\{\frac{e(X)G_{1}(k\mid X)}{e^{*}(X)G_{1}^{*}(k\mid X)}-1\right\}\left\{\mu_{1}(X)S_{1}(k\mid X)-\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)\right\},$
which is equal to 0 if either $\{e^{*}(X)=e(X),G_{1}^{*}(t\mid X)=G_{1}(t\mid X)\}$ or $\{\mu_{1}^{*}(X)=\mu_{1}(X),S_{1}^{*}(t\mid X)=S_{1}(t\mid X)\}$ for $t\leq k$.
By symmetry, the part corresponding to the control arm $a=0$ also holds, thus the double robustness in Theorem 3 holds. $\square$
A.5 Proof of Theorem 4
Following the von Mises expansion (Hines et al.,, 2022), let $P$ and $P_{n}$ denote the true and empirical distribution of the observed data, respectively, and $\hat{P}_{n}$ denotes an estimated $P$, we have
$n^{1/2}(\hat{\tau}^{\textup{eif}}-\tau)\ =\ n^{-1/2}\sum_{i=1}^{n}D_{\tau}(O_{i})+n^{1/2}(P_{n}-P)\{D_{\tau}(O,\hat{P}_{n})-D_{\tau}(O)\}-n^{1/2}R(P,\hat{P}_{n}),$
where $D_{\tau}(O,\hat{P}_{n})$ is the EIF for $\tau$ when plugging in the estimated values of the nuisance parameters and $R(P,\hat{P}_{n})$ denotes the higher order remainder term. Under Assumption 4, the empirical process term $n^{1/2}(P_{n}-P)\{D_{\tau}(O,\hat{P}_{n})-D_{\tau}(O)\}=o_{P}(1)$ (Van der Vaart,, 2000). The remainder term corresponding to the treatment arm $a=1$ satisfies
$R_{1}(P,\hat{P}_{n})$ $=$ $-E\{D_{\tau}(O,\hat{P}_{n})\}-E\{\hat{\mu}_{1}(X)\hat{S}_{1}(k\mid X)-\mu_{1}(X)S_{1}(k\mid X)\}$
$=$ $-E\left[\frac{A}{\hat{e}(X)}\left\{\frac{Y1(T\wedge C>k)}{\hat{G}_{1}(k\mid X)}+\hat{\mu}_{1}(X)\hat{S}_{1}(k\mid X)\int_{0}^{k}\frac{\textnormal{d}M_{\hat{G}_{1}}(t)}{\hat{G}_{1}(t\mid X)\hat{S}_{1}(t\mid X)}\right\}\right.$
$\quad\quad\ \left.-\frac{A-\hat{e}(X)}{\hat{e}(X)}\hat{\mu}_{1}(X)\hat{S}_{1}(k\mid X)-\hat{\mu}_{1}(X)\hat{S}_{1}(k\mid X)+\hat{\mu}_{1}(X)\hat{S}_{1}(k\mid X)-\mu_{1}(X)S_{1}(k\mid X)\right]$
$=$ $-E\left[\frac{e(X)}{\hat{e}(X)}\hat{\mu}_{1}(X)\hat{S}_{1}(k\mid X)\int_{0}^{k}\left\{\frac{S_{1}(t\mid X)}{\hat{S}_{1}(t\mid X)}-1\right\}\frac{G_{1}(t\mid X)}{\hat{G}_{1}(t\mid X)}\{\textnormal{d}\Lambda_{1}(t\mid X)-\textnormal{d}\hat{\Lambda}_{1}(t\mid X)\}\right]$
$+E\left[\left\{\frac{e(X)G_{1}(k\mid X)}{\hat{e}(X)\hat{G}_{1}(k\mid X)}-1\right\}\left\{\mu_{1}(X)S_{1}(k\mid X)-\hat{\mu}_{1}(X)\hat{S}_{1}(k\mid X)\right\}\right]$
$=$ $E\left\{\frac{e(X)}{\hat{e}(X)}\hat{\mu}_{1}(X)\hat{S}_{1}(k\mid X)\int_{0}^{k}\frac{S_{1}(t\mid X)-\hat{S}_{1}(t\mid X)}{\hat{S}_{1}(t\mid X)}\frac{G_{1}(t\mid X)-\hat{G}_{1}(t\mid X)}{\hat{G}_{1}(t\mid X)}\textnormal{d}t\right\}$
$+E\left[\frac{e(X)G_{1}(k\mid X)-\hat{e}(X)\hat{G}_{1}(k\mid X)}{\hat{e}(X)\hat{G}_{1}(k\mid X)}\left\{\mu_{1}(X)S_{1}(k\mid X)-\hat{\mu}_{1}(X)\hat{S}_{1}(k\mid X)\right\}\right],$
where the third equality follows from a similar derivation as in the Proof of Theorem 3 and the fourth equality follows from the Duhamel equation (Gill and Johansen,, 1990; Westling et al.,, 2024). Let $R_{11}$ and $R_{12}$ denote the two terms in the last two lines, respectively. For the remainder term to be of small order, we need $R_{11}=o_{P}(n^{-1/2})$ and $R_{12}=o_{P}(n^{-1/2})$.
Next, we show that Assumption 5 is a sufficient condition for $R_{11}=o_{P}(n^{-1/2})$ and $R_{12}=o_{P}(n^{-1/2})$. In the following discussion, we suppress the dependency of the nuisance functions and their estimated values on $X$ for ease of notation. By the facts that $e\hat{\mu}_{1}\hat{S}_{1}(k)/\hat{e}$ is bounded and $1/\{\hat{S}_{1}(t)\hat{G}_{1}(t)\}$ is bounded for any $t\leq k$, we employ the Cauchy-Schwarz inequality to upper bound the term $R_{11}$ by a constant times $\int_{0}^{k}||S_{1}(t)-\hat{S}_{1}(t)||_{2}||G_{1}(t)-\hat{G}_{1}(t)||_{2}\textnormal{d}t$, where $||\cdot||_{2}$ denotes the $L_{2}(P)$ norm. It follows that Assumption 5 guarantees $R_{11}=o_{P}(n^{-1/2})$. For $R_{12}$, we further suppress the dependency of $S_{1}(k\mid X)$, $G_{1}(k\mid X)$ and their estimators on both $k$ and $X$, and rewrite the term as $R_{12}=E[\{\hat{e}(G_{1}-\hat{G}_{1})+G_{1}(e-\hat{e})\}\{\hat{\mu}_{1}(S_{1}-\hat{S}_{1})+S_{1}(\mu_{1}-\hat{\mu}_{1})\}/(\hat{e}\hat{G}_{1})]$. Because $\hat{e}/\hat{G}_{1}$ is bounded, by the Cauchy-Schwarz inequality, the first term in $R_{12}$ is upper bounded by $||G_{1}-\hat{G}_{1}||_{2}||S_{1}-\hat{S}_{1}||_{2}=o_{P}(n^{-1/2})$ by Assumption 5. The other three terms in $R_{12}$ can be similarly bounded. Thus, $R_{1}(P,\hat{P}_{n})=o_{P}(n^{-1/2})$.
Symmetric arguments imply the analogous term for the control arm satisfies $R_{0}(P,\hat{P}_{n})=o_{P}(n^{-1/2})$. Therefore, results in Theorem 4 follow from the central limit theorem. $\square$

### v2

@@
 Abstract
-The analysis of randomized controlled trials is often complicated by intercurrent events—events that occur after treatment initiation and may impact outcome assessment.
-These events may lead to patients discontinuing their assigned treatment or dropping out of the trial entirely.
-In an analysis of data from two recent immunology trials, we categorize intercurrent events into two broad types: those unrelated to treatment (e.g., withdrawal from the study due to external factors like pandemics or relocation) and those related to treatment (e.g., adverse events or lack of efficacy).
-We adopt distinct strategies to handle each type, aiming to target a clinically more relevant estimand.
-For treatment-related intercurrent events, they often meaningfully describe the patient’s outcome, we employ a composite variable strategy, where we attribute an outcome value that reflects the lack of treatment success.
-For treatment-unrelated intercurrent events, we adopt a hypothetical strategy that assumes these event times are conditionally independent of the outcome, given treatment and covariates, and envisions a scenario in which the intercurrent events do not occur.
-We establish the nonparametric identification and semiparametric estimation theory for the causal estimand and introduce doubly robust estimators.
-We illustrate our methods through the re-analysis of two randomized trials on baricitinib for Systemic Lupus Erythematosus.
-We classify intercurrent events, apply four estimators, and compare our approach with common ad-hoc methods, highlighting the robustness and practical implications of our framework.
-Keywords: Causal inference; Censoring; Estimand; Post-treatment variable; Potential outcome
+The analysis of randomized controlled trials is often complicated by intercurrent events (ICEs) – events that occur after treatment initiation and affect either the interpretation or existence of outcome measurements.
+Examples include treatment discontinuation or the use of additional medications.
+In two recent clinical trials for systemic lupus erythematosus with complications of ICEs, we classify the ICEs into two broad categories: treatment-related (e.g., treatment discontinuation due to adverse events or lack of efficacy) and treatment-unrelated (e.g., treatment discontinuation due to external factors such as pandemics or relocation).
+To define a clinically meaningful estimand, we adopt tailored strategies for each category of ICEs.
+For treatment-related ICEs, which are often informative about a patient’s outcome, we use the composite variable strategy that assigns an outcome value indicative of treatment failure.
+For treatment-unrelated ICEs, we apply the hypothetical strategy, assuming their timing is conditionally independent of the outcome given treatment and baseline covariates, and hypothesizing a scenario in which such events do not occur.
+A central yet previously overlooked challenge is the presence of competing ICEs, where the first ICE censors all subsequent ones.
+Despite its ubiquity in practice, this issue has not been explicitly recognized or addressed in previous data analyses due to the lack of rigorous statistical methodology.
+In this paper, we propose a principled framework to formulate the estimand, establish its nonparametric identification and semiparametric estimation theory, and introduce weighting, outcome regression, and doubly robust estimators.
+We apply our methods to analyze the two systemic lupus erythematosus trials, demonstrating the robustness and practical utility of the proposed framework.
+Keywords: Causal inference; Clinical trial; International Council for Harmonization; Post-treatment variable; Potential outcomes
 1 Intercurrent events in randomized controlled trials
-Randomized controlled trials (RCTs) are considered the gold standard for evaluating treatment efficacy.
-However, after treatment initiation, various events—referred to as intercurrent events (ICEs)—can arise, impacting the interpretation or availability of the outcome of interest, and posing significant challenges to the analysis of RCTs.
-These ICEs can include treatment discontinuation due to adverse events, lack of efficacy, or study withdrawal due to patient relocation.
-Addressing these ICEs carefully is crucial to ensure the validity and reliability of the causal conclusions drawn from RCTs.
-Given the complexities introduced by ICEs, clear guidance on how to handle them is crucial for drawing valid causal conclusions.
-The International Conference on Harmonization (ICH) has established global standards for drug development to align practices among pharmaceutical companies and regulatory agencies.
-Recognizing the importance of well-defined estimands in the presence of ICEs, the ICH introduced the E9 (R1) Addendum (ICH E9 , 2019, R1), a key resource that provides updated statistical principles to support clear and consistent approaches to analyzing and interpreting RCT results.
-The guidance has become essential for companies in drug research and development, and various strategies have been proposed and utilized to address ICEs following the guidance (Qu et al.,, 2021; Han and Zhou,, 2023).
-One straightforward approach to address the complications from the ICEs is the intention-to-treat (ITT) analysis.
-The ITT strategy includes all participants in their originally assigned groups and uses the observed outcome value, regardless of whether they completed the treatment or adhered to the study protocol.
-This strategy, also known as the treatment policy strategy in ICH E9 , 2019 (R1), is relevant for policy decisions and widely accepted.
-However, it fails to address ICEs that are terminal events, such as death, since these events directly impact the existence of the outcome variable.
-To address the limitation of ITT analyses and approach the issue from a different perspective, the composite variable strategy modifies the causal parameter of interest from the commonly used average causal effect to a comparison between two composite potential outcomes across the treatment arms.
-The composite potential outcome is defined as the endpoint of interest if no ICE occurs before its measurement, and a pre-specified, clinically meaningful value indicates treatment failure otherwise.
-The composite variable strategy is widely used in practice across different types of endpoints: binary endpoints (e.g., the non-responder imputation approach), ordinal or continuous endpoints (e.g., Rosenbaum,, 2006), and time-to-event endpoints (e.g., progression-free survival).
-In addition to the treatment policy strategy and the composite outcome strategy, three additional strategies for addressing ICEs were provided in ICH E9 , 2019 (R1).
-The hypothetical strategy considers treatment effects under a counterfactual scenario where ICEs do not occur or occur differently.
-The remaining two strategies–the while-on-treatment strategy and the principal stratification strategy–are less commonly used in practice.
+Randomized controlled trials (RCTs) are considered the gold standard for evaluating treatment efficacy, primarily because randomization supports assumption-lean inference of the treatment effect.
+However, after treatment initiation, various events, referred to as intercurrent events (ICEs), can arise, impacting the interpretation or availability of outcome measurements and posing significant challenges to the analysis of RCTs.
+Examples of ICEs include treatment discontinuation due to adverse events or lack of efficacy, patient relocation, and the use of additional medications.
+Carefully accounting for ICEs is essential to ensure the validity and reliability of the causal conclusions drawn from RCTs.
+Recognizing the critical need for clearly defined estimands in the presence of ICEs, the International Council for Harmonization (ICH) issued the E9(R1) Addendum (ICH E9 , 2019, R1).
+This addendum introduces a structured estimand framework for clinical trials to obtain precisely defined treatment effects that align with the clinical questions of interest.
+It outlines strategies for addressing ICEs during the formulation of the clinical question and emphasizes that careful specification of the treatment, population, and outcome variable often addresses many of the ICEs raised in discussions between sponsors and regulators.
+Since its release, the ICH E9(R1) Addendum has been widely discussed, increasingly adopted in clinical drug development, and has sparked substantial interest in statistical research (Qu et al.,, 2021; Kang et al.,, 2022; Ionan et al.,, 2023; Han and Zhou,, 2023; Olarte Parra et al.,, 2025).
+A widely accepted strategy is the treatment policy strategy (ICH E9 , 2019, R1), which includes all participants in their originally assigned groups and uses the observed outcome values, regardless of whether or not an ICE occurs.
+This approach aligns with the intention-to-treat principle and reflects the effect of a treatment policy in real clinical settings.
+However, it cannot address ICEs that are terminal events, such as death, because such events preclude the existence of the outcome variable.
+The composite variable strategy is an alternative approach that is well-suited for handling ICEs that are informative of the patient’s outcome such as the terminal events.
+This strategy incorporates the occurrence of ICEs directly into the outcome definition.
+Specifically, it defines a composite outcome: if no ICE occurs, the outcome of interest is used as observed; if an ICE occurs, a pre-specified, clinically meaningful value—typically indicating treatment failure—is assigned.
+This strategy is widely used in practice across various types of outcomes, including binary outcomes (e.g., the non-responder imputation approach), ordinal or continuous outcomes (e.g., Rosenbaum,, 2006), and time-to-event outcomes (e.g., progression-free survival).
+In addition to the treatment policy and composite variable strategies, ICH E9 , 2019 (R1) outlines three additional strategies for handling ICEs.
+The hypothetical strategy evaluates treatment effects under a hypothetical scenario in which the ICE would not occur.
+The while-on-treatment and principal stratification strategies are also described but are less commonly used in current practice due to their reliance on strong assumptions and the potential to introduce bias in treatment comparisons.
 ICH E9 , 2019 (R1) recommends using different strategies based on the specific type of ICEs involved.
-However, very few papers have formally explored how to combine different strategies and rigorously addressed the pitfalls of doing so naively (Lipkovich et al.,, 2020; Qu et al.,, 2021).
-1.1 A phase-3 immunology trial
-To illustrate our setup and notation, we first introduce a real phase-3 immunology trial (Morand et al.,, 2023; Petri et al.,, 2023), which we will revisit in Section 6.
-This trial investigates the causal effect of baricitinib versus placebo on Systemic Lupus Erythematosus.
-Patients in the RCT were assigned either to a baricitinib treatment group or to a placebo control group.
-The primary endpoint of interest is an immune response index measured 52 weeks after treatment initiation.
-Ideally, we would compare this outcome across the two groups at week 52.
-However, 429 out of 1535 (27.95%) patients experienced ICEs between treatment initiation and the outcome measurement, leading to missingness of the outcome.
+While some studies have explored the use of multiple strategies (Lipkovich et al.,, 2020; Qu et al.,, 2021), none have rigorously examined how to combine different strategies, nor have they rigorously addressed the potential pitfalls of doing so without careful consideration.
+Moreover, the challenges posed by competing ICEs remain unrecognized.
+1.1 Two phase-3 trials in systemic lupus erythematosus
+To rigorously examine the challenges posed by ICEs in clinical trials, we analyze data from two recent randomized studies for systemic lupus erythematosus (Morand et al.,, 2023; Petri et al.,, 2023), in which participants were randomly assigned to either the baricitinib treatment group or the placebo control group.
+These twin trials were designed to provide substantiated evidence on the causal treatment effect of baricitinib versus placebo.
+The primary outcome is a response index measured at 52 weeks after treatment initiation.
+Ideally, this outcome would be compared directly between the two treatment groups at week 52.
+However, 429 out of 1,535 patients (27.95%) experienced ICEs during the follow-up period, resulting in unobserved outcome data.
 Figure 1 shows the types and proportions of these ICEs.
-During the 52 weeks, various ICEs happened.
-For instance, some patients discontinued the assigned treatment due to adverse events or lack of efficacy, some had protocol deviation and were excluded from the study, some of them withdrew from the study for unknown reasons, and some were lost to follow-up.
-Figure 1: Pie chart showing the ICE types and proportions
+During the 52-week period, various ICEs occurred: some patients discontinued treatment due to adverse events or lack of efficacy; others discontinued treatment due to study withdrawal for unspecified reasons or were lost to follow-up; and some were excluded from the study due to protocol deviations.
+We revisit this example in greater detail in Section 5.
+Figure 1: Pie chart showing the ICE types and proportions.
+The yellow solid charts represent treatment-related ICEs, the green solid chart represents a treatment-unrelated ICE, and the blue striped charts represent a mix of both types.
 1.2 Our proposal and contribution
-We propose to classify ICEs into two broad types: (1) treatment-unrelated ICEs, such as treatment discontinuation due to relocation or COVID-19 lockdown, which are assumed to be independent of treatment efficacy conditional on the observed covariates; and (2) treatment-related ICEs, such as treatment discontinuation due to adverse events or lack of efficacy, use of rescue medication, and terminating events such as death, which are likely correlated with both treatment and outcomes.
-Figure 1 illustrates these classifications: yellow solid charts represent treatment-related ICEs, the green solid chart represents a treatment-unrelated ICE, and blue stripped charts represent a mix of both types.
-For instance, patients withdrew from the study for various reasons–some due to relocation or the pandemic, which can be considered as unrelated to the treatment, while others withdrew due to concerns about potential side effects, which are likely to be treatment-related.
-When patients did not disclose a detailed reason, we adopt a conservative approach and classify their ICEs as treatment-related.
-Both types of ICEs may occur between treatment initiation and outcome measurement.
-Figure 2 depicts the timeline and three scenarios in our motivating example.
-Before week 52, patient 1 withdrew from the study due to relocation, which is arguably unrelated to the treatment, patient 2 dropped out of the study due to an adverse event, and patient 3 completed the study through the endpoint at week 52, when the outcome of interest was measured.
-Figure 2: Illustration of the motivating immunology trial example
-Both types of ICEs lead to missing primary endpoints measured at a landmark time because the outcome cannot be observed or is not even well-defined when an ICE occurs.
-For the first type of ICE, we apply the hypothetical strategy and assume the ICE event happens at random conditional on the observed covariates.
-It considers the hypothetical scenario in which the ICE did not occur.
-For the second category, we utilize the composite variable strategy, incorporating the ICE itself as indicative of treatment failure, for which we provide a formal definition and discussion in Section 3.
-A main challenge in combining the two strategies and our key insight is that the two types of ICEs can compete with each other, meaning that patients discontinue due to the first reason that occurs.
-Use study withdrawal due to relocation and adverse events as examples to represent the two types of ICEs, respectively.
-For example, suppose a patient discontinues due to relocation like patient 1 in Figure 2.
-In that case, it remains unknown whether an adverse event would have occurred before the final endpoint measurement in the counterfactual scenario had the withdrawal due to relocation ICE not occurred.
-It is important to carefully deal with the issue of competing ICEs, which, to the best of our knowledge, has never been pointed out in the literature.
-We consider this feature and provide the nonparametric identification of our causal parameter of interest under two assumptions: (1) treatment assignment satisfies the ignorability and overlap assumptions, and (2) the treatment-unrelated ICE time is independent of the potential outcomes and the potential values of the treatment-related ICE time, given observed covariates.
-We present two identification formulas relying on different sets of nuisance models.
-We then propose two basic estimators based on these two identification formulas and a third estimator that combines them in a similar way to the construction of the augmented inverse probability weighting estimator to achieve double robustness.
-The combined estimator can be further improved in terms of both robustness and asymptotic efficiency.
-To construct a more principled estimator that simultaneously achieves robustness and semiparametric efficiency, we derive the efficient influence function (EIF) thus the semiparametric efficiency bound of our causal parameter.
-We propose an EIF-based estimator and study its asymptotic properties.
-We show the EIF estimator is doubly robust, meaning its consistency does not require all nuisance parameters to be consistently estimated—only one of the two sets of nuisance models needs to be correctly specified.
-Moreover, the EIF estimator achieves the semiparametric efficiency bound under certain conditions on nuisance estimation.
-The remaining of the paper proceeds as follows.
-In Section 2, we provide a motivating example, introduce the basic setup of our research question, describe the composite variable strategy, define our causal parameter of interest, and illustrate the identification challenge.
-In Section 3, we present the key identification assumptions, provide the nonparametric identification of our causal parameter, and construct three basic estimators.
-In Section 4, we derive the EIF for our causal parameter, propose a doubly robust and asymptotically efficient estimator based on the EIF, and study its asymptotic properties.
-In Section 5, we run Monte Carlo simulations to evaluate the finite sample performance of our proposed estimation procedures.
-In Section 6, we implement the estimators to re-analyze two randomized controlled trials that study the causal effect of baricitinib treatment on systemic lupus erythematosus.
-In Section 7, we discuss two future research directions.
+We propose to classify ICEs into two broad types: (1) treatment-unrelated ICEs, such as treatment discontinuation due to relocation or COVID-19 lockdown, which are assumed to be independent of treatment efficacy conditional on the observed covariates; and (2) treatment-related ICEs, such as treatment discontinuation due to adverse events or lack of efficacy, use of rescue medication, and terminating events such as death, which are often informative about a patient’s outcome.
+In our application studies, the two types of ICEs were classified based on a manual review of the detailed comments collected at the clinical sites.
+Figure 1 summarizes the main categories and illustrates our classification scheme: yellow solid charts represent treatment-related ICEs, the green solid chart represents a treatment-unrelated ICE, and the blue striped charts represent a mix of both types.
+For instance, some patients withdrew from the study due to external factors such as relocation or the COVID-19 pandemic, which were classified as treatment-unrelated.
+Others withdrew due to concerns about potential side effects, which were considered treatment-related.
+When the reason for withdrawal was not clearly stated, we adopted a conservative approach and classified the ICE as treatment-related.
+Both types of ICEs may occur during the follow-up period.
+Figure 2 depicts the timeline and three representative scenarios from our motivating example.
+Before week 52, patient 1 withdrew from the study due to relocation, which is arguably unrelated to treatment.
+Patient 2 dropped out due to an adverse event, reflecting a treatment-related ICE.
+Patient 3 completed the study through week 52, at which point the primary outcome was measured.
+Figure 2: Illustration of three representative scenarios of intercurrent events in the immunology trial.
+The primary outcome is measured at week 52, and the solid lines indicate the observed follow-up period for each patient.
+Patient 1 experienced a treatment-unrelated ICE (represented by an arrow), Patient 2 experienced a treatment-related ICE (represented by a diamond), and Patient 3 had no ICE before week 52.
+This figure highlights the issue of competing ICEs: for instance, had Patient 1 not discontinued early and remained on treatment (illustrated by the gray dashed line), they might have experienced a treatment-related ICE later in the follow-up period.
+Both types of ICEs lead to unmeasured primary outcomes at the pre-specified landmark time because the outcome cannot be observed or may not even be well-defined when a treatment-related ICE occurs.
+To ensure the estimand is clinically interpretable and relevant, we apply different strategies tailored to the nature of the ICE.
+For ICEs that are plausibly unrelated to treatment (e.g., treatment discontinuation due to relocation or administrative withdrawal), we apply the hypothetical strategy, which imagines a scenario where the ICE did not occur.
+This enables estimation of the treatment effect as if the patient had remained on treatment and in the study as planned, under the assumption that the ICE time is conditionally independent of the outcome given treatment and baseline covariates.
+This assumption is reasonable in settings where the ICE arises from external or administrative factors that are not influenced by post-randomization conditions.
+For example, if a patient relocates for personal reasons unrelated to their health status or treatment response, then, conditional on baseline characteristics and assigned treatment, the occurrence of this ICE can be viewed as independent of the patient’s outcome.
+In contrast, for treatment-related ICEs (e.g., treatment discontinuation due to adverse events or lack of efficacy), we apply the composite variable strategy, treating the ICE itself as an indication of treatment failure, as such events signify that the patient is unable to continue treatment.
+A formal definition and discussion of this strategy are provided in Section 3.
+Combining the two strategies enhances the interpretability of the estimand and ensures it better reflects clinically relevant questions, as no single strategy can adequately address both types of ICEs.
+Specifically, it would be inappropriate to classify a patient as a non-responder if they discontinue treatment due to relocation, and it is of limited clinical relevance to consider a hypothetical scenario in which adverse events do not occur.
+This approach is consistent with regulatory guidance, which recommends applying different strategies depending on the type of ICE involved (ICH E9 , 2019, R1; Kang et al.,, 2022).
+However, a central yet previously overlooked challenge is the presence of competing ICEs, where the first ICE censors all subsequent ones.
+For example, if a patient discontinues treatment due to relocation, as illustrated by patient 1 in Figure 2, we only observe that no treatment-related ICE occurred before the discontinuation.
+However, it remains unknown whether an adverse event would have occurred before the final outcome measurement in the hypothetical scenario in which the patient had not relocated and had continued treatment as planned.
+Although such competing ICEs are ubiquitous in practice, this issue has not, to our knowledge, been explicitly recognized or addressed in prior methodological work.
+The central thesis of this paper is to address the challenge of competing ICEs.
+We begin by developing a principled framework that clearly defines the estimand and establishes its nonparametric identification.
+Specifically, we derive two identification formulas, each relying on a different set of nuisance parameters.
+Building on these results, we propose two basic estimators corresponding to the two identification strategies.
+We then introduce an augmented estimator that combines the two, achieving double robustness.
+To further enhance both robustness and efficiency, we derive the efficient influence function (EIF) and construct an EIF-based estimator that attains the semiparametric efficiency bound under appropriate conditions.
+We apply our methods to the two systemic lupus erythematosus trials, both partially impacted by the COVID-19 pandemic, demonstrating the robustness and practical utility of the proposed framework.
+1.3 Organization and notation
+The remainder of the paper is organized as follows.
+In Section 2, we present a motivating example, introduce the basic setup of our research question, define the causal parameter of interest, and highlight the identification challenge.
+In Section 3, we state the key identification assumptions, establish the nonparametric identification of the causal parameter, and construct three estimators.
+In Section 4, we derive the EIF, propose an estimator that is both doubly robust and asymptotically efficient, and examine its asymptotic properties.
+In Section 5, we apply the methods to re-analyze the two systemic lupus erythematosus trials.
+Finally, in Section 6, we conclude with a discussion of two directions for future research.
+The supplementary material includes a simulation study and all proofs.
+We also provide publicly available R code implementing all four proposed estimators.
 We use the following notation.
 Let $\lVert r\rVert_{2}=\{\int r(o)^{2}\textnormal{d}P(o)\}^{1/2}$ denote the $L_{2}(P)$ norm where $P(\cdot)$ denotes the distribution of the observed data $O=o$.
-For the survival and censoring functions, let $\lVert r\rVert_{2}=\{\iint r(t,o)^{2}\textnormal{d}P(o)\textnormal{d}t\}^{1/2}$ denote the $L_{2}(P)$ norm.
+For the survival functions, let $\lVert r\rVert_{2}=\{\iint r(t,o)^{2}\textnormal{d}P(o)\textnormal{d}t\}^{1/2}$ denote the $L_{2}(P)$ norm.
 We write $b_{n}=O_{P}(a_{n})$ if $b_{n}/a_{n}$ is bounded in probability and $b_{n}=o_{P}(a_{n})$ if $b_{n}/a_{n}$ converges to 00 in probability.
-2 Basic setup
-2.1 Setup and notation
-In this subsection, we introduce our formal setup and mathematical notation.
-Let $A$ denote the binary treatment indicator, with $A=1$ for patients in the treatment group, and $A=0$ for those in the control group.
-The primary outcome, denoted as $Y$, is measured at a pre-specified time point $k$.
-In our motivating example, $Y$ is the immune response index and $k$ is 52 weeks.
-We adopt the potential outcomes framework (Neyman,, 1923; Rubin,, 1974), where $Y(a)$ denotes the potential outcome under treatment assignment $A=a$ for $a=0,1$, and $Y=Y(A)=AY(1)+(1-A)Y(0)$ is the observed outcome.
-As discussed, two types of ICEs may occur between the treatment initiation and the measurement of the outcome: treatment-unrelated ICEs, assumed independent of the potential outcomes conditional on covariates, and treatment-related ICEs, which likely correlate with the endpoints.
-To formally define the two types of ICEs, let $C$ and $T$ denote time to treatment-unrelated and time-to-treatment-related ICEs, respectively.
-Both are post-treatment variables, thus we use $C(a)$ and $T(a)$ to denote their potential values under treatment $a$, for $a=0,1$, with observed values $C=C(A)=AC(1)+(1-A)C(0)$ and $T=T(A)=AT(1)+(1-A)T(0)$.
-2.2 Causal parameter of interest and challenge in identification
-To address treatment-related ICEs, we define a composite potential outcome as $Y^{\textup{cp}}(a)=Y(a)1\{T(a)>k\}$ for $a=0,1$ where the superscript stands for “composite” and define our parameter of interest as the mean contrast, i.e.,
+2 Setup and estimand
+2.1 Setup
+Let $A$ denote the binary treatment indicator, where $A=1$ corresponds to assignment to the treatment group and $A=0$ to the control group.
+The primary outcome, denoted by $Y$, is measured at a pre-specified time point $k$.
+In our motivating example, $Y$ represents the response index, and $k$ corresponds to 52 weeks.
+We adopt the potential outcomes framework, in which $Y(a)$ denotes the potential outcome under treatment assignment $A=a$, for $a=0,1$.
+The observed outcome is then given by $Y=Y(A)=AY(1)+(1-A)Y(0)$.
+As discussed in Section 1, two types of ICEs may occur after treatment initiation: treatment-unrelated ICEs and treatment-related ICEs.
+To formally define these, let $C$ and $T$ denote the time to a treatment-unrelated ICE and the time to a treatment-related ICE, respectively.
+Since both are post-treatment variables, we use $C(a)$ and $T(a)$ to denote their potential values under treatment assignment $A=a$, for $a=0,1$.
+The observed times are then given by $C=C(A)=AC(1)+(1-A)C(0)$ and $T=T(A)=AT(1)+(1-A)T(0)$.
+2.2 Causal estimand and challenges in identification
+To address treatment-related ICEs, we define a composite potential outcome as
+$Y^{\textup{cp}}(a)=Y(a)1\{T(a)>k\}$
+for $a=0,1$, where the superscript denotes “composite”.
+This construction defines the outcome as zero (i.e., failure) if a treatment-related ICE occurs before the outcome is measured.
+We then define our causal estimand as the mean contrast:
 $\tau=E\{Y^{\textup{cp}}(1)-Y^{\textup{cp}}(0)\}=E[Y(1)1\{T(1)>k\}]-E[Y(0)1\{T(0)>k\}].$ (1)
-This target parameter is widely used for binary outcomes to capture success or failure, where treatment-related ICEs, such as adverse events, are treated as failure (ICH E9 , 2019, R1).
-In our motivating example, $Y(a)$ measures whether a patient is an immune responder at week 52 after taking the treatment $A=a$, which is unobserved for patients experiencing a treatment-related ICE such as an adverse event or lack of efficacy.
-Such patients are defined as non-responders under the composite variable strategy.
-The composite variable approach extends beyond binary outcomes to accommodate scenarios where a clinically significant value can denote treatment failure.
-More specifically, for outcomes such as quality of life, chronic pain, physical functioning, and cognitive performance that are measured on the ordinal and continuous scale, if there is a predefined clinically meaningful value $v$ that can signal treatment failure, then a composite outcome under treatment arm $A=a$ may be constructed as $Y(a)1\{T(a)>k\}+v1\{T(a)\leq k\}$, in which case the causal parameter is defined as in (1) with $Y(a)$ replaced by $Y(a)-v$ for $a=0,1$.
-Rosenbaum, (2006) discussed similar causal parameters when the ICE is death and the outcome of interest is some measure of life quality after a certain period.
-Identifying $\tau$ in the presence of both ICE types is challenging.
-For clarity, in the remainder of this subsection, we focus on withdrawal due to relocation and adverse events to represent the treatment-unrelated and treatment-related ICEs, respectively.
-The definition of the causal parameter $\tau$ in (1) essentially treats $Y^{\textup{cp}}(a)$ as the relevant outcome, assigning 0 when adverse events occur, i.e., when $T(a)<k$.
-If adverse events were the only reason for missingness in the outcome, we could identify $\tau$ using the difference in means of the observed composite outcomes $Y^{\textup{cp}}=Y^{\textup{cp}}(A)=Y1(T>k)$, i.e.,
-$\tau$ $=$ $E\{Y1(T>k)\mid A=1\}-E\{Y1(T>k)\mid A=0\}.$
-However, this formula is infeasible to learn from observed data because treatment-unrelated ICEs, such as withdrawal due to relocation, could happen before treatment-related ICEs thus censoring $T(a)$.
-Figure 2 shows how withdrawal due to relocation in patient 1 censored their adverse event time and the outcome, creating missingness in the composite outcome $Y1(T>k)$ for such patients.
-The key challenge is that the two types of ICEs compete with each other, and patients discontinue due to the first ICE that occurs.
-Table 1 summarizes observed ICE types and composite outcomes, highlighting that censoring complicates the identification of $\tau$.
-When there is no ICE, we observe the outcome $Y$, which equals the composite outcome because $T>k$.
-When we observe an adverse event, the composite outcome is by definition 0.
-When we observe a withdrawal due to relocation, both the outcome of interest and the time-to-adverse event are missing, and so is the composite outcome.
-We address the challenge and present identification results in the following section.
+This is the estimand under the composite variable strategy.
+It is widely used for binary outcomes to represent treatment success or failure, where treatment-related ICEs such as adverse events or lack of efficacy are treated as failures (ICH E9 , 2019, R1).
+In our motivating example, $Y(a)$ indicates whether a patient would be a responder at week 52 under treatment assignment $A=a$, had no ICE occurred.
+For patients who experience a treatment-related ICE before week 52, $Y(a)$ is unobserved.
+Under the composite variable strategy, these patients are defined as non-responders.
+The composite variable approach extends beyond binary outcomes and is applicable in settings where treatment failure can be defined using a clinically meaningful threshold.
+Specifically, for outcomes such as quality of life, chronic pain, physical functioning, and cognitive performance that are typically measured on an ordinal or continuous scale, if a predefined value $v$ may be used to indicate treatment failure, then a composite outcome under treatment assignment $A=a$ can be constructed as $Y(a)1\{T(a)>k\}+v1\{T(a)\leq k\}$, where the outcome retains its actual value if no treatment-related ICE occurs before time $k$, and takes the failure value $v$ otherwise.
+The corresponding causal estimand can then be defined as in (1), with $Y(a)$ replaced by $Y(a)-v$ for $a=0,1$.
+Rosenbaum, (2006) discussed a similar causal parameter in settings where the ICE is death and the outcome of interest is a measure of quality of life assessed after a fixed period.
+Identifying $\tau$ in the presence of both types of ICEs presents a key methodological challenge.
+For clarity, in the remainder of this subsection, we use treatment discontinuation due to relocation and adverse events to represent treatment-unrelated and treatment-related ICEs, respectively.
+The causal estimand $\tau$ defined in (1) is based on the composite potential outcome $Y^{\textup{cp}}(a)$, which takes the value 00 if an adverse event occurs before the outcome measurement, i.e., when $T(a)<k$.
+If adverse event were the only type of ICE, and treatment is randomized, $\tau$ could be identified from observed data using the difference in means of the composite outcomes: $\tau=E\{Y1(T>k)\mid A=1\}-E\{Y1(T>k)\mid A=0\}$.
+However, this formula is infeasible in the presence of treatment-unrelated ICEs, which may occur before treatment-related ICEs and thus censor both $T(a)$ and $Y$.
+As illustrated in Figure 2, patient 1 discontinued due to relocation, which censored both the occurrence of a potential adverse event and the outcome, resulting in an unobserved composite outcome $Y{1}(T>k)$.
+Therefore, for patients who experience treatment-unrelated ICEs, the problem cannot be treated as standard censoring, and one cannot naïvely impute the outcome $Y$ under a hypothetical scenario where the treatment-unrelated ICE did not occur.
+This is because such patients may still have experienced a treatment-related ICE had they not discontinued due to relocation.
+In other words, a valid identification strategy must recover the expected value of the composite outcome $Y{1}(T>k)$, rather than the outcome $Y$ alone.
+Table 1 summarizes the observed ICE types and corresponding composite outcomes, highlighting that competing ICEs complicate the identification of $\tau$.
+When no ICE occurs, the outcome $Y$ is observed and equals the composite outcome since $T>k$.
+When an adverse event is observed, the composite outcome is, by definition, 0.
+However, when a patient discontinues treatment due to relocation, both the outcome $Y$ and the time to the adverse event $T$ are unobserved, and so is the composite outcome.
+We address this challenge and present formal identification results in the following section.
 Table 1: Summary of the observed ICE types and outcome.
-$C$ is the time to withdrawal due to relocation (WR), $T$ is the time to adverse event (AE), $Y$ is the outcome of interest, and $k$ is the pre-specified time point when the measurement of $Y$ is taken.
+$C$ is the time to treatment discontinuation due to relocation, $T$ is the time to adverse event, $Y$ is the outcome of interest, and $k$ is the pre-specified time point when the measurement of $Y$ is taken.
+A question mark “?” indicates that the corresponding value is unobserved.
 observed ICE type $(T,C,k)$-relationship $Y$ $1(T>k)$ composite outcome $Y^{\textup{cp}}$
-no AE/WR $C\wedge T>k$ $Y$ 1 $Y$
-AE $C\wedge k>T$ 0 0
-WR $T\wedge k>C$
-3 Identification and three basic estimators
+no any type of ICE $C\wedge T>k$ $Y$ 1 $Y$
+adverse event $C\wedge k>T$ 0 0
+treatment discontinuation due to relocation $T\wedge k>C$
+3 Nonparametric identification and basic estimators
 3.1 Assumptions
-We assume the joint distribution of $\{X_{i},A_{i},T_{i}(1),T_{i}(0),C_{i}(1),C_{i}(0),Y_{i}(1),Y_{i}(0)\}$ for patient $i$ is independently and identically distributed from a super population and will drop the subscript $i$ whenever there is no confusion.
-We first assume the following ignorability and overlap assumption for the identification of our causal parameter of interest.
+We assume that the joint distribution of $\{X_{i},A_{i},T_{i}(1),T_{i}(0),C_{i}(1),C_{i}(0),Y_{i}(1),Y_{i}(0)\}$ for patient $i$ is independently and identically distributed from a superpopulation.
+For notational simplicity, we omit the subscript $i$ when there is no confusion.
+We begin by stating the following assumption on treatment assignment.
 Assumption 1 (Treatment assignment)
-We assume the following conditional independence and overlap conditions for the treatment assignment:
+We assume the following conditional independence and overlap conditions:
   • (a)
@@
   • (b)
-    For some constant $\eta\in(0,0.5)$, $\eta<e(X)<1-\eta$ with probability 1, where $e(X)=\textup{pr}(A=1\mid X)$ denotes the propensity score.
-Assumption 1 is guaranteed by the design of an RCT, allowing our results to apply directly to RCTs like our motivating example.
-However, our formulation is more general as it can also accommodate observational studies.
-Under Assumption 1, we have the following equations for $\tau$
+    For some constant $\eta\in(0,0.5)$, $\eta<e(X)<1-\eta$ with probability 1, where $e(X)=\textup{pr}(A=1\mid X)$ denotes the propensity score (Rosenbaum and Rubin,, 1983).
+Assumption 1 holds in RCTs by design, allowing our results to apply directly to settings such as our motivating example.
+However, our formulation is more general and can also accommodate cases where treatment assignment is not completely randomized, such as in stratified randomized experiments and observational studies.
+Under Assumption 1, the causal estimand $\tau$ can be expressed as:
 $\tau$ $=$ $E[E\{Y1(T>k)\mid A=1,X\}-E\{Y1(T>k)\mid A=0,X\}]$
 $=$ $E\left\{\frac{AY1(T>k)}{e(X)}-\frac{(1-A)Y1(T>k)}{1-e(X)}\right\},$
-which are the classic outcome regression and the inverse probability weighting identification formulas when we treat the composite potential outcomes $Y^{\textup{cp}}(a)=Y(a)1\{T(a)>k\}$ as the “new” outcome variable of interest with $Y^{\textup{cp}}=Y1(T>k)$.
-However, as discussed in Section 2.2, they do not serve as identification formulas anymore with the presence of the competing treatment-unrelated ICEs because $Y1(T>k)$ is no longer fully observed if treatment-unrelated ICE happened before both $T$ and $k$.
-Therefore, we need further identification assumptions on the time of treatment-unrelated ICE.
-Specifically, we assume it is independent of all other potential values, including the potential outcomes, which is equivalent to the censoring at random assumption in the survival analysis literature when we treat treatment-unrelated ICEs as censoring.
-In addition, we need the censoring probability to be bounded away from 1.
-We summarize them in the following assumption.
-Assumption 2 (Time of treatment-unrelated ICE)
-We assume the following conditional independence and positivity conditions for the time of treatment-unrelated ICE.
+which correspond to the standard outcome regression and inverse probability weighting identification formulas, respectively, when treating the composite potential outcome $Y^{\textup{cp}}(a)=Y(a){1}\{T(a)>k\}$ as the “new” outcome of interest.
+However, as discussed in Section 2.2, these expressions no longer serve as feasible identification formulas in the presence of competing treatment-unrelated ICEs.
+This is because $Y{1}(T>k)$ is not fully observed if a treatment-unrelated ICE occurs.
+To address this challenge, we introduce additional assumptions regarding treatment-unrelated ICEs.
+Specifically, we assume that the potential time of a treatment-unrelated ICE under treatment $a$ is conditionally independent of the potential outcomes and the potential time to treatment-related ICEs, given baseline covariates.
+This assumption is analogous to the censoring-at-random assumption in survival analysis when treating treatment-unrelated ICEs as censoring events (Robins et al.,, 1994; Tsiatis,, 2006).
+Additionally, we require that the probability of no occurrence of treatment-unrelated ICE by time $k$, given observed covariates, is strictly positive.
+These conditions are summarized in the following assumption.
+Assumption 2 (Treatment-unrelated ICE)
+We assume the following conditional independence and positivity conditions for the time of treatment-unrelated ICE:
   • (a)
@@
     For some constant $\eta_{C}>0$, $\textup{pr}\{C(a)>k\mid X\}>\eta_{C}$ with probability 1.
-In our motivating example, Assumption 2 states that the time to treatment-unrelated ICEs, such as withdrawal due to relocation, is independent of both the potential outcome of whether a patient will be a responder at week 52 and the time to an adverse event, conditional on observed baseline covariates.
-In other words, after accounting for factors such as region, baseline corticosteroid use, and the physician global assessment score, the decision to withdraw due to relocation does not provide additional information about the patients’ eventual response or the time to adverse events.
-Additionally, it assumes that every patient has a nonzero probability of not experiencing a treatment-unrelated ICE.
+In our motivating example, Assumption 2(a) states that the time to a treatment-unrelated ICE, such as treatment discontinuation due to relocation, is independent of both the potential outcome (i.e., whether a patient would be a responder at week 52) and the time to a treatment-related ICE (e.g., an adverse event), conditional on observed baseline covariates.
+In other words, after accounting for baseline covariates such as geographic region, baseline corticosteroid use, and the physician’s global assessment score, the decision to discontinue treatment due to relocation is assumed to be unrelated to the patient’s potential outcomes or potential timing of adverse events.
+Assumption 2(b) requires that every patient has a strictly positive probability of not experiencing a treatment-unrelated ICE before time $k$, given their covariates.
+This standard positivity condition rules out deterministic treatment-unrelated ICE happening in any subpopulation and is reasonable in our application.
 3.2 Two identification formulas
-Using the composite variable strategy for the ICEs that are related to the treatment, recall that we define the parameter of interest as:
-$\tau=\mu_{1}-\mu_{0}=E[Y(1)1\{T(1)>k\}]-E[Y(0)1\{T(0)>k\}].$
-Under Assumptions 1 and 2, we show the parameter of interest $\tau$ is nonparametrically identified and provide two identification formulas in the following theorem.
+Using the composite variable strategy to handle treatment-related ICEs, recall that we define the causal estimand as in (1).
+Under Assumptions 1 and 2, we show that the estimand $\tau$ is nonparametrically identifiable and present two identification formulas in the following theorem.
 Theorem 1 (Nonparametric identification of $\tau$)
@@
 $\tau$ $=$ $E\left\{\mu_{1}(X)S_{1}(k\mid X)-\mu_{0}(X)S_{0}(k\mid X)\right\},$ (2)
-where $\mu_{a}(X)=E(Y\mid T\wedge C>k,X,A=a)$ is the conditional mean of observed outcome among those with no ICE and $A=a$ and $S_{a}(t\mid X)=\textup{pr}(T>t\mid X,A=a)$ is the conditional survival function of treatment-related ICE time among those with $A=a$.
+where for $a=0,1$, $\mu_{a}(X)=E(Y\mid T\wedge C>k,X,A=a)$ is the conditional mean of observed outcome among those with no ICE and $A=a$, and $S_{a}(t\mid X)=\textup{pr}(T>t\mid X,A=a)$ is the conditional survival function of treatment-related ICE in treatment arm $a$.
 Second,
 $\tau$ $=$ $E\left[\frac{AY1(T\wedge C>k)}{e(X)G_{1}(k\mid X)}-\frac{(1-A)Y1(T\wedge C>k)}{\left\{1-e(X)\right\}G_{0}(k\mid X)}\right],$ (3)
-where $e(X)=\textup{pr}(A=1\mid X)$ is the propensity score and $G_{a}(t\mid X)=\textup{pr}(C>t\mid X,A=a)$ for $a=0,1$ is the survival function of $C$ in the subsample $A=a$.
-The first formula (2) expresses $\tau$ using a standard outcome regression approach.
-It models the conditional mean of the outcome among individuals who have not experienced ICE by $k$, adjusting for the probability of this event occurring.
-Equation (3) provides an alternative identification using inverse probability weighting.
-It identifies $\tau$ by averaging observed outcomes only among individuals who have not experienced ICE by time $k$ while correcting for the bias due to differences in the probability of observing a composite outcome.
-The weights adjust for two key factors: $e(X)$ that accounts for difference in treatment assignment probabilities, and $G_{a}(k\mid X)$ that adjusts for the likelihood of censoring in each treatment group.
-For instance, units with a lower probability of being observed are up-weighted, while those more likely to remain observed are down-weighted.
-Both equations (2) and (3) are valid identification formulas because the nuisance parameters involved are either functions of observed data or identified from observed data.
-More specifically, nuisance parameters $e(X)$ and $\mu_{a}(X)$ are functions of the observed data.
-The survival functions of the treatment-related time $S_{a}(t\mid X)$ and the censoring mechanism $G_{a}(t\mid X)$ for $t\leq k$, although not direct functions of observed data, are both identifiable due to the classic result on survival data with censoring at random (Robins and Rotnitzky,, 1992; Robins and Finkelstein,, 2000).
-For example, $S_{1}(t\mid X)=\textup{pr}(T>t\mid X,A=1)$ for $t\leq k$ can be estimated using observations in the treatment subgroup from treating $T\wedge C\wedge k$ as the observed event time and $1(C\wedge k\geq T)$ as the event status for the treatment-related ICE, and estimating using a Cox proportional hazard model (Cox,, 1972).
-We can use a similar procedure to estimate $G_{1}(t\mid X)$ by switching the roles of $T$ and $C$, and use similar procedures to estimate $S_{0}(t\mid X)$ and $G_{0}(t\mid X)$ by using observations in control subgroup.
+where $e(X)=\textup{pr}(A=1\mid X)$ is the propensity score and for $a=0,1$, $G_{a}(t\mid X)=\textup{pr}(C>t\mid X,A=a)$ is the conditional survival function of treatment-unrelated ICE in treatment arm $a$.
+The first identification formula (2) expresses $\tau$ using a standard outcome regression approach.
+It decomposes the conditional expectation of the composite outcome given covariates and treatment as $E\{YI(T>k)\mid X,A\}=E(Y\mid T>k,X,A)P(T>k\mid X,A)=E(Y\mid T\wedge C>k,X,A)P(T>k\mid X,A)$.
+This formulation models the conditional mean outcome $Y$ among individuals who do not have any ICE by time $k$, and weights it by the probability of not experiencing a treatment-related ICE by that time.
+The second identification formula (3) provides an alternative identification strategy based on inverse probability weighting.
+It identifies $\tau$ by reweighting observed outcomes among individuals who remain free of ICEs up to time $k$, separately within each treatment arm.
+The weights adjust for the different probability of observing the composite outcome and consist of two components: $e(X)$, which accounts for differences in treatment assignment probabilities, and $G_{a}(k\mid X)$, which adjusts for the probability of remaining free of treatment-unrelated ICEs up to time $k$ within each treatment group.
+Effectively, individuals with a lower probability of being observed are up-weighted, while those with a higher probability are down-weighted, ensuring an unbiased estimate of the treatment effect.
+Both equations (2) and (3) are valid identification formulas, as the nuisance parameters involved are either functions of observed data or identifiable from observed data.
+Specifically, the propensity score $e(X)$ and the conditional outcome model $\mu_{a}(X)$ are functions of observed data.
+Although the survival functions for both types of ICEs, $S_{a}(t\mid X)$ for treatment-related ICEs and $G_{a}(t\mid X)$ for treatment-unrelated ICEs, are not direct functions of the observed data, they are identifiable under standard results in survival analysis with censoring at random (Robins and Rotnitzky,, 1992; Robins and Finkelstein,, 2000; Ebrahimi et al.,, 2003).
+For example, $S_{1}(t\mid X)=\textup{pr}(T>t\mid X,A=1)$ for $t\leq k$ can be estimated using data from the treatment group by viewing $T\wedge C\wedge k$ as the observed event time and using $1(C\wedge k\geq T)$ as the event indicator for the treatment-related ICE.
+Estimation can then proceed using parametric or semiparametric methods, such as the Cox proportional hazards model (Cox,, 1972), or more flexible approaches (Wolock et al.,, 2024).
+A similar approach can be used to estimate $G_{1}(t\mid X)$ by switching the roles of $T$ and $C$.
+The corresponding functions $S_{0}(t\mid X)$ and $G_{0}(t\mid X)$ can be estimated in the same way using data from the control group.
 3.3 Two basic estimators
-Identification formulas (2) and (3) in Theorem 1 motivate two basic estimators based on estimating the corresponding nuisance functions.
-Let $\hat{S}_{a}(t\mid X)$ and $\hat{G}_{a}(t\mid X)$ denote the fitted models for $t\leq k$ and $a=0,1$, $\hat{\mu}_{a}(X)$ denote the fitted outcome models for $a=0,1$, and $\hat{e}(X)$ denote the fitted propensity score model.
-Plugging in the fitted values of the nuisance parameters and taking empirical analogs of equations (2) and (3) motivate the following two estimators, respectively,
+The identification formulas (2) and (3) in Theorem 1 motivate two basic estimators, each relying on estimates of the corresponding nuisance functions.
+Let $\hat{S}_{a}(t\mid X)$ and $\hat{G}_{a}(t\mid X)$ denote the fitted survival models for treatment-related and treatment-unrelated ICEs, respectively, for $t\leq k$ and $a=0,1$.
+Let $\hat{\mu}_{a}(X)$ be the fitted outcome regression model for each treatment arm, and $\hat{e}(X)$ be the fitted propensity score model.
+By substituting these fitted values into the identification formulas and taking empirical analogs, we obtain the following two estimators:
 $\hat{\tau}^{\textup{out}}$ $=$ $n^{-1}\sum_{i=1}^{n}\hat{\mu}_{1}(X_{i})\hat{S}_{1}(k\mid X_{i})-n^{-1}\sum_{i=1}^{n}\hat{\mu}_{0}(X_{i})\hat{S}_{0}(k\mid X_{i}),$
 $\hat{\tau}^{\textup{ipw}}$ $=$ $n^{-1}\sum_{i=1}^{n}\frac{A_{i}Y_{i}1(T_{i}\wedge C_{i}>k)}{\hat{e}(X_{i})\hat{G}_{1}(k\mid X_{i})}-n^{-1}\sum_{i=1}^{n}\frac{(1-A_{i})Y_{i}1(T_{i}\wedge C_{i}>k)}{\left\{1-\hat{e}(X_{i})\right\}\hat{G}_{0}(k\mid X_{i})}.$
-The outcome regression estimator $\hat{\tau}^{\textup{out}}$ is consistent if the outcome model and the survival function are correctly specified.
-The weighting estimator $\hat{\tau}^{\textup{ipw}}$ is consistent if the propensity score model and the censoring mechanism are correctly specified.
-3.4 Propensity score-augmented weighting estimator
-Combining the two identification formulas (2) and (3), we have the following augmented weighting identification formula,
+The outcome regression estimator $\hat{\tau}^{\textup{out}}$ is consistent if both $\hat{\mu}_{a}(X)$ and $\hat{S}_{a}(k\mid X)$ are consistently estimated.
+The inverse probability weighting estimator $\hat{\tau}^{\textup{ipw}}$ is consistent if both $\hat{e}(X)$ and $\hat{G}_{a}(k\mid X)$ are consistently estimated.
+3.4 An augmented estimator
+By combining the two identification formulas (2) and (3), we derive the following augmented identification formula:
 $\tau$ $=$ $E\left\{\frac{AY1(T\wedge C>k)}{e(X)G_{1}(k\mid X)}-\frac{A-e(X)}{e(X)}\mu_{1}(X)S_{1}(k\mid X)\right\}$ (4)
 $-E\left[\frac{(1-A)Y1(T\wedge C>k)}{\left\{1-e(X)\right\}G_{0}(k\mid X)}-\frac{e(X)-A}{1-e(X)}\mu_{0}(X)S_{0}(k\mid X)\right].$
+Equation (4) can be interpreted both as a modified form of the outcome regression identification formula (2), and as an augmented form of the weighting identification formula in (3).
+Under the true model at the population level, the correction terms have mean 00, and thus (4) holds by construction.
+However, it becomes meaningful in the presence of possible model misspecification, where the augmentation plays a crucial role in improving robustness.
 Based on (4), we construct an estimator that augments the weighting estimator $\hat{\tau}^{\textup{ipw}}$ using the estimated outcome models,
@@
 Suppose Assumptions 1 and 2 hold, and assume that $G_{a}(k\mid X)$ is correctly specified for $a=0,1$.
-$\hat{\tau}^{\textup{aug}}$ is a consistent estimator for $\tau$ if either the propensity score model $e(X)$ is correct, or both the outcome regression $\mu_{a}(X)$ and the survival models $S_{a}(k\mid X)$ are correct for $a=0,1$.
-Proposition 1 shows that the consistency of $\hat{\tau}^{\textup{aug}}$ depends on the correct model of $G_{a}(k\mid X)$ for $a=0,1$ but remains robust to misspecification of the propensity score model $e(X)$.
-To build intuition, Equation (4) can be viewed as a projection of the weighting identification formula (3) onto the nuisance tangent space of the propensity score model $A\mid X$, thus we refer to $\hat{\tau}^{\textup{aug}}$ as a propensity score-augmented estimator.
-However, since it is not further projected onto the tangent space of the $G_{a}(k\mid X)$ model, the estimator does not retain robustness to misspecification of $G_{a}(k\mid X)$.
-We will further address the issue in Section 4, where we introduce an estimator that is robust to the misspecification of both models.
-$\hat{\tau}^{\textup{aug}}$ dominates the weighting estimator $\hat{\tau}^{\textup{ipw}}$ in terms of robustness.
+$\hat{\tau}^{\textup{aug}}$ is a consistent estimator for $\tau$ if either $e(X)$ is correct, or both $\mu_{a}(X)$ and $S_{a}(k\mid X)$ are correct for $a=0,1$.
+Proposition 1 shows that the consistency of $\hat{\tau}^{\textup{aug}}$ depends on correctly specifying $G_{a}(k\mid X)$ for $a=0,1$, while it remains robust to misspecification of the propensity score $e(X)$.
+To build intuition, from the perspective of semiparametric efficiency theory, Equation (4) can be viewed as a projection of the weighting identification formula (3) onto the nuisance tangent space of the propensity score model $A\mid X$.
+However, because it is not further projected onto the nuisance tangent space of the $G_{a}(k\mid X)$ model, the resulting estimator does not retain robustness to misspecification of $G_{a}(k\mid X)$.
+We will further address the issue in Section 4, where we introduce an alternative estimator that is robust to the misspecification of both models.
+The augmented estimator $\hat{\tau}^{\textup{aug}}$ dominates the weighting estimator $\hat{\tau}^{\textup{ipw}}$ in terms of robustness.
 The consistency of $\hat{\tau}^{\textup{aug}}$ is contingent on less restrictive requirements on nuisance parameter estimation in the sense that $\hat{\tau}^{\textup{aug}}$ is consistent whenever $\hat{\tau}^{\textup{ipw}}$ is.
-Consistency of $\hat{\tau}^{\textup{aug}}$ requires the correct specification of the censoring mechanism $G_{a}(k\mid X)$ for $a=0,1$.
-Given that $G_{a}(k\mid X)$ is correctly modeled, $\hat{\tau}^{\textup{aug}}$ is doubly robust because it is consistent if either the propensity score model is correct, or the outcome regression and the survival models are both correct.
+Consistency of $\hat{\tau}^{\textup{aug}}$ requires the correct specification of the survival function for treatment-unrelated ICE $G_{a}(k\mid X)$ for $a=0,1$.
+Given that $G_{a}(k\mid X)$ is correctly modeled, $\hat{\tau}^{\textup{aug}}$ is doubly robust because it is consistent if either the propensity score model is correct, or the outcome regression and the survival model for treatment-related ICEs are both correct.
 There is no clear dominance between $\hat{\tau}^{\textup{aug}}$ and $\hat{\tau}^{\textup{out}}$ as they require consistent estimation of different sets of nuisance parameters.
 For all three estimators $\hat{\tau}^{\textup{out}}$, $\hat{\tau}^{\textup{ipw}}$, and $\hat{\tau}^{\textup{aug}}$, we can construct variance estimators using nonparametric bootstrap.
-4 Another doubly robust and semiparametrically efficient estimator
-The propensity score-augmented weighting estimator $\hat{\tau}^{\textup{aug}}$ improves robustness.
-However, it is not robust to the misspecification of the censoring mechanism, nor does it achieve the semiparametric efficiency bound.
+4 A semiparametrically efficient and doubly robust estimator
+The augmented weighting estimator $\hat{\tau}^{\textup{aug}}$ improves robustness of $\hat{\tau}^{\textup{ipw}}$.
+However, it is not robust to the misspecification of the survival function for treatment-unrelated ICE, nor does it achieve the semiparametric efficiency bound.
 In this section, we show that $\hat{\tau}^{\textup{aug}}$ can be further improved.
@@
 First, for a given treatment arm $a=0,1$, the treatment-unrelated ICE time $C(a)$ censors the full data because $Y^{\textup{cp}}(a)$ is only observable if $C(a)\geq\{T(a)\wedge k\}$.
-Within each treatment group, we do not observe the full data due to censoring and only observe $(\Delta(a)=1\{C(a)\geq T(a)\wedge k\},\tilde{T}(a)=C(a)\wedge T(a)\wedge k,\Delta(a)Y^{\textup{cp}}(a))$.
+Within each treatment group, we do not observe the full data due to censoring and only observe
+$(\Delta(a)=1\{C(a)\geq T(a)\wedge k\},\tilde{T}(a)=C(a)\wedge T(a)\wedge k,\Delta(a)Y^{\textup{cp}}(a)).$
 Second, the treatment assignment generates another level of missingness, because, for each observation, we never simultaneously observe both composite potential outcomes $\{Y^{\textup{cp}}(1),Y^{\textup{cp}}(0)\}$ even without censoring.
-Therefore, the observed data is $O=(A,\Delta=\Delta(A),\tilde{T}=\tilde{T}(A),\Delta Y^{\textup{cp}}=\Delta(A)Y^{\textup{cp}}(A))$.
-4.1 EIF and estimator
+Therefore, the observed data is
+$O=(A,\Delta=\Delta(A),\tilde{T}=\tilde{T}(A),\Delta Y^{\textup{cp}}=\Delta(A)Y^{\textup{cp}}(A)).$
+4.1 EIF and EIF-based estimator
 In the following theorem, we provide the EIF for $\tau$.
@@
 $-\frac{e(X)-A}{1-e(X)}\mu_{0}(X)S_{0}(k\mid X)-\mu_{0},$
-and thus, the EIF for $\tau$ is $D_{\tau}(O)=D_{1}(O)-D_{0}(O)$, where $\textnormal{d}M_{G_{a}}(t)=1(C\in\textnormal{d}t,\Delta=0)-1(\tilde{T}\geq t)\textnormal{d}\Lambda_{a}(t\mid X)$ with $\Lambda_{a}(t\mid X)$ denoting the conditional cumulative hazard function for the censoring $C$ in the treatment group $A=a$ for $a=0,1$.
+and thus, the EIF for $\tau$ is $D_{\tau}(O)=D_{1}(O)-D_{0}(O)$, where $\textnormal{d}M_{G_{a}}(t)=1(C\in\textnormal{d}t,\Delta=0)-1(\tilde{T}\geq t)\textnormal{d}\Lambda_{a}(t\mid X)$ with $\Lambda_{a}(t\mid X)$ denoting the conditional cumulative hazard function for the treatment-unrelated ICE $C$ in the treatment group $A=a$ for $a=0,1$.
 The $M_{G_{a}}(t)$ in the EIF is the martingale constructed from the censoring counting process.
 Intuitively, $1(C\in\textnormal{d}t,\Delta=0)$ is the actual observed increment in the censoring counting process at time $t$, which records whether a censoring event has occurred, while $1(\tilde{T}\geq t)\textnormal{d}\Lambda_{a}(t\mid X)$ represents the expected increment in the counting process, given the history up to time $t$.
-The martingale $M_{G_{a}}(t)$ is essentially capturing the difference between the actual observed events and their expected occurrences.
+The martingale $M_{G_{a}}(t)$ captures the difference between the actual observed events and their expected occurrences.
 The EIF implies another identification formula for $\tau$ by the property that $E\{D_{\tau}(O)\}=0$.
@@
 $-E\left[\frac{1-A}{1-e(X)}\mu_{0}(X)S_{0}(k\mid X)\int_{0}^{\tilde{T}}\frac{\textnormal{d}M_{G_{0}}(t)}{S_{0}(t\mid X)G_{0}(t\mid X)}\right],$
-where the first two lines are the same as in the propensity score-augmented inverse probability weighting identification formula (4).
-Intuitively, we further augment (4) by the term (7) to achieve robustness to misspecification of the censoring model.
-The EIFs in Theorem 2 motivate the following estimators for $\tau$,
+where the first two lines are the same as in the augmented weighting identification formula (4).
+Intuitively, we further augment (4) by the last two terms in (7) to achieve robustness to misspecification of the survival function for treatment-unrelated ICE.
+These augmentation terms are zero at the population level under true $G_{a}(t\mid X)$.
+However, their empirical counterparts may deviate significantly from zero, providing diagnostic insight into possible misspecification of $G_{a}(t\mid X)$.
+The EIFs in Theorem 2 motivate the following estimator for $\tau$,
 $\hat{\tau}^{\textup{eif}}$ $=$ $\hat{\tau}^{\textup{aug}}+n^{-1}\sum_{i=1}^{n}\frac{A_{i}}{\hat{e}(X_{i})}\hat{\mu}_{1}(X_{i})\hat{S}_{1}(k\mid X_{i})\int_{0}^{\tilde{T}_{i}}\frac{\textnormal{d}M_{\hat{G}_{1}}(t)}{\hat{S}_{1}(t\mid X_{i})\hat{G}_{1}(t\mid X_{i})}$
 $-n^{-1}\sum_{i=1}^{n}\frac{1-A_{i}}{1-\hat{e}(X_{i})}\hat{\mu}_{0}(X_{i})\hat{S}_{0}(k\mid X_{i})\int_{0}^{\tilde{T}_{i}}\frac{\textnormal{d}M_{\hat{G}_{0}}(t)}{\hat{S}_{0}(t\mid X_{i})\hat{G}_{0}(t\mid X_{i})}.$
-More explicitly, taking the first modification term in $\hat{\tau}^{\textup{eif}}$ as an example, because the observed time points are discrete, the estimator of the integration part for observation $i$ is
+To gain insights, consider the first augmentation term in $\hat{\tau}^{\textup{eif}}$. Since the observed time points are discrete, the integral for observation $i$ can be estimated as follows:
 $\int_{0}^{\tilde{T}_{i}}\frac{\textnormal{d}M_{\hat{G}_{1}}(t)}{\hat{S}_{1}(t\mid X_{i})\hat{G}_{1}(t\mid X_{i})}$ $=$ $\sum_{t\leq\tilde{T}_{i}}\frac{1(\Delta_{i}=0,C_{i}=t)-\hat{\lambda}_{1}(t\mid X_{i})}{\hat{S}_{1}(t\mid X_{i})\hat{G}_{1}(t\mid X_{i})}$ (8)
 $=$ $-\sum_{t\leq\tilde{T}_{i}}\frac{\hat{\lambda}_{1}(t\mid X_{i})}{\hat{S}_{1}(t\mid X_{i})\hat{G}_{1}(t\mid X_{i})}+\frac{1(\Delta_{i}=0)}{\hat{S}_{1}(\tilde{T}_{i}\mid X_{i})\hat{G}_{1}(\tilde{T}_{i}\mid X_{i})},$
-where $\hat{\lambda}_{1}(t\mid X_{i})$ denotes the estimator of the conditional hazard function of censoring $\lambda_{1}(t\mid X)$.
-The first term in (8) is a summation of the ratio $-\hat{\lambda}_{1}(t\mid X_{i})/\{\hat{S}_{1}(t\mid X_{i})\hat{G}_{1}(t\mid X_{i})\}$, where the summation is over all observed event time points before $\tilde{T}_{i}$.
-The second term is 0 for observations that are not right-censored by the treatment-unrelated ICE, and is $1/\{\hat{S}_{1}(\tilde{T}_{i}\mid X_{i})\hat{G}_{1}(\tilde{T}_{i}\mid X_{i})\}$ for observations with a treatment-unrelated ICE happened at time $\tilde{T}_{i}$.
-To construct the estimator $\hat{\tau}^{\textup{eif}}$, we need to estimate the following nuisance parameters: the propensity score model $e(X)$, the outcome model $\mu_{a}(X)$ for $a=0,1$, the survival function $S_{a}(t\mid X)$, and the censoring mechanism $G_{a}(t\mid X)$ for $t\leq k$ and $a=0,1$.
-Nonetheless, the consistency of $\hat{\tau}^{\textup{eif}}$ does not require the correct specification of all four nuisance parameters, as shown in Theorem 3.
+where $\hat{\lambda}_{1}(t\mid X_{i})$ denotes the estimated conditional hazard function for the treatment-unrelated ICE at time $t$ given covariates $X_{i}$.
+In (8), the first term is a sum of the ratio $-\hat{\lambda}_{1}(t\mid X_{i})/\{\hat{S}_{1}(t\mid X_{i})\hat{G}_{1}(t\mid X_{i})\}$, where the summation is over all observed event time points before $\tilde{T}_{i}$.
+The second term is 0 for individuals who are not right-censored by the treatment-unrelated ICE, and is $1/\{\hat{S}_{1}(\tilde{T}_{i}\mid X_{i})\hat{G}_{1}(\tilde{T}_{i}\mid X_{i})\}$ for observations with a treatment-unrelated ICE occurred at time $\tilde{T}_{i}$.
+To construct the estimator $\hat{\tau}^{\textup{eif}}$, we need to estimate the following nuisance parameters: the propensity score model $e(X)$, the outcome model $\mu_{a}(X)$ for $a=0,1$, and the survival functions for treatment-related and treatment-unrelated ICEs $S_{a}(t\mid X)$ and $G_{a}(t\mid X)$, respectively, for $t\leq k$ and $a=0,1$.
+Importantly, as shown in Theorem 3, the consistency of $\hat{\tau}^{\textup{eif}}$ does not require correct specification of all four nuisance parameters.
 4.2 Asymptotic properties of the EIF-based estimator
@@
 We first introduce additional notation.
-For $t\leq k$ and $a=0,1$, let $e^{*}$, $G_{a}^{*}$, $\mu_{a}^{*}$, and $S_{a}^{*}$ denote the probability limit of the estimated propensity score, estimated censoring mechanism, estimated outcome model, and estimated survival function of the AE time, respectively, i.e., $\lVert\hat{e}-e^{*}\rVert=o_{P}(1)$, $\lVert\hat{G}_{a}-G_{a}^{*}\rVert=o_{P}(1)$, $\lVert\hat{\mu}_{a}-\mu_{a}^{*}\rVert=o_{P}(1)$, and $\lVert\hat{S}_{a}-S_{a}^{*}\rVert=o_{P}(1)$.
-If the propensity score model is consistently estimated, we have $e^{*}=e$, and analogous results for the other three nuisance models.
-In the following theorem, we provide the double robustness of the EIF estimator $\hat{\tau}^{\textup{eif}}$.
+For $t\leq k$ and $a=0,1$, let $e^{*}$, $G_{a}^{*}$, $\mu_{a}^{*}$, and $S_{a}^{*}$ denote the probability limit of the estimated nuisance functions $\hat{e}$, $\hat{G}_{a}$, $\hat{\mu}_{a}$, $\hat{S}_{a}$, respectively, i.e., $\lVert\hat{e}-e^{*}\rVert=o_{P}(1)$, $\lVert\hat{G}_{a}-G_{a}^{*}\rVert=o_{P}(1)$, $\lVert\hat{\mu}_{a}-\mu_{a}^{*}\rVert=o_{P}(1)$, and $\lVert\hat{S}_{a}-S_{a}^{*}\rVert=o_{P}(1)$.
+If a given nuisance model is correctly specified, its corresponding limit equals the true function.
+For example, if the propensity score model is consistent, then $e^{*}=e$, and similar results hold for the other three nuisance components.
+In the following theorem, we provide the double robustness of the EIF-based estimator $\hat{\tau}^{\textup{eif}}$.
 Theorem 3 (Double robustness of $\hat{\tau}^{\textup{eif}}$)
 Under Assumptions 1 and 2, $\hat{\tau}^{\textup{eif}}$ is doubly robust in the sense that it is consistent for $\tau$ if either $\{e^{*}(X)=e(X),G_{a}^{*}(t\mid X)=G_{a}(t\mid X)\}$ or $\{\mu_{a}^{*}(X)=\mu_{a}(X),S_{a}^{*}(t\mid X)=S_{a}(t\mid X)\}$ for $t\leq k$ and $a=0,1$.
-Theorem 3 shows that $\hat{\tau}^{\textup{eif}}$ is consistent if at least one of the following two sets of nuisance parameters are consistently estimated: (1) the outcome model $\mu_{a}(X)$ and the survival function $S_{a}(t\mid X)$; (2) the propensity score model $e(X)$ and the censoring mechanism $G_{a}(t\mid X)$, for $t\leq k$ and $a=0,1$.
-When our data is from an RCT, i.e., the propensity score $e(X)$ is known, and thus $\hat{e}(X)$ is always correctly specified, the EIF estimator $\hat{\tau}^{\textup{eif}}$ is a consistent estimator for $\tau$ if either the censoring mechanism $G_{a}(t\mid X)$ is consistently estimated, or both the outcome model $\mu_{a}(X)$ and the survival function $S_{a}(t\mid X)$ are consistently estimated, for $t\leq k$ and $a=0,1$.
+Theorem 3 shows that $\hat{\tau}^{\textup{eif}}$ is consistent if at least one of the following two sets of nuisance parameters are consistently estimated: (1) the outcome model $\mu_{a}(X)$ and the survival function for the treatment-related ICE $S_{a}(t\mid X)$; (2) the propensity score model $e(X)$ and the survival function for the treatment-unrelated ICE $G_{a}(t\mid X)$, for $t\leq k$ and $a=0,1$.
+When our data is from an RCT, i.e., the propensity score $e(X)$ is known by design, so $\hat{e}(X)$ can be correctly specified.
+Consequently, the EIF-based estimator $\hat{\tau}^{\textup{eif}}$ is a consistent estimator for $\tau$ if either $G_{a}(t\mid X)$ is consistently estimated, or both $\mu_{a}(X)$ and $S_{a}(t\mid X)$ are consistently estimated, for $t\leq k$ and $a=0,1$.
 To conduct statistical inference, we provide the asymptotic distribution of $\hat{\tau}^{\textup{eif}}$.
@@
 Assumption 3 (Consistency of the nuisance parameters)
-Assume that either $\{e^{*}(X)=e(X),\allowbreak G_{a}^{*}(t\mid X)=G_{a}(t\mid X)\}$ or $\{\mu_{a}^{*}(X)=\mu_{a}(X),S_{a}^{*}(t\mid X)=S_{a}(t\mid X)\}$ for $t\leq k$ and $a=0,1$ is satisfied.
+Assume that either $\{e^{*}(X)=e(X),G_{a}^{*}(t\mid X)=G_{a}(t\mid X)\}$ or $\{\mu_{a}^{*}(X)=\mu_{a}(X),S_{a}^{*}(t\mid X)=S_{a}(t\mid X)\}$ for $t\leq k$ and $a=0,1$ is satisfied.
 Assumption 4 (Donsker Condition)
@@
 Assumption 5 (Convergence rates of nuisance parameters)
-The convergence rate of the nuisance parameters estimation satisfy $\{||\hat{e}-e^{*}||+||\hat{G}_{a}-G_{a}^{*}||\}\{||\hat{\mu}_{a}-\mu_{a}^{*}||+||\hat{S}_{a}-S_{a}^{*}||\}=o_{P}(n^{-1/2})$ for $a=0,1$.
-Assumption 3 requires either the propensity score and censoring models or the survival and outcome models to be consistently estimated.
+The convergence rate of the nuisance parameters estimation satisfies $\{||\hat{e}-e^{*}||+||\hat{G}_{a}-G_{a}^{*}||\}\{||\hat{\mu}_{a}-\mu_{a}^{*}||+||\hat{S}_{a}-S_{a}^{*}||\}=o_{P}(n^{-1/2})$ for $a=0,1$.
+Assumption 3 requires either the propensity score and survival function of the treatment-unrelated ICE or the outcome and survival function of the treatment-related ICE to be consistently estimated.
 It guarantees the consistency of $\hat{\tau}^{\textup{eif}}$.
 Assumption 4 imposes restrictions on the nuisance model complexity and is a standard regularity condition (Van der Vaart,, 2000).
-Recent literature documents that we can use cross-fitting in the estimation of nuisance parameters to relax the Donsker condition (van der Laan et al.,, 2011; Chernozhukov et al.,, 2018; Díaz,, 2019).
-Assumption 5 imposes additional restrictions on the rate of convergence of the nuisance parameters, in addition to all of them being consistently estimated.
+We can also employ flexible machine learning models with cross-fitting techniques in the estimation of nuisance parameters to relax the Donsker condition (Pfanzagl and Wefelmeyer,, 1985; Klaassen,, 1987; Zheng and van der Laan,, 2011; Chernozhukov et al.,, 2018).
+Finally, Assumption 5 imposes additional restrictions on the rate of convergence of the nuisance parameters, in addition to all of them being consistently estimated.
 Theorem 4 (Asymptotic distribution)
@@
 $n^{1/2}(\hat{\tau}^{\textup{eif}}-\tau)\ =\ n^{-1/2}\sum_{i=1}^{n}D_{\tau}(O_{i})+o_{P}(1).$
-$\hat{\tau}^{\textup{eif}}$ is a consistent and asymptotically Gaussian estimator of $\tau$ with asymptotic variance equal to $E\{D_{\tau}^{2}(O)\}$, thus achieves the semiparametric efficiency bound.
+The estimator $\hat{\tau}^{\textup{eif}}$ is a consistent and asymptotically Normal estimator of $\tau$ with asymptotic variance equal to $E\{D_{\tau}^{2}(O)\}$, thus achieving the semiparametric efficiency bound.
 Therefore, we can construct the variance estimator based on the semiparametric efficiency bound by taking the empirical analog of the plug-in estimation $n^{-1}\sum_{i=1}^{n}\hat{D}_{\tau}^{2}(O_{i})$.
 We can also use a nonparametric bootstrap to estimate variance and conduct statistical inference if the nuisance estimators satisfy certain smoothness conditions.
-Compare $\hat{\tau}^{\textup{eif}}$ with the propensity score-augmented weighting estimator $\hat{\tau}^{\textup{aug}}$.
-On the one hand, $\hat{\tau}^{\textup{eif}}$ is more robust in the sense that it allows for misspecification of the censoring model and is asymptotically more efficient.
+4.3 Comparing the estimators
+In real data analysis, we recommend implementing and comparing all four proposed estimators to assess whether they provide coherent scientific implications.
+Theoretical comparisons among these estimators primarily focus on two key dimensions: robustness to nuisance model misspecification and asymptotic efficiency.
+In practice, examining how the point estimates differ across these estimators can provide insight into the presence of nuisance model misspecification.
+Specifically, certain pairwise differences among the estimators can serve as informal diagnostics.
+For example, if $e(X)$ is correctly specified, the difference between $\hat{\tau}^{\textup{ipw}}$ and $\hat{\tau}^{\textup{aug}}$ should converge to zero.
+Similarly, if $\mu_{a}(X)$ and $S_{a}(k\mid X)$ are correctly specified, the difference between $\hat{\tau}^{\textup{out}}$ and $\hat{\tau}^{\textup{aug}}$ should converge to zero.
+Finally, if $G_{a}(t\mid X)$ is correctly specified, the difference between $\hat{\tau}^{\textup{eif}}$ and $\hat{\tau}^{\textup{aug}}$ should converge to zero.
+Therefore, when implementing all four estimators in practice, the observed difference among them may serve as an informal diagnostic tool to detect possible nuisance model misspecification.
+We provide more details in Section A.1 in the supplementary material.
+Among the four estimators, a key comparison lies between the EIF-based estimator $\hat{\tau}^{\textup{eif}}$ with the augmented weighting estimator $\hat{\tau}^{\textup{aug}}$.
+On the one hand, $\hat{\tau}^{\textup{eif}}$ is more robust and asymptotically efficient: it allows for misspecification of the censoring model $C\mid A,X$ and it achieves the semiparametric efficiency bound under correct nuisance specification.
 Continuing the intuition discussed in Section 3.4, $\hat{\tau}^{\textup{aug}}$ only orthogonalizes the part corresponding to $A\mid X$ but not $C\mid X,A$, making it robust only to misspecification of the $A\mid X$ model, not the $C\mid X,A$ model.
@@
 We consider the choice between $\hat{\tau}^{\textup{eif}}$ and $\hat{\tau}^{\textup{aug}}$ as a trade-off between implementation simplicity and robustness and efficiency of the estimators.
-5 Simulation
+Finally, we note the theoretical possibility of applying a martingale-based correction to the basic weighting estimator $\hat{\tau}^{\textup{ipw}}$, in a manner analogous to the correction in the EIF-based estimator $\hat{\tau}^{\textup{eif}}$ compared with $\hat{\tau}^{\textup{aug}}$.
+Such a corrected estimator would be robust to misspecification of $G_{a}(t\mid X)$.
+However, it remains less attractive in practice: it is neither robust to misspecification of the propensity score nor asymptotically efficient, and it lacks the implementation simplicity of $\hat{\tau}^{\textup{aug}}$.
+For these reasons, we do not recommend it, though we include its explicit form in Section A.2 of the supplementary material only for theoretical completeness.
+4.4 Simulation results
+We conduct Monte Carlo simulations to evaluate the finite-sample performance of four estimators under various combinations of correctly specified and misspecified nuisance models.
+Consistent with our theory, $\hat{\tau}^{\textup{eif}}$ demonstrates the greatest robustness across all misspecification scenarios, maintaining low bias and valid coverage even when multiple nuisance models are incorrect.
+We relegate the detailed results to Section B in the supplementary material.
+5 Analyzing two phase-3 trials
+5.1 Data analysis
+In this section, we re-analyze data from two trials on systemic lupus erythematosus (Morand et al.,, 2023; Petri et al.,, 2023).
+Both trials are double-blinded, randomized, placebo-controlled phase-3 trials.
+There are three treatment arms: 2mg baritinib, 4mg baritinib, and placebo.
+The primary outcome in both studies is the Systemic lupus erythematosus Responder Index 4 (SRI4) at week 52, a binary composite endpoint that reflects clinical response.
+A participant is classified as a responder if they show a predefined improvement in disease activity, without overall worsening or the emergence of significant disease activity in new organ systems.
+Baseline covariates used for model fitting include geographic region, baseline corticosteroid use, and the Physician’s Global Assessment score measured at baseline.
+A substantial proportion of participants in both trials experienced ICEs: 218 out of 760 in one trial and 211 out of 775 in the other.
+As described in Section 1.2, we classify these ICEs into two categories: treatment-related ICEs (82.6% and 84.4% in the two trials, respectively) and treatment-unrelated ICEs (17.4% and 15.6%).
+We conduct separate comparisons of the 2mg and 4mg baricitinib treatment arms versus the placebo group.
+For each comparison, we apply all four estimators: $\hat{\tau}^{\textup{out}}$, $\hat{\tau}^{\textup{ipw}}$, $\hat{\tau}^{\textup{aug}}$, and $\hat{\tau}^{\textup{eif}}$.
+We use logistic regression to estimate the propensity score and outcome models, and use Cox proportional hazards regression to separately estimate the survival functions for the two types of ICEs.
+We conduct variance estimation and inference using a nonparametric bootstrap with 500 replicates.
+Table 2 presents the resulting point estimates, their estimated standard errors, and corresponding $p$-values.
+Table 2: Treatment effect of Baricitinib on the primary outcome SRI4.
+The first four columns report the results of Trial 1 and the last four columns report those of Trial 2.
+The first three rows report the point estimators, standard errors, and the $p$-value of the composite outcome average treatment effect corresponding to the 2mg treatment arm, and the last three rows report those of the 4mg treatment arm.
+For each trial and each treatment arm, we report the results based on all four estimators.
+Trial 1 (Petri et al.,, 2023) Trial 2 (Morand et al.,, 2023)
+$\hat{\tau}^{\textup{out}}$ $\hat{\tau}^{\textup{ipw}}$ $\hat{\tau}^{\textup{aug}}$ $\hat{\tau}^{\textup{eif}}$ $\hat{\tau}^{\textup{out}}$ $\hat{\tau}^{\textup{ipw}}$ $\hat{\tau}^{\textup{aug}}$ $\hat{\tau}^{\textup{eif}}$
+2mg point 0.030 0.029 0.026 0.026 0.019 0.019 0.022 0.022
+se 0.042 0.043 0.042 0.043 0.042 0.043 0.043 0.042
+$p$-value 0.479 0.504 0.534 0.540 0.643 0.662 0.602 0.606
+4mg point 0.113 0.120 0.115 0.113 $-0.002$ $-0.002$ $-0.002$ $-0.002$
+se 0.046 0.046 0.046 0.046 0.042 0.042 0.042 0.042
+$p$-value 0.013 0.008 0.012 0.013 0.961 0.962 0.966 0.969
+5.2 Interpretation of the statistical results
+Across both trials and treatment arms, the point estimates and $p$-values from the four methods are generally consistent.
+Our results suggest a statistically significant positive effect of 4mg baricitinib on the SRI4 outcome in Trial 1, while in Trial 2, the effect is not significant and the point estimate is slightly negative.
+No significant treatment effect is observed for the 2 mg baricitinib dose in either trial.
+As discussed in Section 4.3, though not a formal statistical test, qualitatively, the robustness of estimates across the four estimators in Table 2 provides empirical evidence that our nuisance models may not be severely misspecified.
+5.3 Comparison with ad-hoc methods
+Next, we use the two commonly used ad-hoc methods and compare them with our proposed methods.
+The first method is the non-responder imputation method that assigns an outcome as 00 whenever an ICE occurs.
+It is equivalent to using the composite outcome strategy that treats all ICEs as treatment-related, leading to a composite outcome equal to 00.
+The second method naively applies a hypothetical strategy to all ICEs, assuming they are independent of potential outcomes conditional on observed covariates.
+Table 3 reports the estimated treatment effect of both treatment arms across the two trials using the inverse probability weighting estimator.
+The non-responder imputation method generally yields slightly smaller effect estimates than our proposed methods, although the differences are modest, likely due to the relatively low proportion of treatment-unrelated ICEs in both trials.
+Nevertheless, our proposed approach targets a clinically more meaningful causal estimand.
+The second ad hoc method naively applies the hypothetical strategy to all ICEs, assuming that their occurrence is conditionally independent of the potential outcomes given baseline covariates.
+This assumption is untestable and likely violated in practice.
+For example, treatment discontinuation due to adverse events or lack of efficacy is plausibly related to a patient’s potential outcome.
+As shown in Table 3, this method produces treatment effect estimates that differ substantially from those obtained using our proposed approach, potentially leading to misleading clinical conclusions.
+These empirical findings are consistent with the large-sample bias observed in the simulation study and underscore the importance of using appropriate strategies tailored to different types of ICEs.
+Table 3: Treatment effect of Baricitinib on the primary outcome SRI4 using non-responder imputation (NRI) and hypothetical strategy (HS).
+The first two columns report the results of Trial 1 and the last two columns report those of Trial 2.
+The first three rows report the point estimators, standard errors, and the $p$-value of the estimated treatment effect corresponding to the 2mg treatment arm, and the last three rows report those of the 4mg treatment arm.
+For each trial and each treatment arm, we report the results based on the inverse probability-weighting estimator.
+Trial 1 Trial 2
+NRI HS NRI HS
+2mg point 0.039 0.011 0.010 0.022
+se 0.043 0.049 0.042 0.050
+$p$-value 0.362 0.815 0.817 0.654
+4mg point 0.093 0.073 $-0.008$ 0.044
+se 0.044 0.050 0.041 0.051
+$p$-value 0.036 0.147 0.847 0.382
+6 Future directions
+In this study, we address the challenges posed by ICEs in RCTs by proposing methods to handle competing ICEs.
+We classify ICEs into treatment-related and treatment-unrelated events and apply different strategies to identify a clinically meaningful causal effect.
+For treatment-related ICEs, which are often informative about a patient’s outcome, we use a composite variable strategy that assigns an outcome value indicative of treatment failure.
+For treatment-unrelated ICEs, we apply a hypothetical strategy, assuming their timing is conditionally independent of the outcome given treatment and baseline covariates, and envisioning a scenario in which such events do not occur.
+The central thesis of this paper is to address the challenge of competing ICEs, where the first ICE censors all subsequent ones.
+In this paper, we propose a principled framework that carefully formulates the estimand, establishes its nonparametric identification and semiparametric estimation theory, and introduces weighting, outcome regression, and doubly robust estimators.
+While our proposed framework provides a rigorous and flexible approach for handling competing ICEs in RCTs, several challenges and extensions remain open for future research.
+6.1 Data collection and ICE classification
+Our proposed methods have broad applicability across various therapeutic areas, including immunology, oncology, and cardiology, where treatment discontinuation and other ICEs frequently occur.
+However, the effectiveness of these approaches relies on the accurate classification of ICEs, which should be performed in collaboration with clinicians to ensure clinical relevance in trial analyses.
+To support this, a modernized case report form is needed to enable more granular documentation of the timing, reasons, and magnitude of treatment discontinuation.
+For example, the reason for the event should be specified, such as discontinuation due to toxicity versus lack of efficacy.
+In some cases, the event may need to meet a threshold of magnitude, such as the use of additional medication exceeding a specified duration or dose.
+Additionally, the timing of the event may be relevant, particularly in relation to its proximity to outcome assessment.
+Fortunately, a cross-industry PHUSE working group is tackling this problem, and a recommended new case report form will be available soon (PHUSE Working Group,, 2024).
+6.2 Random $K$
+In practice, the measurement of the outcome of interest $Y$ does not necessarily happen at a fixed point $k$.
+For instance, the time $K$ that a patient visits the clinic and takes the measurement can be treated as a random variable independent of treatment $A$, potential outcomes $Y(a)$, and both types of ICEs $T(a)$ and $C(a)$ for $a=0,1$.
+For a random $K$, the analogous formulas to both the previous identification formulas (2) and (3) hold, with a replacement of $k$ to $K$, i.e.,
+$\tau$ $=$ $E\left\{\mu_{1}(X)S_{1}(K\mid X)-\mu_{0}(X)S_{0}(K\mid X)\right\}$ (9)
+$=$ $E\left[\frac{AY1(T\wedge C>K)}{e(X)G_{1}(K\mid X)}-\frac{(1-A)Y1(T\wedge C>K)}{\left\{1-e(X)\right\}G_{0}(K\mid X)}\right].$ (10)
+If $K$ is a pre-treatment covariate that is observed, the proposed four estimators carry over by replacing the fixed $k$ with the observed values of $K$.
+If $K$ is also treated as a post-treatment variable, we can construct a weighting estimator following (10).
+However, (9) no longer provides a feasible identification formula unless $S_{a}(K\mid X)$ is identified for $a=0,1$.
+Considering the presence of three competing time-to-event random variables, it becomes necessary to employ competing risks models.
+We defer this analysis to future work.
+Acknowledgement
+Sizhu Lu and Peng Ding were partially supported by the U.S.
+National Science Foundation grants # 1745640 and #1945136.
+Ting Ye was partially supported by the HIV Prevention Trials Network (HPTN) and NIH grant: NIAID 5 UM1 AI068617.
+Supplementary material
+The supplementary material includes additional technical details, simulation results, and proofs of all theorems and propositions.
+Supplementary Material
+Section A provides additional results on comparing the four proposed estimators and another martingale-corrected weighting estimator, discussed in Section 4.3 of the main paper.
+Section B reports detailed results from Monte Carlo simulations that assess the finite-sample performance of our proposed estimators.
+Section C provides proofs of all theorems and propositions.
+Appendix A Additional results
+A.1 Pairwise differences between the estimators
+[…]
+[…]
+[…]
+[…]
+These probability limits indicate the impact of nuisance model misspecification on the pairwise comparisons.
+For example, if $G_{1}(t\mid X)$ is correctly specified, the expected difference between $\hat{\mu}_{1}^{\textup{eif}}$ and $\hat{\mu}_{1}^{\textup{aug}}$ should be small.
+Similar logic applies to other comparisons.
+Therefore, when implementing all four estimators in practice, the observed difference among them may serve as an informal diagnostic tool to detect possible nuisance model misspecification.
+A.2 The martingale-corrected weighting estimator
+In this subsection, we present the explicit form of the martingale-corrected weighting estimator introduced in Section 4.3.
+Drawing from the proof of Theorem 2 in Section C.3, if we do not project onto the propensity score nuisance tangent space in Step 2, the following identification formula for $\mu_{1}$ holds:
+$\mu_{1}$ $=$ $E\left[\frac{A}{e(X)}\left\{\frac{Y1(T\wedge C>k)}{G_{1}(k\mid X)}+\mu_{1}(X)S_{1}(k\mid X)\int_{0}^{\tilde{T}}\frac{\mathrm{~{}d}M_{G_{1}}(t)}{S_{1}(t\mid X)G_{1}(t\mid X)}\right\}\right],$
+which motivates the following estimator for $\mu_{1}$:
+$\hat{\mu}_{1}^{\textup{mc-ipw}}$ $=$ $n^{-1}\sum_{i=1}^{n}\frac{A_{i}Y_{i}1(T_{i}\wedge C_{i}>k)}{\hat{e}(X_{i})\hat{G}_{1}(k\mid X_{i})}+n^{-1}\sum_{i=1}^{n}\frac{A_{i}}{\hat{e}(X_{i})}\hat{\mu}_{1}(X_{i})\hat{S}_{1}(k\mid X_{i})\int_{0}^{\tilde{T}_{i}}\frac{\textnormal{d}M_{\hat{G}_{1}}(t)}{\hat{S}_{1}(t\mid X_{i})\hat{G}_{1}(t\mid X_{i})}.$
+The control counterpart, $\hat{\mu}_{0}^{\textup{mc-ipw}}$, is defined analogously, and the estimator for $\tau$ is given by $\hat{\tau}^{\textup{mc-ipw}}=\hat{\mu}_{1}^{\textup{mc-ipw}}-\hat{\mu}_{0}^{\textup{mc-ipw}}$.
+Under Assumuptions 1 and 2, and assuming correct specification of $e(X)$, the estimator $\hat{\tau}^{\textup{mc-ipw}}$ is consistent for $\tau$ if either $G_{a}(t\mid X)$ is correct, or both $\mu_{a}(X)$ and $S_{a}(t\mid X)$ are correct for $t\leq k$ and $a=0,1$.
+Although the martingale correction enhances robustness to misspecification of $G_{a}(t\mid X)$, $\hat{\tau}^{\textup{mc-ipw}}$ is still less attractive than $\hat{\tau}^{\textup{eif}}$: it does not achieve the efficiency bound and is not robust to misspecification of $e(X)$.
+Moreover, it does not inherit the implementation simplicity of $\hat{\tau}^{\textup{aug}}$.
+In particular, it requires estimating the full survival curves for both the treatment-unrelated ICEs $G_{a}(t\mid X)$ and treatment-related ICEs $S_{a}(t\mid X)$ over $t\leq k$, making it computationally demanding without offering clear practical advantages.
+Appendix B Simulation
 In this section, we conduct Monte Carlo simulations to study the finite sample performance of our proposed estimators.
+B.1 Data generating processes
 We first generate the covariates $X=(X_{1},X_{2},X_{3})^{\tiny\textsc{t}}\in\mathbb{R}^{3}$ from three independent standard Gaussian distributions and denote $\tilde{X}_{j}=\{(X_{j}+2)^{2}-1\}/\sqrt{12}$ for $j=1,2,3$.
 Next, we generate the treatment assignment following $A\mid X\sim\textup{Bernoulli}(\{e(X)\})$ with $e(X)$ being the propensity score model, and generate the potential outcomes following $Y(a)\mid X\sim\mathcal{N}(\mu_{a}(X),\sigma_{a}^{2})$ with $\mu_{a}(X)$ being the outcome model for $a=0,1$.
-We then generate the potential values of the treatment-related ICE survival time $T(a)$ from the distribution with a survival function $S_{a}(t\mid X)=\exp[-0.002t^{1.2}\exp\{\gamma_{a}(X)\}]$ and the potential values of the treatment-unrelated ICE survival time $C(a)$ from the distributions with a survival function $G_{a}(t\mid X)=\exp[-\varrho_{a}(t)\exp\{\delta_{a}(X)\}]$, for $a=0,1$.
+We then generate the potential values of the treatment-related ICE survival time $T(a)$ from the distribution with a survival function
+$S_{a}(t\mid X)=\exp[-0.002t^{1.2}\exp\{\gamma_{a}(X)\}]$
+and the potential values of the treatment-unrelated ICE survival time $C(a)$ from the distributions with a survival function
+$G_{a}(t\mid X)=\exp[-\varrho_{a}(t)\exp\{\delta_{a}(X)\}],$
+for $a=0,1$.
 For each of the nuisance parameters, we consider two scenarios when it will be correctly modeled and misspecified, therefore, we generate data following two different choices of each.
-We summarize the detailed data-generating choices in Table 2.
-Table 2: Model choices for the nuisance parameters
+We summarize the detailed data-generating choices in Table S1.
+Table S1: Model choices for the nuisance parameters
 Correctly specified Misspecified
@@
 Generate the observed event time as $T\wedge C\wedge k$, the observed event type indicator, and the observed outcome $Y$ if $T\wedge C>k$.
+B.2 Simulation results
 We compare the finite sample performance of the outcome estimator $\hat{\tau}^{\textup{out}}$, the weighting estimator $\hat{\tau}^{\textup{ipw}}$, the augmented weighting estimator $\hat{\tau}^{\textup{aug}}$, and the EIF estimator $\hat{\tau}^{\textup{eif}}$ based on $1000$ Monte Carlo samples with a sample size of $n=1000$.
@@
 We use $X$ in all regression models, therefore, when the true data-generating processes involve non-linear functions of $X$, the fitted nuisance models suffer from misspecification.
-We evaluate the performance of the four estimators with reported bias, standard deviation, and coverage probability using a nonparametric bootstrap with $200$ bootstrap samples in Table 3.
-Table 3: Finite sample performance of the four estimators.
+We evaluate the performance of the four estimators with reported bias, standard deviation, and coverage probability using a nonparametric bootstrap with $200$ bootstrap samples in Table S2.
+Table S2: Finite sample performance of the four estimators.
 For each estimator, we report the finite-sample bias, standard deviation (SD), and the coverage rate (CR) of a 95% confidence interval, which is constructed using a nonparametric bootstrap with 200 bootstrap iterations.
@@
 The second method naively applies a hypothetical strategy to all ICEs, assuming they are independent of potential outcomes conditional on observed covariates.
-We generate data under the “$\mu\_S$_wrong regime, ensuring correct specification of both the propensity score $e(X)$ and the treatment-unrelated ICE survival model $G_{a}(t\mid X)$ for $a=0,1$.
+We generate data under the “$\mu\_S$_wrong” regime, ensuring correct specification of both the propensity score $e(X)$ and the treatment-unrelated ICE survival model $G_{a}(t\mid X)$ for $a=0,1$.
 Based on $1000$ Monte Carlo simulations with a sample size of $n=1000$, the inverse probability-weighting estimator using the non-responder imputation method exhibits a finite sample bias of $-1.195$ with a standard deviation of $0.077$, and that using the hypothetical strategy for all ICEs has a bias of $1.188$ with a standard deviation of $0.388$.
 Both ad-hoc methods yield inconsistent estimators.
-6 Application
-In this section, we re-analyze data from two immunology trials to study the effect of baricitinib on Systemic Lupus Erythematosus (Morand et al.,, 2023; Petri et al.,, 2023).
-Both trials are double-blinded, randomized, placebo-controlled phase-3 trials.
-There are three treatment arms: 2mg baritinib, 4mg baritinib, and placebo.
-The primary endpoint of interest in both studies is the Systemic Lupus Erythematosus Responder Index 4 (SRI4) measured at week 52.
-It is a binary composite responder index based on both the level of improvement in disease activity and an indicator of no worsening of the overall condition or the development of substantial disease activity in new organ systems.
-The baseline covariates used in model-fitting include the region each patient is from, the baseline corticosteroid usage, and the baseline value of the Physician Global Assessment score measured at week 0.
-There are significant numbers of observations with ICEs happening in both trials (218/760 and 211/775).
-We classify the observed intercurrent events into two types: those likely to be correlated with the potential outcomes such as adverse events (82.6% and 84.4%), and those conditionally independent of the potential outcomes given covariates and the treatment arm such as withdrawal due to relocation (17.4% and 15.6%).
-We compare both 2mg and 4mg treatment arms with the placebo group separately.
-We implemented all four estimators $\hat{\tau}^{\textup{out}}$, $\hat{\tau}^{\textup{ipw}}$, $\hat{\tau}^{\textup{aug}}$, and $\hat{\tau}^{\textup{eif}}$.
-We used logistic regression for estimating the propensity score and outcome models and used Cox proportional hazards regression to estimate the survival function and censoring model.
-We used the nonparametric bootstrap with 500 iterations to compute the variance estimator and conduct inference.
-We summarize the point estimators, their estimated standard errors, and the $p$-values in Table 4.
-Table 4: Treatment effect of Baricitinib on the primary endpoint SRI4.
-The first four columns report the results of Trial 1 and the last four columns report those of Trial 2.
-The first three rows report the point estimators, standard errors, and the $p$-value of the composite outcome average treatment effect corresponding to the 2mg treatment arm, and the last three rows report those of the 4mg treatment arm.
-For each trial and each treatment arm, we report the results based on all four estimators.
-Trial 1 (Petri et al.,, 2023) Trial 2 (Morand et al.,, 2023)
-$\hat{\tau}^{\textup{out}}$ $\hat{\tau}^{\textup{ipw}}$ $\hat{\tau}^{\textup{aug}}$ $\hat{\tau}^{\textup{eif}}$ $\hat{\tau}^{\textup{out}}$ $\hat{\tau}^{\textup{ipw}}$ $\hat{\tau}^{\textup{aug}}$ $\hat{\tau}^{\textup{eif}}$
-2mg point 0.030 0.029 0.026 0.026 0.019 0.019 0.022 0.022
-se 0.042 0.043 0.042 0.043 0.042 0.043 0.043 0.042
-$p$-value 0.479 0.504 0.534 0.540 0.643 0.662 0.602 0.606
-4mg point 0.113 0.120 0.115 0.113 $-0.002$ $-0.002$ $-0.002$ $-0.002$
-se 0.046 0.046 0.046 0.046 0.042 0.042 0.042 0.042
-$p$-value 0.013 0.008 0.012 0.013 0.961 0.962 0.966 0.969
-In both trials and both treatment arms, the point estimators are close to each other in most of the cases, and the statistical inference based on different estimators is remarkably robust across methods.
-Our results suggest a significantly positive effect of 4mg baricitinib treatment on SRI4 in trial 1, although that in trial 2 is insignificant with a slightly negative point estimator.
-Neither trial observes a significant effect of 2mg baricitinib treatment on SRI4.
-Next, we use the two commonly used ad-hoc methods discussed at the end of Section 5 and compare them with our proposed methods.
-Table 5 reports the estimated treatment effect of both treatment arms across the two trials using the inverse probability-weighting estimator.
-The non-responder imputation method yields smaller estimates than our proposed methods most of the time, although the difference is modest due to the relatively low proportion of treatment-unrelated ICEs in both trials.
-Nonetheless, our proposed method targets a clinically more meaningful causal parameter.
-The second ad-hoc method naively applies the hypothetical strategy to all ICEs, assuming they are independent of potential outcomes conditional on the observed covariates.
-Such independence assumption is untestable and likely to be violated in clinical trials.
-For instance, patients discontinued treatment due to adverse events or lack of efficacy.
-As shown in Table 5, this method produces estimated treatment effects substantially different from the obtained results using our proposed methods, leading to misleading clinical conclusions.
-It aligns with the large sample bias it exhibits in the simulation study.
-Table 5: Treatment effect of Baricitinib on the primary endpoint SRI4 using non-responder imputation (NRI) and hypothetical strategy (HS).
-The first two columns report the results of Trial 1 and the last two columns report those of Trial 2.
-The first three rows report the point estimators, standard errors, and the $p$-value of the estimated treatment effect corresponding to the 2mg treatment arm, and the last three rows report those of the 4mg treatment arm.
-For each trial and each treatment arm, we report the results based on the inverse probability-weighting estimator.
-Trial 1 Trial 2
-NRI HS NRI HS
-2mg point 0.039 0.011 0.010 0.022
-se 0.043 0.049 0.042 0.050
-$p$-value 0.362 0.815 0.817 0.654
-4mg point 0.093 0.073 $-0.008$ 0.044
-se 0.044 0.050 0.041 0.051
-$p$-value 0.036 0.147 0.847 0.382
-7 Discussion
-In this study, we address the challenges posed by ICEs in RCTs by proposing methods to handle competing ICEs.
-We classify ICEs into treatment-related and treatment-unrelated events and apply different strategies for the estimation of a clinically meaningful causal effect.
-For treatment-related ICEs, we use the composite outcome approach, assigning values that capture treatment failure.
-For treatment-unrelated ICEs, we adopt a hypothetical strategy that assumes their occurrence is conditionally independent of the potential outcome given covariates.
-Our key insight is that patients discontinue treatment following the first ICE that occurs, which in turn censors the potential occurrence of ICEs of another type.
-In this paper, we rigorously address the issue of competing ICEs by employing a formal causal inference framework.
-We provide identification formulas for the causal parameter of interest, derive its semiparametric efficient influence function, and propose multiple estimators.
-While our proposed framework provides a rigorous and flexible approach to handling competing ICEs, several challenges and extensions remain open for future research.
-7.1 Data collection and ICE classification
-Our proposed methods have broad applicability across various therapeutic areas, including immunology, oncology, and cardiology, where treatment discontinuation and other ICEs frequently occur.
-However, the effectiveness of these approaches relies on the accurate classification of ICEs, which should be performed in collaboration with clinicians to ensure clinical relevance in trial analyses.
-In our application studies, the two types of ICEs were classified based on a manual review of the detailed comments collected at the clinic.
-A modernized case report form for a more accurate collection of reasons for treatment discontinuation is required for each application of the proposed method in clinical trials.
-Fortunately, a cross-industry PHUSE working group is tackling this problem, and a recommended new case report form will be available soon (PHUSE Working Group,, 2024).
-7.2 Random $K$
-In practice, the measurement of the endpoint of interest $Y$ does not necessarily happen at a fixed point $k$.
-For instance, the time $K$ that a patient visits the clinic and takes the measurement can be treated as a random variable independent of treatment $A$, potential outcomes $Y(a)$, and both types of ICEs $T(a)$ and $C(a)$ for $a=0,1$.
-For a random $K$, the analogous formulas to both the previous identification formulas (2) and (3) hold, with a replacement of $k$ to $K$, i.e.,
-$\tau$ $=$ $E\left\{\mu_{1}(X)S_{1}(K\mid X)-\mu_{0}(X)S_{0}(K\mid X)\right\}$ (9)
-$=$ $E\left[\frac{AY1(T\wedge C>K)}{e(X)G_{1}(K\mid X)}-\frac{(1-A)Y1(T\wedge C>K)}{\left\{1-e(X)\right\}G_{0}(K\mid X)}\right].$ (10)
-If $K$ is a pre-treatment covariate that we observe, the previous analyses carry over.
-If $K$ is also treated as a post-treatment variable, we can construct a weighting estimator following (10).
-However, (9) ceases to serve as an identification formula unless $S_{a}(K\mid X)$ is identified for $a=0,1$.
-Considering the presence of three competing time-to-event random variables, it becomes necessary to employ competing risks models.
-We defer this analysis to future work.
-7.3 Comparison between composite variable and principal stratification
-We propose to use a combination of the composite variable and hypothetical strategies to deal with ICEs in this project.
-Following ICH E9 , 2019 (R1), we can also use the combination of principal stratification and hypothetical strategies to deal with the treatment-related and treatment-unrelated ICEs, respectively.
-Using the principal stratification strategy will target different causal parameters and rely on additional identification assumptions such as monotonicity and principal ignorability of the two potential values of the AE time (Ding and Lu,, 2017; Jiang and Ding,, 2021; Jiang et al.,, 2022).
-We leave this for future work.
-Supplementary Material
-We provide proof of all theorems and propositions in the supplementary material.
-Appendix A Proofs
-A.1 Proof of Theorem 1
-First, equation (2) holds because
-LHS $=$ $E[E\{Y(1)1(T(1)>k)\mid X\}]$
+Appendix C Proofs
+C.1 Proof of Theorem 1
+First, equation (2) holds because its left-hand side equals
+$E[E\{Y(1)1(T(1)>k)\mid X\}]$
 $=$ $E[E\{Y1(T>k)\mid X,A=1\}]$
 $=$ $E\{E(Y\mid T>k,X,A=1)\textup{pr}(T>k\mid X,A=1)\}$
-$=$ $E\{E(Y\mid T>k,C>k,X,A=1)\textup{pr}(T>k\mid X,A=1\}=\textup{RHS},$
-where the first equality is by the law of iterated expectations, the second equality is by the randomization assumption 1, and the fourth equality is by the censoring at random assumption 2.
-Next, equation (3) holds because
-RHS $=$ $E\left[E\left\{\frac{AY1(T\wedge C>k)}{e(X)\textup{pr}(C>k\mid X,A=1)}\mid X\right\}\right]$
+$=$ $E\{E(Y\mid T>k,C>k,X,A=1)\textup{pr}(T>k\mid X,A=1\},$
+which is equal to the right-hand side of (2), where the first equality is by the law of iterated expectations, the second equality is by the randomization assumption 1, and the fourth equality is by the censoring at random assumption 2.
+Next, equation (3) holds because its right-hand side equals
+$E\left[E\left\{\frac{AY1(T\wedge C>k)}{e(X)\textup{pr}(C>k\mid X,A=1)}\mid X\right\}\right]$
 $=$ $E\left(\frac{1}{e(X)\textup{pr}(C>k\mid X,A=1)}E\left[AY(1)1\{T(1)\wedge C(1)>k\}\mid X\right]\right)$
@@
 $=$ $E\left(\frac{1}{\textup{pr}(C>k\mid X,A=1)}E[1\{C(1)>k\}\mid X]E\left[Y(1)1\{T(1)>k\}\mid X\right]\right)$
-$=$ $E[E\{Y(1)1(T(1)>k)\mid X\}]=\textup{LHS},$
-where the first equality is by the law of iterated expectations, the third equality is by the randomization assumption 1, the fourth equality is by the censoring at random assumption 2, and the last equality is again by the law of iterated expectations.
+$=$ $E[E\{Y(1)1(T(1)>k)\mid X\}],$
+which is equal to the left-hand side of (3), where the first equality is by the law of iterated expectations, the third equality is by the randomization assumption 1, the fourth equality is by the censoring at random assumption 2, and the last equality is again by the law of iterated expectations.
 $\square$
-A.2 Proof of Proposition 1
-We prove that when the censoring model is correctly specified, $\hat{\tau}^{\textup{aug}}$ is doubly robust in the sense that it is consistent to $\tau$ if either $e^{*}=e$ or $(\mu_{1}^{*}=\mu_{1},S_{1}^{*}=S_{1})$.
-Want to show
+C.2 Proof of Proposition 1
+We prove that under the correct specification of the censoring model, $\hat{\tau}^{\textup{aug}}$ is doubly robust in the sense that it is consistent for $\tau$ if either $e^{*}=e$ or $(\mu_{1}^{*}=\mu_{1},S_{1}^{*}=S_{1})$.
+We have
+$E\left\{\frac{AY1(T\wedge C>k)}{e^{*}(X)G_{1}(k\mid X)}-\frac{A-e^{*}(X)}{e^{*}(X)}\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)\right\}$ (S1)
+$=$ $E\left[\frac{E\{AY1(T\wedge C>k)\mid X\}}{e^{*}(X)G_{1}(k\mid X)}-\frac{e(X)-e^{*}(X)}{e^{*}(X)}\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)\right]$
+$=$ $E\left\{\frac{e(X)}{e^{*}(X)}\mu_{1}(X)S_{1}(k\mid X)-\frac{e(X)-e^{*}(X)}{e^{*}(X)}\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)\right\},$
+where the first equality follows from the law of integrated expectations and the second equality follows from similar derivations as in the proof of Theorem 1.
+Now observe that the final expression of (S1) is equal to $\mu_{1}$ if either $e^{*}=e$ or $(\mu_{1}^{*}=\mu_{1},S_{1}^{*}=S_{1})$.
+Therefore, we have
 $E\left\{\frac{AY1(T\wedge C>k)}{e^{*}(X)G_{1}(k\mid X)}-\frac{A-e^{*}(X)}{e^{*}(X)}\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)-\mu_{1}\right\}$ $=$ 00 0
 if either $e^{*}=e$ or $(\mu_{1}^{*}=\mu_{1},S_{1}^{*}=S_{1})$ holds.
-$E\left\{\frac{AY1(T\wedge C>k)}{e^{*}(X)G_{1}(k\mid X)}-\frac{A-e^{*}(X)}{e^{*}(X)}\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)\right\}$
-$=$ $E\left[\frac{E\{AY1(T\wedge C>k)\mid X\}}{e^{*}(X)G_{1}(k\mid X)}-\frac{e(X)-e^{*}(X)}{e^{*}(X)}\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)\right]$
-$=$ $E\left\{\frac{e(X)}{e^{*}(X)}\mu_{1}(X)S_{1}(k\mid X)-\frac{e(X)-e^{*}(X)}{e^{*}(X)}\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)\right\},$
-where the first equality is by the law of integrated expectations and the second equality follows from similar derivations as in the proof of Theorem 1.
-The last line is equal to $\mu_{1}$ if either $e^{*}=e$ or $(\mu_{1}^{*}=\mu_{1},S_{1}^{*}=S_{1})$.
+A parallel argument applies to the control counterpart.
+Therefore, combining the treated and control components, we conclude that $\hat{\tau}^{\textup{aug}}$ is consistent for $\tau$ if either the propensity score model is correct, or the outcome and survival models for the treatment-related ICE are both correct.
+This establishes the double robustness of $\hat{\tau}^{\textup{aug}}$.
 $\square$
-A.3 Proof of Theorem 2
+C.3 Proof of Theorem 2
 We follow the semiparametric theory in Bickel et al., (1993) to derive the EIF for $\mu_{1}$.
@@
 Following the steps in Hubbard et al., (2000), we first consider the case when every observation is assigned to the treatment group, i.e., when there is no missingness generated by the treatment assignment, and then project the derived EIF onto the nuisance tangent space of the propensity score to get the final form of EIF for $\mu_{1}$.
+Step 1.
 With full data $(X,Y^{\textup{cp}}(1),Y^{\textup{cp}}(0))$, the estimating equation for $\mu_{1}$ is $D_{\textup{full}}=Y^{\textup{cp}}(1)-\mu_{1}$.
@@
 $D$ $=$ $\frac{\Delta Y^{\textup{cp}}}{G(\tilde{T}\mid X)}+E(Y^{\textup{cp}}\mid X)\int_{0}^{\tilde{T}}\frac{\textnormal{d}M_{G}(t)}{S(t\mid X)G(t\mid X)}-\mu_{1}$ (S2)
+Step 2.
 Next, we follow steps in Section 3 of Hubbard et al., (2000) and compute the EIF when the real observed data is $O=(X,A,\Delta,\tilde{T},\Delta Y^{\textup{cp}})$.
@@
 has the given form in Theorem 2. $\square$
-A.4 Proof of Theorem 3
+C.4 Proof of Theorem 3
 We prove the result by showing
-$\mu_{1}=E\left[\frac{A}{e^{*}(X)}\left\{\frac{Y1(T\wedge C>k)}{G_{1}^{*}(k\mid X)}+\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)\int_{0}^{\tilde{T}}\frac{\textnormal{d}M_{G_{1}^{*}}(t)}{S_{1}^{*}(t\mid X)G_{1}^{*}(t\mid X)}\right\}-\frac{A-e^{*}(X)}{e^{*}(X)}\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)\right]$ (S2)
-if either $\{e^{*}(X)=e(X),G_{1}^{*}(t\mid X)=G_{1}(t\mid X)\}$ or $\{\mu_{1}^{*}(X)=\mu_{1}(X),S_{1}^{*}(t\mid X)=S_{1}(t\mid X)\}$ for $t\leq k$.
-Denote
+$\mu_{1}$ $=$ $E\left[\frac{A}{e^{*}(X)}\left\{\frac{Y1(T\wedge C>k)}{G_{1}^{*}(k\mid X)}+\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)\int_{0}^{\tilde{T}}\frac{\textnormal{d}M_{G_{1}^{*}}(t)}{S_{1}^{*}(t\mid X)G_{1}^{*}(t\mid X)}\right\}\right.$ (S3)
+$\quad\ \left.-\frac{A-e^{*}(X)}{e^{*}(X)}\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)\right]$
+if either $\{e^{*}(X)=e(X),G_{1}^{*}(t\mid X)=G_{1}(t\mid X)\}$ or $\{\mu_{1}^{*}(X)=\mu_{1}(X),S_{1}^{*}(t\mid X)=S_{1}(t\mid X)\}$ for $t\leq k$. Define
 $\mathcal{T}_{1}$ $=$ $\frac{A}{e^{*}(X)}\frac{Y1(T\wedge C>k)}{G_{1}^{*}(k\mid X)}-\frac{A-e^{*}(X)}{e^{*}(X)}\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)-\mu_{1}(X)S_{1}(k\mid X),$
@@
 $=$ $S_{1}(t\mid X)G_{1}(t\mid X)\{\textnormal{d}\Lambda_{1}(t\mid X)-\textnormal{d}\Lambda_{1}^{*}(t\mid X)\},$
-therefore,
+and therefore,
 $E(\mathcal{T}_{2}\mid X)$ $=$ $\frac{e(X)}{e^{*}(X)}\mu_{1}^{*}(X)S_{1}^{*}(k\mid X)E\left\{\int_{0}^{k}\frac{\textnormal{d}M_{G_{1}^{*}}(t)}{S_{1}^{*}(t\mid X)G_{1}^{*}(t\mid X)}\mid X\right\}$
@@
 By symmetry, the part corresponding to the control arm $a=0$ also holds, thus the double robustness in Theorem 3 holds. $\square$
-A.5 Proof of Theorem 4
+C.5 Proof of Theorem 4
 Following the von Mises expansion (Hines et al.,, 2022), let $P$ and $P_{n}$ denote the true and empirical distribution of the observed data, respectively, and $\hat{P}_{n}$ denotes an estimated $P$, we have
@@
 In the following discussion, we suppress the dependency of the nuisance functions and their estimated values on $X$ for ease of notation.
-By the facts that $e\hat{\mu}_{1}\hat{S}_{1}(k)/\hat{e}$ is bounded and $1/\{\hat{S}_{1}(t)\hat{G}_{1}(t)\}$ is bounded for any $t\leq k$, we employ the Cauchy-Schwarz inequality to upper bound the term $R_{11}$ by a constant times $\int_{0}^{k}||S_{1}(t)-\hat{S}_{1}(t)||_{2}||G_{1}(t)-\hat{G}_{1}(t)||_{2}\textnormal{d}t$, where $||\cdot||_{2}$ denotes the $L_{2}(P)$ norm.
+By the facts that $e\hat{\mu}_{1}\hat{S}_{1}(k)/\hat{e}$ is bounded and $1/\{\hat{S}_{1}(t)\hat{G}_{1}(t)\}$ is bounded for any $t\leq k$, we employ the Cauchy-Schwarz inequality to upper bound the term $R_{11}$ by a constant times
+$\int_{0}^{k}||S_{1}(t)-\hat{S}_{1}(t)||_{2}||G_{1}(t)-\hat{G}_{1}(t)||_{2}\textnormal{d}t,$
+where $||\cdot||_{2}$ denotes the $L_{2}(P)$ norm.
 It follows that Assumption 5 guarantees $R_{11}=o_{P}(n^{-1/2})$.
-For $R_{12}$, we further suppress the dependency of $S_{1}(k\mid X)$, $G_{1}(k\mid X)$ and their estimators on both $k$ and $X$, and rewrite the term as $R_{12}=E[\{\hat{e}(G_{1}-\hat{G}_{1})+G_{1}(e-\hat{e})\}\{\hat{\mu}_{1}(S_{1}-\hat{S}_{1})+S_{1}(\mu_{1}-\hat{\mu}_{1})\}/(\hat{e}\hat{G}_{1})]$.
+For $R_{12}$, we further suppress the dependency of $S_{1}(k\mid X)$, $G_{1}(k\mid X)$ and their estimators on both $k$ and $X$, and rewrite the term as
+$R_{12}=E[\{\hat{e}(G_{1}-\hat{G}_{1})+G_{1}(e-\hat{e})\}\{\hat{\mu}_{1}(S_{1}-\hat{S}_{1})+S_{1}(\mu_{1}-\hat{\mu}_{1})\}/(\hat{e}\hat{G}_{1})].$
 Because $\hat{e}/\hat{G}_{1}$ is bounded, by the Cauchy-Schwarz inequality, the first term in $R_{12}$ is upper bounded by $||G_{1}-\hat{G}_{1}||_{2}||S_{1}-\hat{S}_{1}||_{2}=o_{P}(n^{-1/2})$ by Assumption 5.
@@
 Thus, $R_{1}(P,\hat{P}_{n})=o_{P}(n^{-1/2})$.
-Symmetric arguments imply the analogous term for the control arm satisfies $R_{0}(P,\hat{P}_{n})=o_{P}(n^{-1/2})$. Therefore, results in Theorem 4 follow from the central limit theorem. $\square$
+Symmetric arguments imply the analogous term for the control arm satisfies $R_{0}(P,\hat{P}_{n})=o_{P}(n^{-1/2})$. Therefore, the results in Theorem 4 follow from the central limit theorem. $\square$